### v1

# Debiased machine learning for combining probability and non-probability survey data

Abstract
We consider the problem of estimating the finite population mean $\bar{Y}$ of an outcome variable $Y$ using data from a nonprobability sample and auxiliary information from a probability sample. Existing double robust (DR) estimators of this mean $\bar{Y}$ require the estimation of two nuisance functions: the conditional probability of selection into the nonprobability sample given covariates $X$ that are observed in both samples, and the conditional expectation of $Y$ given $X$. These nuisance functions can be estimated using parametric models, but the resulting estimator of $\bar{Y}$ will typically be biased if both parametric models are misspecified. It would therefore be advantageous to be able to use more flexible data-adaptive / machine-learning estimators of the nuisance functions. Here, we develop a general framework for the valid use of DR estimators of $\bar{Y}$ when the design of the probability sample uses sampling without replacement at the first stage and data-adaptive / machine-learning estimators are used for the nuisance functions. We prove that several DR estimators of $\bar{Y}$, including targeted maximum likelihood estimators, are asymptotically normally distributed when the estimators of the nuisance functions converge faster than the $n^{1/4}$ rate and cross-fitting is used. We present a simulation study that demonstrates good performance of these DR estimators compared to the corresponding DR estimators that rely on at least one correctly specified parametric model.
Keywords: cross-fitting; data-adaptive; design-based inference; double robust; finite population; machine learning; sample surveys; targeted maximum likelihood.
1 Introduction
Probability sampling methods are the gold standard for conducting surveys. They are designed to yield samples that are representative of the finite population from which they are drawn. Nevertheless, there has been considerable interest in using data from nonprobability samples, e.g. web-based volunteer surveys, due to their increasing availability and the ease and relatively low cost with which such data can be collected. Such samples may, however, not be representive of the population, leading to bias when using them to estimate the finite population mean $\bar{Y}$ of some variable $Y$ of interest.
Several methods have recently been proposed for using data from a probability sample (Sample A) as auxiliary information to address nonrepresentative of a nonprobability sample (Sample B). These methods fall into three classes, all of which require data on some covariates $X$ that are observed on the individuals in both samples. The first class of methods, which include inverse probability weighting (IPW) methods, involve estimating weights for the individuals in Sample B such that the weighted Sample B is representative of the population. The second class, known as mass imputation methods, involve using the relation between $X$ and $Y$ in Sample B to impute the $Y$ values in Sample A. The third class, known as double robust (DR) methods, combine IPW and mass imputation. IPW methods require a consistent estimator of the selection probability function for Sample B, i.e. the conditional probability that an individual belongs to Sample B given $X$. Mass imputation methods require a consistent estimator of the conditional expectation of $Y$ given $X$. We shall refer to this conditional probability and expectation as ‘nuisance’ functions, because they are not of direct interest. DR methods use estimators of both nuisance functions but only require one of these estimators to be consistent.
A parametric model could be used for each of the two nuisance functions, but the resulting estimator would typically be inconsistent if that parametric model — or, in the case of DR estimators, both parametric models — were misspecified. There is therefore considerable interest in using more flexible ‘data-adaptive’ or ‘machine-learning’ estimators of the nuisance functions, in order to minimise this risk of model misspecification. A number of researchers have done this. Ferri-Garcia et al. (2020, 2024)[16, 17], Rueda et al. (2023)[13] and Rueda et al. (2024)[12] use IPW and estimate the conditional probability nuisance function with k-nearest neighbours, random forest or XGboost. Castro-Martin et al. (2020)[2] use IPW or mass imputation, and estimate the conditional probability or expectation nuisance function using random forest or XGboost. Ferri-Garcia et al. (2022)[15] use IPW and estimate the nuisance function using classification and regression trees (CART). Castro-Martin et al. (2021)[3] use IPW, mass imputation or a DR estimator, and estimate the nuisance functions (or functions) using XGboost. Cobo et al. (2025)[11] use a DR estimator and estimate the nuisance functions using XGboost, k-nearest neighbours or neural nets. To reduce the variability of the estimated inverse probability weights, Castro-Martin et al. (2021)[3], Rueda et al. (2024)[12] and Cobo et al. (2025)[11] apply kernel smoothing to them. Wang et al. (2020,2022)[27, 28] developed a theoretical justification for this kernel smoothing approach when logistic regression is used, but not when data-adaptive methods are used. Chen et al. (2022)[5] use mass imputation with kernel smoothing, show that the resulting estimator of $\bar{Y}$ is asymptotically normal and provide a formula for the variance. They note, however, that this method is subject to the curse of dimensionality.
A closely related problem to that of estimating the mean of $Y$ in a finite population by combining data from probability and nonprobability samples is that of estimating $E(Y)$, the expectation of $Y$ in an infinite population, when data $(X_{i},Y_{i})$ $(i=1,\ldots,n)$ are independent and identically distributed (iid), $X$ is fully observed and $Y$ is missing at random given $X$. IPW, imputation and DR methods exist for this problem. The first two require consistent estimators of, respectively, the conditional probability that $Y$ is observed given $X$ and the conditional expectation of $Y$ given $X$. DR methods involve estimating both these nuisance functions, but require only that (at least) one of them is estimated consistently. For IPW and imputation, it is known that using data-adaptive methods to estimate the nuisance function can cause problems. Data-adaptive methods typically yield nuisance function estimators that converge slowly, and this slow convergence is inherited by the IPW or imputation estimator of $E(Y)$, which can then have considerable finite-sample bias; confidence intervals can also have poor coverage[26]. The DR approach (also known as ‘debiased machine learning’) can overcome this issue, because DR estimators do not inherit the slow convergence of the nuisance function estimators[26, 14, 23, 20, 8, 7, 9]. Given the similarity of the problem of estimating $E(Y)$ for an infinite population with iid data to the problem of estimating the mean of $Y$ in a finite population using a nonprobability sample, there is reason to be concerned that the same issue may well also apply to the latter problem. In this article, we study the use of a DR estimator of the mean of $Y$ in a finite population when combining data from a probability sample and a nonprobability sample and using data-adaptive estimators of the nuisance functions. Our goal is to provide a theoretical justification for the validity of this approach. An important difference between the problem of estimating $E(Y)$ for an infinite population with iid data and the problem we study here is that when, as is common, the probability sample is obtained using sampling without replacement, Sample A data are not iid.
We build on the work of Chen, Li and Wu (2020)[6] (henceforth, CLW). CLW developed a DR estimator for $\bar{Y}$ using parametric models for the nuisance functions. Yang et al. (2020)[29] extended this method to allow for variable selection by smoothly clipped absolute deviation (SCAD). We generalise the work of CLW (and Yang et al.) to allow for general data-adaptive estimation of the nuisance functions. Our approach uses cross-fitting, a technique popular in the field of debiased machine learning.
The structure of the article is as follows. In Section 2, we describe our design-model-based inference framework. In Section 3, we describe a von Mises expansion and cross-fitting, both fundamental to our approach. In Section 4, we consider the scenario where clusters (or ‘primary sampling units’) are sampled using simple random sampling (i.e. with equal probabilities) without replacement and then individuals within clusters are sampled using an arbitrary (possibly multi-stage) sampling design that obeys the independence and invariance conditions[24]. Section 5 considers the more general scenario where the cluster sampling probabilities vary. In Section 6, we describe a ratio-style DR estimator of $\bar{Y}$ that is typically more efficient than the DR estimator described in earlier sections. Section 7 covers targeted maximum likelihood estimators (TMLEs), which are an alternative to the estimating equations-style DR estimators described in earlier sections. Possible data-adaptive estimators of the nuisance functions are discussed in Section 8. In Section 9, we present a simulation study that compares DR methods using data-adaptive estimators of nuisance functions with those that use parametric estimators. In this article, we do not use data on the $Y$ values of individuals in Sample A. When such data are available, $\bar{Y}$ can be estimated using only Sample A, without needing Sample B. However, efficiency may be gained by combining the DR estimator that uses the $Y$ values only of individuals in Sample B with the estimator that uses only Sample A[18, 13, 25]. This, along with other issues, is discussed in Section 10.
2 Data-generating mechanism and estimator
Our assumed data-generating mechanism consists of a superpopulation model for generating the finite population, a model for drawing Sample A from the finite population, and a model for drawing Sample B from the finite population. The superpopulation model generates a finite population of $J$ clusters (or ‘primary sampling units’) as follows. Cluster sizes $N_{1},\ldots,N_{J}$ are independently generated from a distribution $f(n)$. For each cluster $j$ ($j=1,\ldots,J$) independently, a $N_{j}\times p$ covariate matrix $(X_{j1},\ldots,X_{jN_{j}})$ is generated conditionally on $N_{j}$ from a distribution $f(x_{1},\ldots,x_{N}\mid N=N_{j})$. For each individual $i=1,\ldots,N_{j}$ in each cluster $j=1,\ldots,J$, an outcome $Y_{ji}$ is independently generated from a distribution $f(y\mid X=X_{ji})$ with expectation $m_{0}(X_{ji})=E(Y\mid X=X_{ji})$. Let $\mathcal{F}_{J}=(N_{1},\ldots,N_{J},X_{11},\ldots$, $X_{1N_{1}},\ldots,X_{J1},\ldots,X_{JN_{J}})$. Our goal will be to estimate the population mean of $Y$, i.e. $\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}Y_{ji}\left/\sum_{j=1}^{J}N_{j}\right$..
Note that, as is usual in model-design-based inference for the mean of $Y$, we shall consider repeated-sampling properties of an estimator of this mean conditional on $\mathcal{F}_{J}$[6, 22]. We have assumed the data-generating mechanism for $\mathcal{F}_{J}$ described above only for the asymptotics; this mechanism describes how the finite population grows as $J\rightarrow\infty$.
The model for drawing Sample A from this finite population is as follows. The variable $R^{A}_{ji}$ will be a binary variable indicating whether individual $i$ in cluster $j$ ($i=1,\ldots,N_{j};\;j=1,\ldots,J$) is included in Sample A. First, a sample of size $M$ ($M<J$) clusters is drawn from the $J$ clusters in the population, with the probability that cluster $j$ is included in the sample being proportional to $h(N_{j},X_{j1},\ldots,X_{jN_{j}})$ for some function $h$ of $N_{j}$ and $X_{j1},\ldots,X_{jN_{i}}$. A possible example would be $h(N_{j},X_{j1},\ldots,X_{jN_{j}})=N_{j}$. Let $R^{C}_{j}=1$ if cluster $j$ is drawn, and $R^{C}_{j}=0$ otherwise. If $R^{C}_{j}=0$, then $R^{A}_{ji}=0$ ($i=1,\ldots,N_{j}$). If $R^{C}_{j}=1$, then $(R^{A}_{j1},\ldots,R^{A}_{jN_{j}})$ is drawn according to some (possibly multistage) sampling design that can depend on $N_{j}$ and $(X_{j1},\ldots,X_{jN_{j}})$ but which, given $N_{j}$, $(X_{j1},\ldots,X_{jN_{j}})$ and $R^{C}_{j}=1$, does not depend on $(Y_{j1},\ldots,Y_{jN_{j}})$, $\{(N_{k},X_{k1},\ldots,X_{kN_{k}},Y_{k1},\ldots,Y_{kN_{k}}):\;k=1,\ldots,j-1,j+1,\ldots J\}$ or $(R^{C}_{1},\ldots,R^{C}_{j-1},R^{C}_{j+1},\ldots,R^{C}_{J}$). This sampling of $(R^{A}_{j1},\ldots,R^{A}_{jN_{j}})$ is done independently of the sampling of $(R^{A}_{k1},\ldots,R^{A}_{kN_{k}})$ within any other cluster $k$ with $R^{C}_{k}=1$. Note this means we are assuming independence and invariance (see pp134–135 of [24]) for the sampling design. Let $\pi^{C}_{j}=P(R^{C}_{j}=1\mid\mathcal{F}_{J})$ and $\pi^{A\mid C}_{ji}=P(R^{A}_{ji}=1\mid R^{C}_{j}=1,N_{j},X_{j1},\ldots,X_{jN_{j}})$. Let $\pi^{A}_{ji}=\pi^{C}_{j}\times\pi^{A\mid C}_{ji}$ denote the first-order inclusion probability for individual $i$ in cluster $j$. For simplicity, we do not consider stratified sampling here, but this is discussed in Section 10.
The model for drawing Sample B from the finite population is as follows. The variable $R^{B}_{ji}$ will be a binary variable indicating whether individual $i$ in cluster $j$ is included in Sample B. Let $\pi^{B}_{0}(X)$ be some function of $X$. Given $\mathcal{F}_{J}$, each $R^{B}_{ji}$ ($i=1,\ldots,N_{j};\;j=1,\ldots,J$) is independently sampled from a Bernoulli distribution with probability parameter $\pi^{B}_{0}(X_{ji})$ independently of $Y_{11},\ldots,Y_{1N_{1}},\ldots Y_{J1},\ldots,Y_{JN_{J}}$, $R^{C}_{1},\ldots,R^{C}_{J}$ and $R^{A}_{11},\ldots,R^{A}_{1N_{1}},\ldots R^{A}_{J1},\ldots,R^{A}_{JN_{J}}$.
Now relabel the $\sum_{j=1}^{J}N_{j}$ individuals so that, for each $j=1,\ldots,J$, the $N_{j}$ individuals in cluster $j$ are labelled as individuals $\sum_{k=1}^{j-1}N_{j}+1,\sum_{k=1}^{j-1}N_{j}+2,\ldots,\sum_{k=1}^{j}N_{j}$. Let $D_{i}$ ($1\leq D_{i}\leq J$) denote the number of the original cluster to which individual $i$ belongs. Finally, let $n=n(\mathcal{F}_{J})=\sum_{j=1}^{J}N_{j}$ denote the total number of individuals in the population. We can now write the population mean as $\bar{Y}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$.
The observed data are
$\{(R^{A}_{i},R^{A}_{i}X_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}D_{i},R^{B}_{i},R^{B}_{i}X_{i},R^{B}Y_{i},R^{B}_{i}D_{i});\;i=1,\ldots,n\},$
i.e. the $X$ values and sampling probabilities of individuals in Sample A, the $X$ and $Y$ values of individuals in Sample B, and the index of the cluster to which each of these individuals belongs. Note that our notation suggests that we know whether an individual in Sample A also appears in Sample B, and vice versa. In fact, this is unnecessary; this notation is used only for convenience. We shall, however, need to know, for each individual in Sample A and each individual in Sample B, whether they belong to the same cluster (this is used in the cross-fitting procedure — see Section 3).
[…]
$\hat{\theta}_{1}(\pi^{B},m)=\frac{1}{n}\sum_{i=1}^{n}U_{i}(\pi^{B},m),$
$U(\pi^{B},m)=U(\pi^{B},m;X,Y,R^{A},R^{B},\pi^{A})=\frac{R^{B}}{\pi^{B}(X)}Y+\left\{\frac{R^{A}}{\pi^{A}}-\frac{R^{B}}{\pi^{B}(X)}\right\}m(X),$
[…]
3 von Mises expansion
3.1 Independent, identically distributed data and infinite population
[…]
$\hat{\theta}_{\rm iid}(\pi_{\rm iid},m_{\rm iid})=\frac{1}{n}\sum_{i=1}^{n}\frac{R_{i}}{\pi_{\rm iid}(X_{i})}Y_{i}+\left\{1-\frac{R_{i}}{\pi_{\rm iid}(X_{i})}\right\}m_{\rm iid}(X_{i}).$
Here, $R$ is an indicator that $Y$ is observed, $\hat{\pi}_{\rm iid}(X)$ is an estimator of $\pi_{\rm iid0}(X)=P(R=1\mid X)$ and $\hat{m}_{\rm iid}(X)$ is an estimator of $m_{\rm iid0}(X)=E(Y\mid X)$. Hines et al. (2022)[20] give an account of the properties of $\hat{\theta}_{\rm iid}$ and of the von Mises expansion that can be used to derive these properties. They describe how, if $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ converge to $\pi_{\rm iid0}$ and $m_{\rm iid0}$ sufficiently rapidly as $n\rightarrow\infty$, then $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})=\hat{\theta}_{\rm iid}(\pi_{\rm iid0},m_{\rm iid0})+o_{p}(n^{-1/2})$, and so $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$ has the same asymptotic normal distribution as $\hat{\theta}_{\rm iid}(\pi_{\rm iid0},m_{\rm iid0})$. The required convergence rates of $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ are slower than the parametric $n^{1/2}$ rate, which enables the use of data-adaptive estimators. Often, $\hat{\theta}_{\rm iid}$ is used with cross-fitting, a technique similar to cross-validation. This involves splitting the sample of $n$ individuals into subsamples, called folds, calculating estimates $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ using the data on all but fold $k$, and then calculating $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$ using these estimates and the data on fold $k$. This procedure is repeated for $k=1,\ldots,K$, and the resulting $K$ values of $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$ are averaged. This ensures that $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ are independent of the data in fold $k$, which avoids the need for $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ to satisfy the Donsker condition[20].
3.2 Sampling without replacement from a finite population
In this article, we obtain an analogous asymptotic equivalence result for our sample survey setting. We make use of cross-fitting. Randomly partition the set of $n$ individuals into $K$ (e.g. $K=5$) subsets, called folds. The way in which is done will be described later. Let $\mathcal{S}_{k}\subset\{1,\ldots,n\}$ denotes the set of indices of the individuals belonging to fold $k$ ($k=1,\ldots,K$). Let $n_{k}=|\mathcal{S}_{k}|$ denote the number of individuals in $\mathcal{S}_{k}$. Let $\mathcal{S}_{-k}$ denote the indices of all individuals except those in fold $k$ (so, $\mathcal{S}_{k}\cup\mathcal{S}_{-k}=\{1,\ldots,n\}$). Let $\hat{m}_{k}$ be an estimator of $m_{0}$ calculated using only the data $\{(X_{i},Y_{i}):R^{B}_{i}=1\mbox{ and }i\in\mathcal{S}_{-k}\}$. Let $\hat{\pi}^{B}_{k}$ be an estimator of $\pi^{B}_{0}$ obtained using only the data $\{(R^{A}_{i},R^{B}_{i}):i\in\mathcal{S}_{-k}\}$ and $\{X_{i}:R^{A}_{i}+R^{B}_{i}\geq 1\mbox{ and }i\in\mathcal{S}_{-k}\}$. Write $\underline{\pi}^{B}=(\pi^{B}_{1},\ldots,\pi^{B}_{K})$, $\underline{m}=(m_{1},\ldots,m_{K})$, $\underline{\hat{\pi}}^{B}=(\hat{\pi}^{B}_{1},\ldots,\hat{\pi}^{B}_{K})$ and $\underline{\hat{m}}=(\hat{m}_{1},\ldots,\hat{m}_{K})$. In a slight abuse of notation, we define
$\hat{\theta}_{1}(\underline{\pi}^{B},\underline{m})=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U_{i}(\pi^{B}_{k},m_{k}).$
[…]
Condition C1: There exist $c_{\pi}>0$ and $c_{m}>0$ such that $c_{\pi}+c_{m}=1$,
$E_{X}\left[\left\{\frac{\pi^{B}_{0}(X)}{\hat{\pi}^{B}_{k}(X)}-1\right\}^{2}\right]=o_{p}(M^{-c_{\pi}})$ (1)
and
$E_{X}\left[\{\hat{m}_{k}(X)-m_{0}(X)\}^{2}\right]=o_{p}(M^{-c_{m}}).$ (2)
Condition C1 allows convergence rates that are considerably slower than those of parametric estimators [20]. For example, Condition C1 is satisfied when both $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ converge faster than the $M^{1/4}$ rate.
Our proof uses the following von Mises expansion of $\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}$.
$\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}$
$\hskip 14.22636pt=\frac{1}{n}\sum_{i=1}^{n}\{U_{i}(\pi^{B}_{0},m_{0})-\bar{Y}\}$
$\hskip 28.45274pt+\sum_{k=1}^{K}\frac{n_{k}}{n}\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\left[\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\right$.
$\hskip 113.81102pt-\left.E\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}\right]$
$\hskip 28.45274pt+\sum_{k=1}^{K}\frac{n_{k}}{n}\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}E\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}.$
The first term on the right-hand side of equation (3) is just $\hat{\theta}_{1}(\pi^{B}_{0},m_{0})-\bar{Y}$. The second and third terms will be referred to as, respectively, the ‘empirical process term’ and the ‘remainder term’. We shall prove that these two terms are $o_{p}(M^{-1/2})$ provided that Condition C1 holds. It follows from this that
$\sqrt{M}\{\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}\}$
$\hskip 28.45274pt=\sqrt{M}\{\hat{\theta}_{1}(\pi^{B}_{0},m_{0})-\bar{Y}\}+o_{p}(1)$
$\hskip 28.45274pt=\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}m_{0}(X_{i})+\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{B}_{i}}{\pi^{B}_{0}(X_{i})}\{Y_{i}-m_{0}(X_{i})\}+o_{p}(1).$
The asymptotic variance of $\hat{\theta}_{1}(\pi^{B}_{0},m_{0})-\bar{Y}$ is given by CLW, along with an estimator of this variance. This is described in more detail by Seaman et al. (2025)[25] (specifically, the formula for $\mu_{\rm DR1}$ when both nuisance models are correctly specified). Moreover, subject to regularity conditions, $\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}$ is asymptotically normally distributed[4].
An important complication when proving that the empirical process and remainder term are $o_{p}(M^{-1/2})$ is that, in general, $\hat{\pi}^{B}_{k}(X)$ is not independent of the data in fold $k$. This is because $\hat{\pi}^{B}_{k}(X)$ is calculated using the $R^{A}_{i}$ values of individuals in $\mathcal{S}_{-k}$, $\sum_{i\in\mathcal{S}_{k}}U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})$ is calculated using the $R^{A}_{i}$ values of individuals in $\mathcal{S}_{k}$, and $\{R^{A}_{i}:\;i\in\mathcal{S}_{-k}\}$ and $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ are, in general, not independent when clusters are sampled without replacement. For example, the number of clusters with $R^{C}_{j}=1$ in fold $k$ must equal $M$ minus the number of clusters with $R^{C}_{j}=1$ in the remaining folds.
4 Simple random sampling of clusters without replacement
Consider the case of simple random sampling of clusters without replacement (so, $\pi^{C}_{1}=\ldots=\pi^{C}_{J}$). As described in Section 2, sampling of individuals within sampled clusters is done using some (possibly multistage) sampling design. Suppose that $M$ and $J$ are both integer multiples of $K$ (the case where $M$ and/or $J$ are not multiples of $K$ is considered as a special case in Section 5).
Randomly partition the $M$ sampled clusters (i.e. those with $R^{C}_{j}=1$) evenly into $K$ folds and, likewise, the JMJ-M unsampled clusters (i.e. those with $R^{C}_{j}=0$) evenly into the $K$ folds, with all such partitions being equally probable. Each fold contains $M/K$ sampled clusters and (JM)/K(J-M)/K unsampled clusters.
This ensures that the conditional distribution of $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J}$, $\mathcal{S}_{k}$ and $\{R^{A}_{i}:\;i\in\mathcal{S}_{-k}\}$ is the same as the distribution of $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$, and corresponds to simple random sampling of $M/K$ clusters without replacement from the $J/K$ clusters in fold $k$ followed by the original second-stage sampling mechanism for individuals within clusters.
Recall that $Y_{i}$ is assumed to be conditionally independent of $R^{A}_{i^{\prime}}$, $R^{C}_{j}$ and $Y_{i^{\prime\prime}}$ given $\mathcal{F}_{J}$ for all $i$, $i^{\prime}$ and $j$ and all $i^{\prime\prime}\neq i$. Recall also that $R^{B}_{i}$ is assumed to be conditionally independent of $R^{A}_{i^{\prime}}$, $R^{C}_{j}$, $Y_{i^{\prime}}$ and $R^{B}_{i^{\prime\prime}}$ given $\mathcal{F}_{J}$ for all $i$, $i^{\prime}$ and $j$ and all $i^{\prime\prime}\neq i$. Because the choice of folds does not depend on $Y$ or $R^{B}$ values, these conditional independences also hold conditional on $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$.
All this implies that the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{k}\}$ on fold $k$ are conditionally independent of the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{-k}\}$ on the remaining folds given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$. Hence, since $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ are calculated using only the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{-k}\}$, the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{k}\}$ are conditionally independent of $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$. In Appendix A1 we show that this ensures that the empirical process and remainder terms are $o_{p}(M^{-1/2})$, as required, provided that Condition C1 holds.
5 Sampling of clusters without replacement with varying cluster sampling probabilities
Now consider the more general case where there are a small number, $L$, of distinct values of the cluster sampling probability $\pi^{C}_{j}$, which we denote as $\pi^{C(1)},\ldots,\pi^{C(L)}$. Simple random sampling of clusters without replacement is the special case of this with $L=1$. Let $M^{(l)}=M^{(l)}(R^{C}_{1},\ldots,R^{C}_{J})=\sum_{j=1}^{J}R^{C}_{j}\;I(\pi^{C}_{j}=\pi^{C(l)})$ and $J^{(l)}=\sum_{j=1}^{J}I(\pi^{C}_{j}=\pi^{C(l)})$ denote, respectively, the number of sampled clusters and the total number of clusters with cluster sampling probability $\pi^{C(l)}$ ($l=1,\ldots,L$). Clearly, $\sum_{l=1}^{L}J^{(l)}=J$ and $\sum_{l=1}^{L}M^{(l)}=M$. Assume that the cluster sampling mechanism satisfies the following condition:
$P(R^{C}_{1}=r_{1},\ldots,R^{C}_{1}=r_{J}\mid\mathcal{F}_{J})=P(R^{C}_{1}=r_{1}^{\prime},\ldots,R^{C}_{1}=r_{J}^{\prime}\mid\mathcal{F}_{J})$
for all values $(r_{1},\ldots,r_{J})$ and $(r_{1}^{\prime},\ldots,r_{J}^{\prime})$ of $(R^{C}_{1},\ldots,R^{C}_{J})$ such that
$M^{(l)}(r_{1},\ldots,r_{J})=M^{(l)}(r_{1}^{\prime},\ldots,r_{J}^{\prime})\hskip 22.76228pt\forall\;l=1,\ldots,L.$
This condition is satisfied by, for example, conditional Poisson sampling, Sampford sampling, Pareto sampling and randomised systematic sampling[19, 1].
For each of $l=1,\ldots,L$, randomly partition the $M^{(l)}$ sampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ evenly into $K$ folds, and randomly partition the $J^{(l)}-M^{(l)}$ unsampled clusters evenly into the same $K$ folds. For each $l$, this partition begins by randomly choosing a set of $K\times\lfloor M^{(l)}/K\rfloor$ sampled clusters and a set of $K\times\lfloor(J^{(l)}-M^{(l)})/K\rfloor$ unsampled clusters ($\lfloor x\rfloor$ denotes the integer part of $x$) and evenly partitioning each of these two sets between the folds. This leaves fewer than $K$ sampled clusters and fewer than $K$ unsampled clusters so far unassigned to folds. Each of the unassigned sampled clusters is randomly assigned to the $K$ folds in such a way that no fold receives more than one of these clusters, and the same is done with the unassigned unsampled clusters.
[…]
As in Section 4, $\hat{m}_{k}$ is calculated using the $Y$ values of all individuals with $R^{B}=1$ not in fold $k$, and $\hat{\pi}^{B}_{k}$ is calculated using $R^{B}$ values of all the individuals not in fold $k$. However, unlike in Section 4, this calculation of $\hat{\pi}^{B}_{k}$ uses the $R^{A}$ values of only individuals in a subset of the clusters not in fold $k$. We shall refer to this subset as the ‘active’ subset. The reason for using only this subset is to ensure that information about the $R^{A}$ values of individuals in fold $k$ provided by the value of $\hat{\pi}^{B}_{k}$ vanishes asymptotically. We now describe the procedure for choosing the active subset, first for the case where $L=1$ (simple random sampling without replacement) and then for the case where $L>1$.
If $L=1$, then at most $\lceil M/K\rceil$ of the $M$ sampled clusters are in fold $k$ (here, $\lceil x\rceil$ denotes the smallest integer that is greater than or equal to $x$). Hence, at least $M-\lceil M/K\rceil$ sampled clusters must be in the remaining folds. The active subset is chosen by randomly subsampling $M-\lceil M/K\rceil$ of the $M-M_{k}$ sampled clusters not in fold $k$. To compensate for the subsampling, we multiply the $\pi^{A}$ values of individuals not in fold $k$ by $(M-\lceil M/K\rceil)/(M-M/K)$ when calculating $\hat{\pi}^{B}_{k}$.
Now consider the case where $L>1$. Choose a number $0<\delta<1$ and define
$C^{(l)}=\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor\wedge(M^{(l)}-M^{(l)}_{k})\hskip 28.45274pt(l=1,\ldots,L)$ (5)
($a\wedge b$ denotes the minimum of $a$ and $b$). Choose the active subset by, for each $l=1,\ldots,L$, randomly subsampling $C^{(l)}$ of the $M^{(l)}-M^{(l)}_{k}$ sampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ not in fold $k$ To compensate for the subsampling, multiply the $\pi^{A}$ values of individuals with $\pi^{C}=\pi^{C(l)}$ not in fold $k$ by
$\frac{\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor}{\pi^{C(l)}(J^{(l)}-J^{(l)}_{k})}$ (6)
when calculating $\hat{\pi}^{B}_{k}$. The motivation for calculating $\hat{\pi}^{B}_{k}$ using the $R^{A}$ values only of individuals in the active clusters is that when $M$ is large, $C^{(l)}$ is very likely to equal $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor$. If $C^{(l)}$ were guaranteed to equal $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor$, then knowing $\hat{\pi}^{B}_{k}$ would tell us nothing about the $(R^{A},R^{B},Y)$ values of individuals in fold $k$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$.
Since, when $M$ is large, $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor\approx(M^{(l)}-M^{(l)}_{k})(1-\delta)$, the price we pay to achieve this noninformativeness of $\hat{\pi}^{B}_{k}$ when $M$ is large is the discarding of about $100\delta$% of the individuals with $R^{A}_{i}=1$ when calculating $\hat{\pi}^{B}_{k}$. This motivates us to choose $\delta$ close to zero, e.g. $\delta=0.01$.
In Appendix A2 we prove that when the folds and active subset are chosen in the way described above, the empirical process and remainder terms in the von Mises expansion (equation (3)) are $o_{p}(M^{-1/2})$, as required, provided that Condition C1 is satisfied.
In practice, the number of distinct cluster sampling probabilities will often not be small. In this situation, we propose dividing the $J$ clusters into a small number $L$ of sets according to the ranks of their sampling probabilities $\pi^{C}_{j}$. For example, one might divide them by quartile of $\pi^{C}_{j}$ into $L=4$ sets. When subsampling the active subset from the $l$th set of clusters and compensating for this subsampling, replace $\pi^{C(l)}$ in expressions (5) and (6) by the mean $\pi^{C}_{j}$ value of clusters in set $l$. This is what we do in the simulation study of Section 9. Provided that the values of $\pi^{C}_{j}$ do not vary greatly within each of the $L$ sets, the results in Appendix A2 should be approximately valid.
6 Ratio estimator
In the context where parametric nuisance models are used, CLW also proposed the estimator
[…]
$\hat{n}^{A}$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\frac{R^{A}_{i}}{\pi^{A}_{i}}$
$\hat{n}^{B}$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\frac{R^{B}_{i}}{\hat{\pi}^{B}(X_{i})}.$
This estimator $\hat{\theta}_{{\rm CLW}2}$ may be more efficient than $\hat{\theta}_{1}$, for the same reason that the Hajek estimator may be more efficient than the Horwitz-Thompson estimator. In CLW’s simulation study with parametric nuisance models, $\hat{\theta}_{{\rm CLW}2}$ did indeed have smaller variance than $\hat{\theta}_{1}$.
Here we propose instead the ratio estimator
$\hat{\theta}_{2}=\hat{\theta}_{2}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})=\frac{n}{\hat{n}^{A}}\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})=\frac{1}{\hat{n}^{A}}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\left[\frac{R^{A}_{i}}{\pi^{A}_{i}}\hat{m}_{k}(X_{i})+\frac{R^{B}_{i}}{\hat{\pi}^{B}_{k}(X_{i})}\{Y_{i}-\hat{m}_{k}(X_{i})\}\right].$
In Appendix A7, we show that if Condition C1 is satisfied,
$\sqrt{M}\{\hat{\theta}_{2}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}\}$ $=$ $\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}\left\{m_{0}(X_{i})-\frac{1}{n}\sum_{i^{\prime}=1}^{n}m_{0}(X_{i^{\prime}})\right\}$ (7)
$+\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{B}_{i}}{\pi^{B}_{0}(X_{i})}\{Y_{i}-m_{0}(X_{i})\}$
$+\sqrt{M}\frac{1}{n}\sum_{i=1}^{n}\{m_{0}(X_{i})-Y_{i}\}+o_{p}(1),$
and $\hat{\theta}_{2}=\hat{\theta}_{{\rm CLW}2}+o_{p}(M^{-1/2})$, i.e. $\hat{\theta}_{2}$ is asymptotically equivalent to $\hat{\theta}_{{\rm CLW}2}$.
The mean-zero term $\sqrt{M}n^{-1}\sum_{i=1}^{n}\{m_{0}(X_{i})-Y_{i}\}$ in expression (7) is $O_{p}(1)$ and does not vanish asymptotically. However, when the population is large compared to Samples A and B, the variance of this term is small compared to the variance of the other terms in expression (7). We see that if this term is ignored, equation (7) is the same as equation (4) but with $m_{0}(X_{i})$ in the first term replaced by $m_{0}(X_{i})-n^{-1}\sum_{i^{\prime}=1}^{n}m_{0}(X_{i^{\prime}})$.
Ignoring the term $\sqrt{M}n^{-1}\sum_{i=1}^{n}\{m_{0}(X_{i})-Y_{i}\}$ in equation (7), the asymptotic variance of $\hat{\theta}_{{\rm CLW}2}-\bar{Y}$ (and hence of $\hat{\theta}_{2}-\bar{Y}$) is given by CLW, along with an estimator of this variance. This is described in more detail by Seaman et al. (2025)[25] (specifically, the formulae for $\mu_{\rm DR2}$ when both nuisance models are correctly specified).
7 Targeted maximum likelihood
A popular alternative to estimator $\hat{\theta}_{\rm iid}$ for iid data is the targeted maximum likelihood estimator (TMLE). The TMLE method involves modifying initial estimate $\hat{m}_{\rm iid}$ of $E(Y\mid X)$ to $\hat{m}^{*}_{\rm iid}$ in such a way that the simple regression imputation estimator $n^{-1}\sum_{i=1}^{n}\hat{m}_{\rm iid}^{*}(X_{i})$ of $E(Y)$ can be used. We now describe an analogous TMLE method for our survey sample situation.
Define $\hat{m}_{k}(x;\epsilon_{k})$ as
$\hat{m}_{k}(x;\epsilon_{k})=\hat{m}_{k}(x)+\epsilon_{k}\;\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}$ (8)
for some $\epsilon_{k}$. Alternatively, if $Y$ is bounded by zero and one, i.e. $P(0<Y<1)=1$, then $\hat{m}_{k}(x;\epsilon_{k})$ can be defined by equation (8) or equation (9):
$\hat{m}_{k}(x;\epsilon_{k})=\frac{\exp\{\mbox{logit }\hat{m}_{k}(x)+\epsilon_{k}/\hat{\pi}^{B}_{k}(X_{i})\}}{1+\exp\{\mbox{logit }\hat{m}_{k}(x)+\epsilon_{k}/\hat{\pi}^{B}_{k}(X_{i})\}}.$ (9)
Define $\hat{m}^{*}_{k}(x)=\hat{m}_{k}(x;\hat{\epsilon}_{k})$, where $\hat{\epsilon}_{k}$ is the solution to the estimating equation
$\sum_{i\in\mathcal{S}_{k}}\frac{R^{B}_{i}}{\hat{\pi}^{B}_{k}(X_{i})}\{Y_{i}-\hat{m}_{k}(X_{i};\hat{\epsilon}_{k})\}=0.$ (10)
Note that if $\hat{m}_{k}(x;\epsilon_{k})$ is defined by equation (8), $\hat{\epsilon}_{k}$ can be calculated by fitting a linear regression model with no intercept and only the covariate $1/\hat{\pi}^{B}$ to the outcome $Y-\hat{m}_{k}(X)$ in individuals with $R^{B}=1$ in fold $k$. If $\hat{m}_{k}(x;\epsilon_{k})$ is defined by equation (9), $\hat{\epsilon}_{k}$ can be calculated by fitting a logistic regression model with offset $\mbox{logit }\hat{m}_{k}(X)$, no intercept and only the covariate $1/\hat{\pi}^{B}$ to the outcome $Y$ in those same individuals.
It follows from equation (10) that
$\hat{\theta}_{\rm TMLE1}=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U_{i}(\hat{\pi}^{B}_{k},\hat{m}^{*}_{k})=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}\;\hat{m}^{*}_{k}(X_{i}).$
This mass imputation estimator $\hat{\theta}_{\rm TMLE1}$ is our TMLE estimator. In Appendix A8 we show that $\hat{\theta}_{\rm TMLE1}$ is asymptotically equivalent to $\hat{\theta}_{1}$, i.e. $\hat{\theta}_{\rm TMLE1}=\hat{\theta}_{1}+o_{p}(M^{-1/2})$, provided that Condition C1 holds.
We also define the corresponding TMLE estimator $\hat{\theta}_{\rm TMLE2}=\hat{\theta}_{\rm TMLE1}\times n/\hat{n}^{A}$, which has the same asymptotic distribution as $\hat{\theta}_{2}$, i.e. $\hat{\theta}_{\rm TMLE2}=\hat{\theta}_{2}+o_{p}(M^{-1/2})$.
8 Calculating $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$
So far, we have been agnostic about how to calculate $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$, except to specify that they use, respectively, $(X,R^{A},R^{B})$ values and $(X,Y)$ values of individuals not in fold $k$. Estimating $m_{0}(X)=E(Y\mid X)$ is relatively straightforward, because $Y_{1},\ldots,Y_{n}$ are assumed to be independent given $X_{1},\ldots,X_{n}$. This is a standard problem for which many data-adaptive methods could be used.
Estimating $\pi^{B}_{0}(X)=P(R^{B}=1\mid X)$ is more complicated, because, although $R^{B}_{1},\ldots,R^{B}_{n}$ are assumed independent given $X_{1},\ldots,X_{n}$, we only observe $X$ for individuals with $R^{A}=1$ or $R^{B}=1$. CLW, working with parametric nuisance models, discuss pseudo-maximum likelihood, approximate pseudo-maximum likelihood, calibration and the method of Kim and Haziza (2014)[21]. The log pseudo-likelihood is defined as
$\sum_{i=1}^{n}R^{B}_{i}\;\log\pi^{B}(X_{i})-\left(\frac{R^{A}_{i}}{\pi^{A}}-R^{B}_{i}\right)\log\{1-\pi^{B}(X_{i})\}.$ (11)
This is the log likelihood for a binomial regression model of $R^{B}$ on $X$ in the population of $n$ individuals, using the individuals in Sample A weighted by their inverse sampling probabilities to represent the population, so that the ‘failures’ or ‘non-events’ are represented by the weighted Sample A minus the individuals in Sample B. Ferri-Garcia et al. (2022)[15] (see also [10]) use the pseudo-likelihood together with a classification and regression tree to estimate $\pi^{B}_{0}(X)$.
The approximate log pseudo-likelihood is the same as expression (11) but omitting the $R^{B}_{i}\{1-\pi^{B}(X_{i})\}$ term. When Sample B is small compared to the population, this term will be negligible compared to the other terms in expression (11). The approximate pseudo-likelihood has the advantage that it can be used quite generally with data-adaptive methods that predict a binary outcome and that allow for weights. We acknowledge CLW’s criticism of the approximate pseudo-likelihood in situations where Sample B is not a small fraction of the population, but we focus in Section 9 on scenarios where Sample B is small compared to the population, a situation which would be quite common in practice.
9 Simulation study
We simulated a finite population with $J=1000$ clusters. Each cluster $j$ contains $H_{jq}$ households of $q$ individuals ($q=1,2,3$), where $H_{jq}$ is negatively binomially distributed with mean 100 and variance 400. Hence, the expected number of individuals in a cluster is 600, and the expected population size is 600,000. For each individual $i$ in cluster $j$, continuous variables $X_{1i}$ and $X_{2i}$ and binary variables $X_{3i}$ and $X_{4i}$ were generated independently, with expectations that depend on the total number of households in the cluster $j$, the size $q$ of the household to which individual $i$ belongs, and a cluster-level random effect. This fixed population was used for all the simulations.
For each simulated dataset, $Y_{1},\ldots,Y_{n}$ were generated independently from a normal distribution with mean that is a function of $X_{i}=(X_{1i},X_{2i},X_{3i},X_{4i})$ and with variance 1. Sample A was drawn by sampling $M$ clusters with replacement from the $J$ clusters using Sampford sampling, with the probability of sampling cluster $j$ being proportional to the total number of households in cluster $j$. From each of these $M$ sampled clusters, $n_{\rm house}$ households were sampled using simple random sampling without replacement. Then one individual was sampled at random from each of these $n_{\rm house}$ sampled households. Thus, Sample A contained $M\times n_{\rm house}$ individuals. The probability of sampling individual $i$ for Sample B was a function of $X_{i}$.
We considered six scenarios. In Scenarios 1 and 2, $m_{0}(X)$ is linear in $X$ and $\pi^{B}_{0}(X)$ follows a logistic regression model with only main effects for $X_{1}$, $X_{2}$, $X_{3}$ and $X_{4}$. Specifically, $m_{0}(X)=X_{1}+X_{2}+2X_{3}+X_{4}$ and $\mbox{logit }\pi^{B}_{0}(X)=\alpha_{\rm int}+0.5X_{1}+X_{2}+0.5X_{3}+X_{4}$. Both these scenarios use $M=150$ and $n_{\rm house}=20$. They differ in the value of $\alpha_{\rm int}$, and hence the expected size of Sample B. In Scenario 1, $\alpha_{\rm int}=-6.2$ and the expected size of Sample B is 7000; in Scenario 2, $\alpha_{\rm int}=-7.5$ and the expected size of Sample B is 2000. In Scenarios 3–6, $m_{0}(X)$ is a non-linear function of $X$ and the logistic regression model for $\pi^{B}_{0}(X)$ includes an interaction and a quadratic term. Specifically, $m_{0}(X)=0.5X_{1}+0.5X_{2}+2X_{3}+X_{4}+2X_{1}X_{3}+X_{2}^{2}$ and $\mbox{logit }\pi^{B}_{0}(X)=-6.4+0.25X_{1}+0.5X_{2}+0.5X_{3}+X_{4}+X_{1}X_{3}+0.5X_{2}^{2}$. The expected size of Sample B is 7500. These four scenarios differ in the values of $M$ and $n_{\rm house}$. Specifically, Scenarios 3–6 use, respectively, $(M,n_{\rm house})=(150,20)$, $(50,20)$, $(150,5)$ and $(50,5)$.
The following estimators were applied to each of 100 simulated datasets for each of the six scenarios.
HT
: Horwitz-Thompson estimator that only uses Sample A
Hajek
: Hajek estimator that only uses Sample A
Naive
: Simple unweighted mean of $Y$ in Sample A.
DR1
: $\hat{\theta}_{1}$ using parametric nuisance models
DR2clw
: $\hat{\theta}_{{\rm CLW}2}$ using parametric nuisance models
DR2
: $\hat{\theta}_{2}$ using parametric nuisance models
TMLE1
: $\hat{\theta}_{\rm TMLE1}$ using parametric nuisance models
TMLE2
: $\hat{\theta}_{\rm TMLE2}$ using parametric nuisance models
KH
: $\hat{\theta}_{1}$ with parametric nuisance models whose parameters are estimated using method of Kim and Haziza[6, 21]
DR1.hal5
: $\hat{\theta}_{1}$ using Highly Adaptive LASSO (HAL) to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
DR2.hal5
: $\hat{\theta}_{2}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
TMLE1.hal5
: $\hat{\theta}_{\rm TMLE1}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
TMLE2.hal5
: $\hat{\theta}_{\rm TMLE2}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
Of these, DR1, DR2clw and KH were described by CLW. For the methods using HAL, we set $\delta=0.01$ and used $L=4$ sets of clusters based on quartiles of the distribution of $\pi^{C}_{j}$ (as described at the end of Section 5).
In addition, we calculated the estimators using generalised boosting models (GBM), in place of HAL, to estimate the nuisance models. These will be denoted ‘DR1.gbm5’ and so on. We also applied the estimators that used HAL or GBM without cross-fitting. These will be denoted ‘DR1.hal1’, ‘DR1.gbm1’, and so on.
Results are shown in Tables 1, 2, 3, 4, 5 and 6 for Scenarios 1–6, respectively.
In Scenarios 1 and 2, the methods using parametric nuisance models, viz. DR1, DR2clw, DR2, TMLE1, TMLE2 and KH, yield approximately unbiased point estimates and standard error estimates, and the coverage of 95% confidence intervals is correct. The estimators DR2clw and TMLE2 are more efficient than the DR1 and TMLE1 estimators, and DR2 is, as expected, as efficient as DR2clw. Using HAL to estimate the nuisance functions provides no benefit in these scenarios, where the parametric nuisance models are correctly specified, but also no loss of efficiency or coverage. Using GBM leads to some bias and a small deterioriation in coverage, but bias is still small and coverage not far from its nominal level.
In Scenarios 3–6, the methods using parametric nuisance models yield biased estimates and confidence intervals with poor coverage. When HAL is used, the bias is small and the coverage close to its nominal level. Results from GBM are similar, although the bias can be slightly greater than for HAL. Using cross-fitting with five folds appears to provide no performance improvement over not using cross-fitting in these scenarios but also no loss of efficiency.
bias empSE SEhat cover
HT 0.002 0.046 0.046 94
Haj 0.000 0.037 0.037 95
naive -0.288 0.035
DR1 0.001 0.044 0.044 94
DR2clw -0.000 0.035 0.035 95
DR2 -0.000 0.035 0.035 95
TMLE1 0.001 0.044 0.044 94
TMLE2 -0.000 0.035 0.035 95
KH 0.001 0.044 0.044 94
DR1.hal5 0.002 0.044 0.044 93
DR2.hal5 0.001 0.035 0.034 94
TMLE1.hal5 0.002 0.044 0.044 93
TMLE2.hal5 0.000 0.035 0.034 94
DR1.hal1 0.002 0.044 0.044 93
DR2.hal1 0.000 0.034 0.034 94
TMLE1.hal1 0.002 0.044 0.044 93
TMLE2.hal1 0.000 0.034 0.034 94
DR1.gbm5 0.012 0.044 0.043 92
DR2.gbm5 0.010 0.034 0.033 91
TMLE1.gbm5 0.009 0.044 0.043 93
TMLE2.gbm5 0.008 0.034 0.033 92
DR1.gbm1 0.010 0.044 0.043 92
DR2.gbm1 0.009 0.034 0.033 92
TMLE1.gbm1 0.007 0.044 0.043 92
TMLE2.gbm1 0.006 0.034 0.033 92
Table 1: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 1.
bias empSE SEhat cover
HT 0.002 0.046 0.046 94
Haj 0.000 0.037 0.037 95
naive -0.288 0.035
DR1 0.001 0.052 0.051 96
DR2clw -0.001 0.044 0.043 95
DR2 -0.001 0.044 0.043 95
TMLE1 0.001 0.052 0.051 96
TMLE2 -0.001 0.044 0.043 95
KH 0.001 0.052 0.051 96
DR1.hal5 0.002 0.052 0.049 95
DR2.hal5 0.001 0.044 0.041 94
TMLE1.hal5 0.002 0.052 0.050 95
TMLE2.hal5 0.000 0.044 0.042 94
DR1.hal1 0.002 0.052 0.050 95
DR2.hal1 0.001 0.044 0.041 94
TMLE1.hal1 0.001 0.052 0.050 95
TMLE2.hal1 0.000 0.044 0.042 94
DR1.gbm5 0.021 0.058 0.053 91
DR2.gbm5 0.019 0.051 0.045 90
TMLE1.gbm5 0.020 0.052 0.054 92
TMLE2.gbm5 0.019 0.044 0.045 91
DR1.gbm1 0.017 0.052 0.048 92
DR2.gbm1 0.016 0.043 0.039 90
TMLE1.gbm1 0.013 0.052 0.049 94
TMLE2.gbm1 0.011 0.044 0.040 92
Table 2: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 2.
bias empSE SEhat cover
HT 0.001 0.062 0.063 94
Haj -0.001 0.053 0.054 95
naive -0.399 0.047
DR1 0.222 0.068 0.078 16
DR2clw 0.184 0.053 0.069 18
DR2 0.220 0.058 0.069 8
TMLE1 0.125 0.064 0.070 58
TMLE2 0.124 0.053 0.060 46
KH 0.089 0.062 0.063 72
DR1.hal5 0.002 0.063 0.062 94
DR2.hal5 0.001 0.053 0.053 94
TMLE1.hal5 0.002 0.063 0.066 96
TMLE2.hal5 0.001 0.053 0.053 94
DR1.hal1 0.002 0.062 0.062 94
DR2.hal1 0.001 0.053 0.053 94
TMLE1.hal1 0.002 0.062 0.067 96
TMLE2.hal1 0.001 0.053 0.053 94
DR1.gbm5 0.020 0.063 0.062 94
DR2.gbm5 0.018 0.053 0.052 92
TMLE1.gbm5 0.016 0.063 0.066 96
TMLE2.gbm5 0.015 0.053 0.052 92
DR1.gbm1 0.019 0.063 0.062 94
DR2.gbm1 0.018 0.053 0.052 92
TMLE1.gbm1 0.016 0.063 0.066 95
TMLE2.gbm1 0.014 0.053 0.052 92
Table 3: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 3.
bias empSE SEhat cover
HT -0.002 0.109 0.110 94
Haj -0.001 0.091 0.094 95
naive -0.399 0.081
DR1 0.238 0.119 0.126 56
DR2clw 0.190 0.084 0.110 60
DR2 0.239 0.099 0.110 42
TMLE1 0.118 0.110 0.116 86
TMLE2 0.119 0.090 0.098 82
KH 0.085 0.108 0.106 87
DR1.hal5 0.000 0.109 0.106 93
DR2.hal5 0.001 0.090 0.089 95
TMLE1.hal5 -0.000 0.108 0.113 94
TMLE2.hal5 0.001 0.090 0.089 95
DR1.hal1 0.000 0.108 0.106 93
DR2.hal1 0.001 0.090 0.089 95
TMLE1.hal1 -0.000 0.108 0.113 94
TMLE2.hal1 0.001 0.090 0.089 95
DR1.gbm5 0.024 0.109 0.105 93
DR2.gbm5 0.025 0.090 0.088 93
TMLE1.gbm5 0.019 0.109 0.113 95
TMLE2.gbm5 0.020 0.090 0.088 94
DR1.gbm1 0.020 0.109 0.105 94
DR2.gbm1 0.021 0.090 0.088 94
TMLE1.gbm1 0.016 0.109 0.113 96
TMLE2.gbm1 0.016 0.090 0.088 94
Table 4: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 4.
bias empSE SEhat cover
HT -0.011 0.111 0.112 94
Haj -0.013 0.092 0.093 94
naive -0.406 0.081
DR1 0.239 0.129 0.123 52
DR2clw 0.186 0.090 0.104 57
DR2 0.237 0.111 0.104 38
TMLE1 0.117 0.108 0.115 85
TMLE2 0.115 0.087 0.094 79
KH 0.082 0.105 0.107 90
DR1.hal5 -0.003 0.105 0.106 95
DR2.hal5 -0.005 0.086 0.086 94
TMLE1.hal5 -0.003 0.105 0.113 96
TMLE2.hal5 -0.005 0.086 0.086 94
DR1.hal1 -0.003 0.105 0.106 95
DR2.hal1 -0.005 0.086 0.086 94
TMLE1.hal1 -0.003 0.105 0.113 96
TMLE2.hal1 -0.005 0.086 0.086 94
DR1.gbm5 0.022 0.106 0.106 96
DR2.gbm5 0.021 0.087 0.085 94
TMLE1.gbm5 0.018 0.106 0.112 97
TMLE2.gbm5 0.016 0.087 0.086 94
DR1.gbm1 0.019 0.105 0.106 96
DR2.gbm1 0.017 0.086 0.085 94
TMLE1.gbm1 0.015 0.106 0.112 97
TMLE2.gbm1 0.013 0.087 0.086 94
Table 5: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 5.
bias empSE SEhat cover
HT 0.008 0.195 0.196 95
Haj 0.004 0.164 0.164 95
naive -0.394 0.142
DR1 0.329 0.262 0.212 68
DR2clw 0.224 0.147 0.178 80
DR2 0.325 0.246 0.178 58
TMLE1 0.123 0.186 0.202 94
TMLE2 0.119 0.149 0.165 91
KH 0.095 0.178 0.184 94
DR1.hal5 0.010 0.180 0.184 95
DR2.hal5 0.007 0.146 0.149 96
TMLE1.hal5 0.011 0.180 0.196 97
TMLE2.hal5 0.007 0.145 0.149 96
DR1.hal1 0.011 0.180 0.184 95
DR2.hal1 0.007 0.145 0.149 96
TMLE1.hal1 0.010 0.180 0.196 97
TMLE2.hal1 0.006 0.146 0.149 96
DR1.gbm5 0.035 0.180 0.184 94
DR2.gbm5 0.031 0.146 0.149 95
TMLE1.gbm5 0.036 0.181 0.196 96
TMLE2.gbm5 0.032 0.146 0.149 94
DR1.gbm1 0.032 0.181 0.184 95
DR2.gbm1 0.029 0.146 0.149 95
TMLE1.gbm1 0.032 0.181 0.196 96
TMLE2.gbm1 0.028 0.146 0.149 95
Table 6: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 6.
10 Discussion
We have described a debiased machine learning approach for estimating the mean of $Y$ from a nonprobability sample when auxiliary information on $X$ are available from a probability sample. This method yields an asymptotically normally distributed estimator with a standard error that can estimated as easily as can the standard error of a Horwitz-Thompson estimator. To prove this, we have relied on cross-fitting and have (unless cluster sampling probabilities are all equal) discarded some information when estimating the nuisance function $\hat{\pi}^{B}$. For the closely related problem of estimating $E(Y)$ for an infinite population using iid data, an alternative to cross-fitting is to rely on the Donsker condition [20]. This limits the choice of data-adaptive estimators of $P(R=1\mid X)$ and $E(Y\mid X)$ to those that satisfy the Donsker condition. Nevertheless, it would be interesting to establish whether the Donsker condition would suffice in the setting we study in this article.
Some researchers have previously used DR estimators with data-adaptive estimation of nuisance functions, e.g. [3] and [11]. However, they did not establish the validity of this approach. Indeed [11] comment (page 276) that “there are hardly any results on the theoretical properties of variance estimators when more complex machine learning models and techniques are used.” Moreover, these researchers used data-adaptive estimators that do not necessarily satisfy the Donsker condition and without using cross-fitting, which is potentially problematic even in the simpler situation of estimating $E(Y)$ for an infinite population using iid data.
To establish asymptotic properties of our estimators, we have assumed that the finite population is generated from a particular superpopulation model, so that it grows in a particular way. It would be interesting to investigate whether asymptotic properties can be established when the finite population is allowed to grow in another way. We also used subsampling of an ‘active subset’ of clusters when calculating $\hat{\pi}^{B}$. This should have no effect on asymptotic efficiency, and in our simulation study we observed no loss of efficiency from doing it. However, if loss of efficiency in finite samples is a concern, it could be mitigated by using repeated cross-fitting [8]. As with other debiased machine learning estimators using cross-fitting, repeated cross-fitting would also have the advantage of reducing the finite-sample Monte Carlo variation induced by the randomness of the process of partitioning units (in our case, clusters) among folds, although at the cost of increasing computation time.
The approach we have presented is for probability samples that do not use stratified sampling. It is straightforward in principle to generalise it to deal with stratified sampling. Now the assignment of clusters to the $K$ folds should be done within each stratum. This should be fairly unproblematic when the probability sample uses simple random sampling of clusters without replacement within each stratum or when the number of clusters sampled from each stratum is large. When, however, the cluster sampling probabilities within strata vary and the number of sampled clusters within strata is small, there may be two problems. First, the subsampling of active subsets of clusters could lead to substantial loss of information for estimating $\hat{\pi}^{B}$ in this situation. Second, even with the subsampling of an active subset, the asymptotic property that $\hat{\pi}^{B}_{k}$ is uninformative about $R^{A}$ values in fold $k$ may be a poor approximation in this situation. The scenario of sampling few clusters per stratum is a problem worthy of further research.
Finally, several researchers have considered the problem of efficiently combining an IPW or DR estimator of the mean of $Y$ with a Horwitz-Thompson or Hajek estimator of this same mean based on Sample A alone. Gao and Yang [18] consider the DR estimator $\hat{\theta}_{1}$ and use parametric nuisance models. Seaman et al. (2025) discuss this approach and extend it to $\hat{\theta}_{{\rm CLW}2}$ and IPW estimators, again using parametric nuisance models. Rueda et al. (2023)[13] used an IPW estimator with random forest or XGboost to estimate the nuisance function. Since, when Condition C1 is satisfied, the debiased machine learning estimators of $\bar{Y}$ that we have presented here have the same asymptotic distribution as the corresponding DR estimators of $\bar{Y}$ that use two correctly specified parametric nuisance models, the combining methods discussed by Seaman et al. (2025) are equally appropriate.
Acknowledgements
This work is supported by the National Survey of Sexual Attitudes and Lifestyles (Natsal) grant from the Wellcome Trust (212931/Z/18/Z), with contributions from the Economic and Social Research Council (ESRC) and the National Institute for Health Research (NIHR). I thank Tommy Nyberg, Anne Presanis and Andrew Copas for helpful discussions, and Stijn Vansteelandt and Andrea Rotnitzky for very early helpful discussions and for teaching me a great deal about semiparametric theory. None of these people are at all responsible for any errors in this manuscript.
For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
ORCID
Shaun R. Seaman https://orcid.org/0000-0003-3726-5937
Appendices for ‘Debiased machine learning for combining probability and non-probability samples’
Shaun R. Seaman
A1 Proof that empirical process and remainder terms are $o_{p}(M^{-1/2})$ for simple random sampling of clusters without replacement
A1.1 Empirical process term
We now introduce the notation
$\mathcal{G}_{J}$ $=$ $(R^{A}_{11},\ldots,R^{A}_{1N_{1}},\ldots,R^{A}_{J1},\ldots,R^{A}_{JN_{J}},R^{B}_{11},\ldots,R^{B}_{1N_{1}},\ldots,R^{B}_{J1},\ldots,R^{B}_{JN_{J}},$
$\hskip 5.69046ptY_{11},\ldots,Y_{1N_{1}},\ldots,Y_{J1},\ldots,Y_{JN_{J}}).$
We shall sometimes use this to help the reader to know what with respect to which random variables an expectation or variance is being taken.
Consider the empirical process term for the $k$th fold ($k=1,\ldots,K$), i.e.
$\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\left[\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\right$.
$\hskip 51.21504pt-\left.E_{\mathcal{G}_{J}}\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}\right].$ (A1)
By the Law of Total Probability, we have
[…]
$\hskip 62.59596pt-\left.\left.E_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]\right|>\Delta\mid\mathcal{F}_{J}\right)$
[…]
$\hskip 11.38092pt\left.\left.-E_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]\bigg{|}>\Delta\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)\mid\mathcal{F}_{J}\right\}.$
By Chebeshev’s Inequality, we have, for any $\Delta>0$ and $\mathcal{F}_{J}$, $\mathcal{S}_{k}$, $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$,
[…]
$\hskip 28.45274pt-\left.E_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\}\right]\bigg{|}>\Delta\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 28.45274pt\leq\frac{{\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]}{\Delta^{2}}$ (A3)
Now,
$U(\hat{\pi}^{B}_{k},\hat{m}_{k})-U(\pi^{B}_{0},m_{0})$ $=$ $R^{B}Y\left\{\frac{1}{\hat{\pi}^{B}_{k}(X)}-\frac{1}{\pi^{B}_{0}(X)}\right\}+\frac{R^{A}}{\pi^{A}}\{\hat{m}_{k}(X)-m_{0}(X)\}$
$-R^{B}\left\{\frac{\hat{m}_{k}(X)}{\hat{\pi}^{B}_{k}(X)}-\frac{m_{0}(X)}{\pi^{B}_{0}(X)}\right\}$
As we argued in Section 4, $(R^{B}_{i},Y_{i})$ is conditionally independent of $R^{B}_{i^{\prime}}$, $Y_{i^{\prime}}$, $R^{A}_{i^{\prime\prime}}$ and $R^{C}_{j}$ for all $i$, all $i^{\prime}\neq i$, and all $i^{\prime\prime}$ and $j$ given $\mathcal{F}_{J}$, $\mathcal{S}_{k}$, $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$. Hence,
${\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$
$={\rm Var}_{\mathcal{G}_{J}}\left(\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\left[R^{B}_{i}Y_{i}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{1}{\pi^{B}_{0}(X_{i})}\right\}-R^{B}_{i}\left\{\frac{\hat{m}_{k}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{m_{0}(X_{i})}{\pi^{B}_{0}(X_{i})}\right\}\right]\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 11.38092pt+{\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right].$ (A4)
If we can show that the conditional expectation given $\mathcal{F}_{J}$ of each the two variances in expression (A4) is $o_{p}(M^{-1})$, then it will follow from equation (A2) that the empirical process term for fold $k$ (i.e. expression (A1)) is $o_{p}(M^{-1/2})$. We shall now look at these two variances in turn.
Consider the first variance in expression (A4). As we argued in Section 4, $R^{B}_{1},Y_{i},\ldots,R^{B}_{n},Y_{n}$ are conditionally independent of $\mathcal{S}_{k}$, $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$, and of each other, given $\mathcal{F}_{J}$. Hence,
${\rm Var}_{\mathcal{G}_{J}}\left(\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\left[R^{B}_{i}Y_{i}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{1}{\pi^{B}_{0}(X_{i})}\right\}-R^{B}_{i}\left\{\frac{\hat{m}_{k}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{m_{0}(X_{i})}{\pi^{B}_{0}(X_{i})}\right\}\right]\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 8.5359pt=\frac{1}{n_{k}^{2}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{1}{\pi^{B}_{0}(X_{i})}\right\}^{2}{\rm Var}(R^{B}_{i}Y_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})$
$\hskip 14.22636pt+\frac{1}{n_{k}^{2}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{\hat{m}_{k}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{m_{0}(X_{i})}{\pi^{B}_{0}(X_{i})}\right\}^{2}{\rm Var}(R^{B}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})$
$\hskip 14.22636pt-\frac{2}{n_{k}^{2}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{1}{\pi^{B}_{0}(X_{i})}\right\}\left\{\frac{\hat{m}_{k}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{m_{0}(X_{i})}{\pi^{B}_{0}(X_{i})}\right\}{\rm Cov}(R^{B}_{i}Y_{i},R^{B}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})$
$\hskip 8.5359pt=\frac{1}{n_{k}^{2}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{1}{\pi^{B}_{0}(X_{i})}\right\}^{2}{\rm Var}(R^{B}_{i}Y_{i}\mid\mathcal{F}_{J})$
$\hskip 14.22636pt+\frac{1}{n_{k}^{2}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{\hat{m}_{k}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{m_{0}(X_{i})}{\pi^{B}_{0}(X_{i})}\right\}^{2}{\rm Var}(R^{B}_{i}\mid\mathcal{F}_{J})$
$\hskip 14.22636pt-\frac{2}{n_{k}^{2}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{1}{\pi^{B}_{0}(X_{i})}\right\}\left\{\frac{\hat{m}_{k}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{m_{0}(X_{i})}{\pi^{B}_{0}(X_{i})}\right\}{\rm Cov}(R^{B}_{i}Y_{i},R^{B}_{i}\mid\mathcal{F}_{J})$
$=\frac{1}{n_{k}^{2}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{1}{\pi^{B}_{0}(X_{i})}\right\}^{2}[\pi^{B}_{0}(X_{i})\;\{1-\pi^{B}_{0}(X_{i})\}\;\{m_{0}(X_{i})\}^{2}+\pi^{B}_{0}(X_{i})\;{\rm Var}(Y_{i}\mid X_{i})]$
$\hskip 14.22636pt+\frac{1}{n_{k}^{2}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{\hat{m}_{k}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{m_{0}(X_{i})}{\pi^{B}_{0}(X_{i})}\right\}^{2}\pi^{B}_{0}(X_{i})\;\{1-\pi^{B}_{0}(X_{i})\}$
$\hskip 14.22636pt-\frac{2}{n_{k}^{2}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{1}{\pi^{B}_{0}(X_{i})}\right\}\left\{\frac{\hat{m}_{k}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{m_{0}(X_{i})}{\pi^{B}_{0}(X_{i})}\right\}\pi^{B}_{0}(X_{i})\;\{1-\pi^{B}_{0}(X_{i})\}\;m_{0}(X_{i}).$
We see that for any fixed functions $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$, expression (A5) is $O_{p}(M^{-1})$.
Now consider the second variance in expression (A4). As we argued in Section 4, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ corresponds to simple random sampling of $M/K$ clusters without replacement from the $J/K$ clusters in $\mathcal{S}_{k}$ followed by the original second-stage sampling of individuals within clusters. Hence, for any fixed function $\hat{m}_{k}$,
${\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$ (A6)
is the variance of the Horwitz-Thompson estimator of the population mean of $\hat{m}_{k}(X)-m_{0}(X)$ in individuals in $\mathcal{S}_{k}$ given $\mathcal{F}_{J}$ when the clusters in $\mathcal{S}_{k}$ are selected by simple random sampling without replacement and individuals within clusters are sampled according to the original second-stage sampling procedure. Subject to some regularity conditions, this variance is $O_{p}(M^{-1})$ (Proposition 2 of Chauvet and Vallee, 2020 [4]).
It follows that, if $\hat{\pi}^{B}_{k}\xrightarrow{p}\pi^{B}_{0}$ and $\hat{m}_{k}\xrightarrow{p}m_{0}$ as $M\rightarrow\infty$, then expressions (A5) and (A6), and hence (A4), are $o_{p}(M^{-1})$, as required.
A1.2 Remainder term
We can write the remainder term for fold $k$ as
$\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}E_{\mathcal{G}_{J}}\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}$
$\hskip 14.22636pt=\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}E\left[\frac{R^{B}_{i}}{\hat{\pi}^{B}_{k}(X_{i})}Y_{i}+\left\{\frac{R^{A}_{i}}{\pi^{A}_{i}}-\frac{R^{B}_{i}}{\hat{\pi}^{B}_{k}(X_{i})}\right\}\hat{m}_{k}(X_{i})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$
$\hskip 28.45274pt-\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}E\left[\frac{R^{B}_{i}}{\pi^{B}_{0}(X_{i})}Y_{i}+\left\{\frac{R^{A}_{i}}{\pi^{A}_{i}}-\frac{R^{B}_{i}}{\pi^{B}_{0}(X_{i})}\right\}m_{0}(X_{i})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$
$\hskip 14.22636pt=\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{\pi^{B}_{0}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}m_{0}(X_{i})+\left\{\frac{E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})}{\pi^{A}_{i}}-\frac{\pi^{B}_{0}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}\right\}\hat{m}_{k}(X_{i})$
$\hskip 28.45274pt-\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})}{\pi^{A}_{i}}\;m_{0}(X_{i}).$ (A7)
As argued in Section 4, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ corresponds to simple random sampling of $M/K$ clusters without replacement from the $J/K$ clusters in $\mathcal{S}_{k}$ followed by the original second-stage sampling of individuals within clusters. Hence, $E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})=\pi^{A}_{i}$. Hence, expression (A7) reduces to
$\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{\pi^{B}_{0}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}m_{0}(X_{i})+\left\{1-\frac{\pi^{B}_{0}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}\right\}\hat{m}_{k}(X_{i})-\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}m_{0}(X_{i})$
$\hskip 14.22636pt=-\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{\pi^{B}_{0}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-1\right\}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}$ (A8)
$=-E\left[\left\{\frac{\pi^{B}_{0}(X)}{\hat{\pi}^{B}_{k}(X)}-1\right\}\{\hat{m}_{k}(X)-m_{0}(X)\}\right]$
$\hskip 14.22636pt-\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{\pi^{B}_{0}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-1\right\}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\right$.
$\hskip 34.14322pt-\left.E\left[\left\{\frac{\pi^{B}_{0}(X)}{\hat{\pi}^{B}_{k}(X)}-1\right\}\{\hat{m}_{k}(X)-m_{0}(X)\}\right]\right].$ (A9)
For fixed functions $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$, the term
$\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{\pi^{B}_{0}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-1\right\}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\right$.
$\hskip 14.22636pt-\left.E\left[\left\{\frac{\pi^{B}_{0}(X)}{\hat{\pi}^{B}_{k}(X)}-1\right\}\{\hat{m}_{k}(X)-m_{0}(X)\}\right]\right]$
in expression (A9) is $O_{p}(M^{-1/2})$. Hence, if $\hat{\pi}^{B}_{k}\xrightarrow{p}\pi^{B}_{0}$ or $\hat{m}_{k}\xrightarrow{p}m_{0}$, then it is $o_{p}(M^{1/2})$. This leaves only the term
$E\left[\left\{\frac{\pi^{B}_{0}(X)}{\hat{\pi}^{B}_{k}(X)}-1\right\}\{\hat{m}_{k}(X)-m_{0}(X)\}\right]$
in expression (A9). By the Cauchy-Schwarz Inequality, the absolute value of this is less than or equal to
$\sqrt{E\left\{\frac{\pi^{B}_{0}(X)}{\hat{\pi}^{B}_{k}(X)}-1\right\}^{2}\times E\{\hat{m}_{k}(X)-m_{0}(X)\}^{2}.}$
So, if Condition C1 holds, then expression (A8) is $o_{p}(M^{-1/2})$, as required.
A2 Proof that empirical process and remainder terms are $o_{p}(M^{-1/2})$ for random sampling of clusters without replacement with unequal probabilities
A2.1 Empirical process term
Consider the empirical process term for fold $k$. Equations (A2) and (A3) still apply when clusters are sampled without replacement with unequal probabilities. Consider the variance in expression (A3).
${\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$
$\hskip 5.69046pt=E\left({\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}\right]\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 8.5359pt+{\rm Var}\left(E_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}\right]\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right).$
If we can show that the conditional expectation given $\mathcal{F}_{J}$ of each of the two summands in expression (A10) is $o_{p}(M^{-1})$, then we shall have shown that expression (A1), the empirical process term for fold $k$, is $o_{p}(M^{-1/2})$. We shall look at the two terms in expression (A10) in turn.
Consider the first of the two terms in expression (A10) and look at the inner variance term, i.e.
${\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}\right$].
Analogously to equation (A4), we have
${\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}\right$]
$={\rm Var}_{\mathcal{G}_{J}}\left(\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\left[R^{B}_{i}Y_{i}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{1}{\pi^{B}_{0}(X_{i})}\right\}-R^{B}_{i}\left\{\frac{\hat{m}_{k}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{m_{0}(X_{i})}{\pi^{B}_{0}(X_{i})}\right\}\right]\right$.
[…]
$\hskip 11.38092pt+{\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}\right$]
$={\rm Var}_{\mathcal{G}_{J}}\left(\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\left[R^{B}_{i}Y_{i}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{1}{\pi^{B}_{0}(X_{i})}\right\}-R^{B}_{i}\left\{\frac{\hat{m}_{k}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}-\frac{m_{0}(X_{i})}{\pi^{B}_{0}(X_{i})}\right\}\right]\right$.
$\hskip 71.13188pt\left.\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 11.38092pt+{\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}\right$] (A11)
The first of the two variance terms in expression (A11) was shown earlier (see expression (A5)) to be $O_{p}(M^{-1})$ and moreover to be $o_{p}(M^{-1})$ if $\hat{\pi}^{B}_{k}\xrightarrow{p}\pi^{B}_{0}$ and $\hat{m}_{k}\xrightarrow{p}m_{0}$.
Consider the second of the two terms in expression (A11). We have
${\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}\right$]
$={\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\frac{R^{A}_{i}}{\pi^{A}_{i}M^{(l)}_{k}/(J^{(l)}_{k}\pi^{C(l)})}\;\frac{M^{(l)}_{k}}{J^{(l)}_{k}\pi^{C(l)}}\right$.
$\hskip 56.9055pt\times\left.\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}\right$]. (A12)
By a very similar argument to that used in Section 4, we see that, because of the way that fold $S_{k}$ has been chosen, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}$ corresponds to stratified simple random sampling of $M^{(l)}_{k}$ clusters without replacement from the stratum of $J^{(l)}_{k}$ clusters with the same value of $\pi^{C(l)}$ followed by the original second-stage sampling of individuals within clusters. The probability of sampling individual $i$ is the original probability $\pi^{A}_{i}$ multiplied by $M^{(l)}_{k}/(J^{(l)}_{k}\pi^{C(l)})$, because the fraction of clusters sampled from this stratum is $M^{(l)}_{k}/J^{(l)}_{k}$, rather than $\pi^{C(l)}$. Hence, for any fixed function $\hat{m}_{k}$, expression (A12) is the variance of the Horwitz-Thompson estimator of the population mean in the $k$th fold of $\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\{M^{(l)}_{k}/(J^{(l)}_{k}\pi^{C(l)})\}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}$ when clusters are sampled in this stratified way. Subject to some regularity conditions, this variance is $O_{p}(M^{-1})$ (Proposition 2 of Chauvet and Vallee, 2020 [4]). Consequently, if $\hat{m}_{k}\xrightarrow{p}m_{0}$ as $M\rightarrow\infty$, then this variance is $o_{p}(M^{-1})$. Hence, the conditional expectation given $\mathcal{F}_{J},S_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ of this variance is also $o_{p}(M^{-1})$. Therefore, the first of the two terms in expression (A10) is $o_{p}(M^{-1})$, as required.
Now consider the second of the two terms in expression (A10). We have
$E_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}\right$]
$\hskip 5.69046pt=\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k})}{\pi^{A}_{i}}-\frac{\pi^{B}_{0}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}\right\}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}$
$\hskip 5.69046pt=\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\left\{\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\frac{M^{(l)}_{k}}{\pi^{C(l)}J^{(l)}_{k}}-\frac{\pi^{B}_{0}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}\right\}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}$
So,
${\rm Var}\left(E_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}\right]\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$={\rm Var}\left[\sum_{l=1}^{L}\frac{M^{(l)}_{k}}{\pi^{C(l)}J^{(l)}_{k}}\;\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$
$={\rm Var}\left[\sum_{l=1}^{L}\frac{M^{(l)}}{\pi^{C(l)}J^{(l)}}\;\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\right$.
$\hskip 39.83368pt+\left.\sum_{l=1}^{L}\frac{1}{\pi^{C(l)}}\left(\frac{M^{(l)}_{k}}{J^{(l)}_{k}}-\frac{M^{(l)}}{J^{(l)}}\right)\right$.
$\hskip 71.13188pt\times\left.\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$
$={\rm Var}\left[\sum_{l=1}^{L}\frac{M^{(l)}_{k}}{\pi^{C(l)}J^{(l)}_{k}}\;\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$
$\hskip 17.07182pt+D$ (A13)
where
$D$ $=$ ${\rm Var}\left[\sum_{l=1}^{L}\frac{1}{\pi^{C(l)}}\left(\frac{M^{(l)}_{k}}{J^{(l)}_{k}}-\frac{M^{(l)}}{J^{(l)}}\right)\right$.
$\hskip 42.67912pt\times\left.\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$
+ $+\;2\;{\rm Cov}\left[\sum_{l=1}^{L}\frac{1}{\pi^{C(l)}}\left(\frac{M^{(l)}_{k}}{J^{(l)}_{k}}-\frac{M^{(l)}}{J^{(l)}}\right)\right.\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\;,$
$\hskip 28.45274pt\left.\sum_{l=1}^{L}\frac{1}{\pi^{C(l)}}\frac{M^{(l)}}{J^{(l)}}\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$
This $D$ term is a term accounts for the possible difference between $M^{(l)}/J^{(l)}$, the overall proportion of sampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$, and $M^{(l)}_{k}/J^{(l)}_{k}$, the proportion in fold $k$. Given the way that clusters are assigned to folds, this difference between proportions will become smaller as $M\rightarrow\infty$. In Appendix A3 we show that
$M^{(l)}_{k}/J^{(l)}_{k}-M^{(l)}/J^{(l)}=O_{p}(M^{-1})$ (A14)
and in Appendix A4, we show that this implies that $D$ is $o_{p}(M^{-1})$.
Now consider the first term in expression (A13). Recall that we aim to show that the conditional expectation given $\mathcal{F}_{J}$ of this is $o_{p}(M^{-1})$. We have
${\rm Var}\left[\sum_{l=1}^{L}\frac{M^{(l)}}{\pi^{C(l)}J^{(l)}}\;\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$
$=\sum_{l=1}^{L}\sum_{l^{\prime}=l}^{L}{\rm Cov}\left[\frac{M^{(l)}}{\pi^{C(l)}J^{(l)}}\;\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\;,\right$.
$\hskip 56.9055pt\left.\frac{M^{(l^{\prime})}}{\pi^{C(l^{\prime})}J^{(l^{\prime})}}\;\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l^{\prime})})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right].$
Using the Covariance Inequality and Cauchy-Schwartz Inequality (see Appendix A5), we obtain
[…]
$\hskip 14.22636pt\left.\left.\left.\frac{M^{(l^{\prime})}}{\pi^{C(l^{\prime})}J^{(l^{\prime})}}\;\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l^{\prime})})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]\mid\mathcal{F}_{J}\right)\right|$
[…]
$\hskip 99.58464pt\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\bigg{]}\mid\mathcal{F}_{J}\bigg{)}\bigg{\}}^{1/2}$
[…]
$\hskip 99.58464pt\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\bigg{]}\mid\mathcal{F}_{J}\bigg{)}\bigg{\}}^{1/2}.$
Hence, if we can show that
[…]
$\hskip 99.58464pt\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\bigg{]}\mid\mathcal{F}_{J}\bigg{)}\;=\;O_{p}(M^{-1}),$ (A16)
then the conditional expectation given $\mathcal{F}_{J}$ of expression (A15) is also $O_{p}(M^{-1})$. From that, it follows that if $\hat{m}_{k}\xrightarrow{p}m_{0}$, then this conditional expectation will be $o_{p}(M^{-1})$, as required.
So, it only remains to show that equation (A16) holds. Now,
[…]
$\hskip 99.58464pt\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\bigg{]}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k}\bigg{)}$
$=\left(\frac{J}{J^{(l)}}\right)^{2}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\right]^{2}$
$\hskip 14.22636pt\times E\left\{{\rm Var}\left(\frac{M^{(l)}}{\pi^{C(l)}J}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k}\right\}$
Also,
$E\left\{{\rm Var}\left(\frac{M^{(l)}}{\pi^{C(l)}J}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k}\right\}$
$={\rm Var}\left(\frac{M^{(l)}}{\pi^{C(l)}J}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k}\right)$
$\hskip 14.22636pt-{\rm Var}\left\{E\left(\frac{M^{(l)}}{\pi^{C(l)}J}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k}\right\}$
$={\rm Var}\left(\frac{M^{(l)}}{\pi^{C(l)}J}\mid\mathcal{F}_{J}\right)$
$\hskip 14.22636pt-{\rm Var}\left\{E\left(\frac{M^{(l)}}{\pi^{C(l)}J}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k}\right\}$
$={\rm Var}\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J}\right)$
$\hskip 14.22636pt-{\rm Var}\left\{E\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k}\right\}$
The first term in expression (A17) is $O_{p}(M^{-1})$. Consider the expectation inside the second term of expression (A17).
$E\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$=E\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J}\right)$
$\hskip 14.22636pt+E\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 14.22636pt-E\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J}\right)$
$=\frac{J^{(l)}}{J}+E\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 14.22636pt-E\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J}\right)$
Hence, the second term in expression (A17) is $O_{p}(M^{-1})$ if
$E\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 14.22636pt-E\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J}\right)=O_{p}(M^{-1/2}).$ (A18)
In Appendix A6, we show that equation (A18) holds. Thus, equation (A16) does hold, as required.
A2.2 Remainder term
Consider the remainder term for fold $k$, i.e.
$\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}E_{\mathcal{G}_{J}}\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}.$
Equation (A7) still applies when clusters are sampled without replacement with unequal probabilities. As before, because of the way that the folds have been chosen, and using equation (A14), we have
$\frac{E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})}{\pi^{A}_{i}}$
$\hskip 8.5359pt=\frac{E\{E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\}}{\pi^{A}_{i}}$
$\hskip 8.5359pt=E\left\{\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\frac{M^{(l)}_{k}}{\pi^{C(l)}J^{(l)}_{k}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}$
$\hskip 8.5359pt=\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\frac{1}{\pi^{C(l)}}\;E\left(\frac{M^{(l)}_{k}}{J^{(l)}_{k}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 8.5359pt=\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\frac{1}{\pi^{C(l)}}\;E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 22.76228pt+\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\frac{1}{\pi^{C(l)}}\;E\left(\frac{M^{(l)}_{k}}{J^{(l)}_{k}}-\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 8.5359pt=\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\frac{1}{\pi^{C(l)}}\;E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)+O_{p}(M^{-1})$ (A19)
$\hskip 8.5359pt=\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\frac{J}{J^{(l)}}\;E\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)+O_{p}(M^{-1}).$
Note that equation (A19) follows from equation (A14). It follows from equation (A18) that equation (A20) reduces to
$\frac{E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})}{\pi^{A}_{i}}$
$\hskip 8.5359pt=\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\frac{J}{J^{(l)}}\;E\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J}\right)+O_{p}(M^{-1/2})+O_{p}(M^{-1})$
$\hskip 8.5359pt=\sum_{l=1}^{L}I(\pi^{C}_{i}=\pi^{C(l)})\times 1+O_{p}(M^{-1/2})$
$\hskip 8.5359pt=1+O_{p}(M^{-1/2}).$
Plugging this into equation (A7), we obtain
$\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}E_{\mathcal{G}_{J}}\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}$
$\hskip 14.22636pt=\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{\pi^{B}_{0}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}m_{0}(X_{i})+\left\{1-\frac{\pi^{B}_{0}(X_{i})}{\hat{\pi}^{B}_{k}(X_{i})}\right\}\hat{m}_{k}(X_{i})-\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}m_{0}(X_{i})$
$\hskip 28.45274pt+\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}O_{p}(M^{-1/2})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}.$ (A21)
The final term in expression (A21) is $O_{p}(M^{-1/2})$ for fixed $\hat{m}_{k}$. If $\hat{m}_{k}\xrightarrow{P}m_{0}$, then this term becomes $o_{p}(M^{-1/2})$, and so can be ignored. The rest of expression (A21) is the same as expression (A8). Hence, we are back to the situation we have where sampling of clusters is by simple random sampling without replacement. Thus, under the conditions given there, the remainder term is $o_{p}(M^{-1/2})$, as required.
A3 Proof that $M^{(l)}_{k}/J^{(l)}_{k}-M^{(l)}/J^{(l)}=O_{p}(M^{-1})$
By the way that the folds are chosen, we have the following bounds on $M^{(l)}_{k}/J^{(l)}_{k}-M^{(l)}/J^{(l)}$.
$\frac{\lfloor M^{(l)}/K\rfloor}{\lfloor J^{(l)}/K\rfloor+1}-\frac{\lfloor M^{(l)}/K\rfloor+1-K^{-1}}{\lfloor J^{(l)}/K\rfloor}\leq\frac{M^{(l)}_{k}}{J^{(l)}_{k}}-\frac{M^{(l)}}{J^{(l)}}\leq\frac{\lfloor M^{(l)}/K\rfloor+1}{\lfloor J^{(l)}/K\rfloor}-\frac{\lfloor M^{(l)}/K\rfloor}{\lfloor J^{(l)}/K\rfloor+1-K^{-1}}.$
Consider the upper bound.
$\frac{\lfloor M^{(l)}/K\rfloor+1}{\lfloor J^{(l)}/K\rfloor}-\frac{\lfloor M^{(l)}/K\rfloor}{\lfloor J^{(l)}/K\rfloor+1-K^{-1}}$ $=$ $\frac{\lfloor J^{(l)}/K\rfloor+\lfloor M^{(l)}/K\rfloor(1-K^{-1})+1-K^{-1}}{\left(\lfloor J^{(l)}/K\rfloor\right)^{2}+\lfloor J^{(l)}/K\rfloor(1-K^{-1})}$
$\rightarrow$ $\frac{\lfloor J^{(l)}/K\rfloor+\lfloor M^{(l)}/K\rfloor(1-K^{-1})}{\left(\lfloor J^{(l)}/K\rfloor\right)^{2}}\hskip 14.22636pt\mbox{as }M\rightarrow\infty$
$=$ $\frac{1}{\lfloor J^{(l)}/K\rfloor}+\frac{\lfloor M^{(l)}/K\rfloor(1-K^{-1})}{\left(\lfloor J^{(l)}/K\rfloor\right)^{2}}$
$=$ $O_{p}(M^{-1}).$
Similarly, the lower bound is $O_{p}(M^{-1})$. Hence,
$M^{(l)}_{k}/J^{(l)}_{k}-M^{(l)}/J^{(l)}=O_{p}(M^{-1}).$
A4 Proof that $D=o_{p}(M^{-1})$
We can write
$D$ $=$ $\sum_{l=1}^{L}\sum_{l^{\prime}=1}^{L}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\right]$
$\hskip 42.67912pt\times\;\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l^{\prime})})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\right]\times\frac{1}{\pi^{C(l)}\pi^{C(l^{\prime})}}$
$\hskip 42.67912pt\times\;{\rm Cov}\left(\frac{M^{(l)}_{k}}{J^{(l)}_{k}}-\frac{M^{(l)}}{J^{(l)}}\;,\;\frac{M^{(l)}_{k}}{J^{(l)}_{k}}-\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
+ $+\;2\;\sum_{l=1}^{L}\sum_{l^{\prime}=1}^{L}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\right]$
$\hskip 56.9055pt\times\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l^{\prime})})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\right]$
$\hskip 56.9055pt\times\frac{1}{\pi^{C(l)}\pi^{C(l^{\prime})}}\times{\rm Cov}\left(\frac{M^{(l)}_{k}}{J^{(l)}_{k}}-\frac{M^{(l)}}{J^{(l)}}\;,\;\frac{M^{(l^{\prime})}}{J^{(l^{\prime})}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right).$
Consider the first of the two summands in expression (LABEL:eq:defineD). Using expression (A14), we have
${\rm Cov}_{M^{(.)}}\left(\frac{M^{(l)}_{k}}{J^{(l)}_{k}}-\frac{M^{(l)}}{J^{(l)}}\;,\;\frac{M^{(l^{\prime})}_{k}}{J^{(l^{\prime})}_{k}}-\frac{M^{(l^{\prime})}}{J^{(l^{\prime})}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)$
$\hskip 8.5359pt={\rm Cov}_{M^{(.)}}\left\{O_{p}(M^{-1}),O_{p}(M^{-1})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}$
$\hskip 8.5359pt=O_{p}(M^{-2})$
$\hskip 8.5359pt=o_{p}(M^{-1}).$
Consider the second of the two summands in expression (LABEL:eq:defineD). We have
$\left|{\rm Cov}\left(\frac{M^{(l)}_{k}}{J^{(l)}_{k}}-\frac{M^{(l)}}{J^{(l)}}\;,\;\frac{M^{(l^{\prime})}}{J^{(l^{\prime})}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)\right|$
$\leq\sqrt{{\rm Var}\left(\frac{M^{(l)}_{k}}{J^{(l)}_{k}}-\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)}\times\sqrt{{\rm Var}\left(\frac{M^{(l^{\prime})}}{J^{(l^{\prime})}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)}$
$=O_{p}(M^{-1})\times\sqrt{{\rm Var}\left(\frac{M^{(l^{\prime})}}{J^{(l^{\prime})}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)}.$ (A23)
Note that equation (A23) follows from expression (A14).
The variance of a proportion cannot be greater than 1, and hence expression (A23) is $O_{p}(M^{-1})$. Hence, if $\hat{m}_{k}\xrightarrow{p}m_{0}$, the second of the two summands in expression (LABEL:eq:defineD) is $o_{p}(M^{-1})$, as required.
Since both of the summands in expression (LABEL:eq:defineD) are $o_{p}(M^{-1})$, $D$ is itself $o_{p}(M^{-1})$.
A5 Application of covariance inequality and Cauchy-Schwartz inequality
For any three random variables $A$, $B$ and $C$, the covariance inequality is that $\{{\rm Cov}(A,B\mid C)\}^{2}\leq{\rm Var}(A\mid C)\times{\rm Var}(B\mid C)$, and hence
$|{\rm Cov}(A,B\mid C)|\leq\sqrt{{\rm Var}(A\mid C)}\times\sqrt{{\rm Var}(B\mid C)}.$
Also, the Cauchy-Schwarz inequality says that $E_{C}\{\sqrt{{\rm Var}(A\mid C)}\times\sqrt{{\rm Var}(B\mid C)}\}]^{2}\leq E_{C}[\{\sqrt{{\rm Var}(A\mid C)}\}^{2}]\times E_{C}[\{\sqrt{{\rm Var}(B\mid C)}\}^{2}]=E_{C}\{{\rm Var}(A\mid C)\}\times E_{C}\{{\rm Var}(B\mid C)\}$, and hence
$E_{C}\{\sqrt{{\rm Var}(A\mid C)}\times\sqrt{{\rm Var}(B\mid C)}\}]\leq\sqrt{E_{C}\{{\rm Var}(A\mid C)\}\times E_{C}\{{\rm Var}(B\mid C)\}}.$
Also,
$|E_{C}\{{\rm Cov}(A,B\mid C)\}|\leq E_{C}\{|{\rm Cov}(A,B\mid C)|\}.$
Putting this together, we obtain
$|E_{C}\{{\rm Cov}(A,B\mid C)\}|\leq\sqrt{E_{C}\{{\rm Var}(A\mid C)\}\times E_{C}\{{\rm Var}(B\mid C)\}}.$ (A24)
Interpret the random variables $A$, $B$ and $C$ as follows.
$A$ $=$ $\frac{M^{(l)}}{\pi^{C(l)}J^{(l)}}\;\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l)})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}$
$B$ $=$ $\frac{M^{(l^{\prime})}}{\pi^{C(l^{\prime})}J^{(l^{\prime})}}\;\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l^{\prime})})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}$
$C$ $=$ $\{\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\}$
Using inequality (A24) and conditioning throughout on $\mathcal{F}_{J}$, we obtain
[…]
$\hskip 28.45274pt\left.\left.\left.\frac{M^{(l^{\prime})}}{\pi^{C(l^{\prime})}J^{(l^{\prime})}}\;\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}I(\pi^{C}_{i}=\pi^{C(l^{\prime})})\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]\mid\mathcal{F}_{J}\right)\right|$
[…]
$\hskip 99.58464pt\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\bigg{]}\mid\mathcal{F}_{J}\bigg{)}\bigg{\}}^{1/2}$
[…]
$\hskip 99.58464pt\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\bigg{]}\mid\mathcal{F}_{J}\bigg{)}\bigg{\}}^{1/2}$
A6 Proof of equation (A18)
By multiplying both sides of equation (A18) by $\pi^{C(l)}J/J^{(l)}$ and noting that $E(M^{(l)}/J^{(l)}\mid\mathcal{F}_{J})=\pi^{C(l)}$, we see that equation (A18) is true if and only if
$\sqrt{M}\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)-\pi^{C(l)}\right\}=O_{p}(1).$ (A25)
We now prove that this holds.
Define $C^{*}=1$ if $C^{(l)}<\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor$ for any $l=1,\ldots,L$, and $C^{*}=0$ otherwise. As explained informally in Section 5, it is very likely that $C^{*}=0$ when $M$ is large.
Now,
$E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},C^{*}=0\right)$
$\hskip 8.5359pt=\frac{1}{J^{(l)}}\sum_{m^{(l)}=0}^{M}m^{(l)}\;P(M^{(l)}=m^{(l)}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},C^{*}=0)$
$\hskip 8.5359pt=\frac{1}{J^{(l)}}\sum_{m^{(l)}=0}^{M}m^{(l)}\;P(M^{(l)}=m^{(l)}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0)$
$\hskip 17.07182pt\times\frac{p(\hat{\pi}^{B}_{k},\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0,M^{(l)}=m^{(l)})}{p(\hat{\pi}^{B}_{k},\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0)}$
$\hskip 8.5359pt=\frac{1}{J^{(l)}}\sum_{m^{(l)}=0}^{M}m^{(l)}\;P(M^{(l)}=m^{(l)}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0)$
$\hskip 28.45274pt\times\frac{p(\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0,M^{(l)}=m^{(l)})}{p(\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0)}\times\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0,M^{(l)}=m^{(l)})}{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0)}$
$\hskip 8.5359pt=\frac{1}{J^{(l)}}\sum_{m^{(l)}=0}^{M}m^{(l)}\;P(M^{(l)}=m^{(l)}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0)$
$\hskip 28.45274pt\times\frac{p(\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k})}{p(\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k})}\times\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0,M^{(l)}=m^{(l)})}{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0)}$
$\hskip 8.5359pt=\frac{1}{J^{(l)}}\sum_{m^{(l)}=0}^{M}m^{(l)}\;P(M^{(l)}=m^{(l)}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0)$
$\hskip 28.45274pt\times\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0,M^{(l)}=m^{(l)})}{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0)}$ (A26)
We shall now show that the ratio of densities in expression (A26) equals 1.
When $C^{*}=0$, define $R^{C*}_{-k}$ to be the set of indices of the $\sum_{l=1}^{L}\lfloor\pi^{C(l)}(1-\epsilon)(J^{(l)}-J^{(l)}_{k})\rfloor$ clusters in $\mathcal{S}_{-k}$ that are used to estimate $\hat{\pi}^{B}_{k}$. Let $\mathcal{R}^{C*}_{-k}$ denote the set of possible values of $R^{C*}_{-k}$ that are compatible with $\mathcal{F}_{J}$, $\mathcal{S}_{k}$ and $C^{*}=0$. The size of this set is
$\prod_{l=1}^{L}\left(\begin{array}[]{c}J^{(l)}-J^{(l)}_{k}\\ \lfloor\pi^{C(l)}(1-\epsilon)(J^{(l)}-J^{(l)}_{k})\rfloor\end{array}\right)$,
i.e. the product over $l=1,\ldots,L$ of the number of ways of choosing $\lfloor\pi^{C(l)}(1-\epsilon)(J^{(l)}-J^{(l)}_{k})\rfloor$ elements from a set of $J^{(l)}-J^{(l)}_{k}$ elements.
Now,
$p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0,M^{(l)}=m^{(l)})$
$\hskip 8.5359pt=\sum_{r^{C*}_{-k}\in\mathcal{R}^{C*}_{-k}}p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0,M^{(l)}=m^{(l)},R^{C*}_{-k}=r^{C*}_{-k})$
$\hskip 56.9055pt\times P(R^{C*}_{-k}=r^{C*}_{-k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0,M^{(l)}=m^{(l)})$
$\hskip 8.5359pt=\sum_{r^{C*}_{-k}\in\mathcal{R}^{C*}_{-k}}p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},R^{C*}_{-k}=r^{C*}_{-k})$
$\hskip 56.9055pt\times P(R^{C*}_{-k}=r^{C*}_{-k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0,M^{(l)}=m^{(l)})$
$\hskip 8.5359pt=\sum_{r^{C*}_{-k}\in\mathcal{R}^{C*}_{-k}}p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},R^{C*}_{-k}=r^{C*}_{-k})\bigg{/}\prod_{l=1}^{L}\left(\begin{array}[]{c}J^{(l)}-J^{(l)}_{k}\\ \lfloor\pi^{C(l)}(1-\epsilon)(J^{(l)}-J^{(l)}_{k})\rfloor\end{array}\right)$. (A29)
Since expression (A30) is not a function of $m^{(l)}$, it follows that
$p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0,M^{(l)}=m^{(l)})=p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=0).$
Therefore, returning to equation (A26), we have
$E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},C^{*}=0\right)$
$\hskip 8.5359pt=\frac{1}{J^{(l)}}\sum_{m^{(l)}=0}^{M}m^{(l)}\;P(M^{(l)}=m^{(l)}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0)$
$\hskip 8.5359pt=E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},,C^{*}=0\right).$ (A31)
Using equation (A31), we have
$\sqrt{M}\;\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)-\pi^{C(l)}\right\}$
$\hskip 14.22636pt=\sqrt{M}\;P(C^{*}=1\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})\;\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},C^{*}=1\right)-\pi^{C(l)}\right\}$
$\hskip 28.45274pt+\sqrt{M}\;P(C^{*}=0\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})\;\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},C^{*}=0\right)-\pi^{C(l)}\right\}$
$\hskip 14.22636pt=\sqrt{M}\;P(C^{*}=1\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})\;\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},C^{*}=1\right)-\pi^{C(l)}\right\}$
$\hskip 28.45274pt+\sqrt{M}\;P(C^{*}=0\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})\;\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0\right)-\pi^{C(l)}\right\}.$
Now,
$P(C^{*}=1\mid\mathcal{F}_{J},\mathcal{S}_{k})$ $\leq$ $\sum_{l=1}^{L}P\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})<\frac{J^{(l)}}{J}(1-\epsilon)\mid\mathcal{F}_{J},\mathcal{S}_{k}\right)$
$=$ $\sum_{l=1}^{L}P\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})-\frac{J^{(l)}}{J}<-\frac{J^{(l)}}{J}\epsilon\mid\mathcal{F}_{J},\mathcal{S}_{k}\right)$
$\leq$ $\sum_{l=1}^{L}P\left(\left|\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})-\frac{J^{(l)}}{J}\right|>\frac{J^{(l)}}{J}\epsilon\mid\mathcal{F}_{J},\mathcal{S}_{k}\right)$
$\leq$ $\sum_{l=1}^{L}\frac{{\rm Var}\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J},\mathcal{S}_{k}\right)}{\left(\frac{J^{(l)}\epsilon}{J}\right)^{2}}$
$=$ $\sum_{l=1}^{L}\left(\frac{J}{J^{(l)}\epsilon}\right)^{2}{\rm Var}\left(\frac{1}{J}\sum_{j=1}^{J}\frac{R^{C}_{j}}{\pi^{C}_{j}}I(\pi^{C}_{j}=\pi^{C(l)})\mid\mathcal{F}_{J}\right)$
$=$ $O_{p}(M^{-1}).$ (A34)
Note that line (A6) uses Chebeshev’s Inequality.
It follows from equation (A34) that
$0$ $=$ $\sqrt{M}\;\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k}\right)-\pi^{C(l)}\right\}$
$=$ $\sqrt{M}\;P(C^{*}=0\mid\mathcal{F}_{J},\mathcal{S}_{k})\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0\right)-\pi^{C(l)}\right\}$
$+\sqrt{M}\;P(C^{*}=1\mid\mathcal{F}_{J},\mathcal{S}_{k})\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=1\right)-\pi^{C(l)}\right\}$
$=$ $\sqrt{M}\;\{1-O_{p}(M^{-1})\}\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0\right)-\pi^{C(l)}\right\}$
$+\sqrt{M}\;O_{p}(M^{-1})\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=1\right)-\pi^{C(l)}\right\}$
$=$ $\sqrt{M}\;\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0\right)-\pi^{C(l)}\right\}+o_{p}(1).$
Hence
$\sqrt{M}\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0\right)-\pi^{C(l)}\right\}=o_{p}(1).$ (A35)
Now, returning to equation (A32) and using equation (A35), we have
$\sqrt{M}\;\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right)-\pi^{C(l)}\right\}$
$\hskip 14.22636pt=\sqrt{M}\;P(C^{*}=1\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})\;\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},C^{*}=1\right)-\pi^{C(l)}\right\}$
$\hskip 28.45274pt+\sqrt{M}\;P(C^{*}=0\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})\;\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=0\right)-\pi^{C(l)}\right\}$
$\hskip 14.22636pt=\sqrt{M}\;P(C^{*}=1\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})\;\left\{E\left(\frac{M^{(l)}}{J^{(l)}}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},C^{*}=1\right)-\pi^{C(l)}\right\}$
$\hskip 28.45274pt+o_{p}(1).$ (A36)
Since the expectation of a proportion is bounded by zero and one, it follows from equation (A36) that equation (A25) holds if
$\sqrt{M}\;P(C^{*}=1\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})=O_{p}(1).$ (A37)
We shall now prove that equation (A37) does hold.
Making use of equation (A34), we have
$\sqrt{M}\;P(C^{*}=1\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})$
$\hskip 8.5359pt=\sqrt{M}\;P(C^{*}=1\mid\mathcal{F}_{J},\mathcal{S}_{k})\;\frac{p(\hat{\pi}^{B}_{k},\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=1)}{p(\hat{\pi}^{B}_{k},\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k})}$
$\hskip 8.5359pt=O_{p}(M^{-1/2})\;\frac{p(\hat{\pi}^{B}_{k},\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=1)}{p(\hat{\pi}^{B}_{k},\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k})}$
$\hskip 8.5359pt=O_{p}(M^{-1/2})\;\frac{p(\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},C^{*}=1)}{p(\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k})}\;\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=1)}{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k})}$
$\hskip 8.5359pt=O_{p}(M^{-1/2})\;\frac{p(\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k})}{p(\hat{m}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k})}\;\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=1)}{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k})}$ (A38)
$\hskip 8.5359pt=O_{p}(1)\;\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=1)}{\sqrt{M}\;p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k})}.$ (A39)
Note that line (A38) follows because $\hat{m}_{k}$ depends only on $\{(X_{i},Y_{i}):i\in\mathcal{S}_{-k}\mbox{ and }R^{B}_{i}=1\}$, which are conditionally independent of the $R^{C}_{j}$’s, and hence of $C^{*}$, given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$.
Hence, equation (A37) holds if
$\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=1)}{\sqrt{M}\;p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k})}=O_{p}(1).$
Now,
$E\left[\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=1)}{\sqrt{M}\;p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k})}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k}\right]$
$=\int\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=1)}{\sqrt{M}\;p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k})}\;p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k})\;d\hat{\pi}^{B}_{k}$
$=\int\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=1)}{\sqrt{M}}\;d\hat{\pi}^{B}_{k}$
$=\frac{1}{\sqrt{M}}.$
Thus, using Markov’s inequality, we have that for any $c>0$,
$P\left[\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=1)}{\sqrt{M}\;p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k})}\geq c\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k}\right]$
$\leq\frac{1}{c}\;E\left[\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=1)}{\sqrt{M}\;p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k})}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k}\right]$
$=\frac{1}{c\;\sqrt{M}}.$ (A40)
Hence,
$\frac{p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k},C^{*}=1)}{\sqrt{M}\;p(\hat{\pi}^{B}_{k}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{m}_{k})}=O_{p}(M^{-1/2}).$
which is a stronger result than we needed. It now follows from equation (A39) that
$\sqrt{M}\;P(C^{*}=1\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})=O_{p}(M^{-1/2}).$
Therefore, equation (A25) holds. In fact, it also holds with $O_{p}(1)$ replaced by $O_{p}(M^{-1/2})$.
A7 Proof of equation (7) and $\hat{\theta}_{2}=\hat{\theta}_{{\rm CLW}2}+o_{p}(M^{-1/2})$
[…]
$\hat{\theta}_{2}(\underline{\pi}^{B},\underline{m},\tau)=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U^{\dagger}_{i}(\pi^{B}_{k},m_{k},\tau),$
and
$\hat{\tau}=\frac{1}{n}\sum_{i=1}^{n}\frac{R^{A}_{i}}{\pi^{A}_{i}},$
with
$U^{\dagger}(\pi^{B}_{k},m_{k},\tau)=\frac{1}{\tau}\left[\frac{R^{A}}{\pi^{A}}m_{k}(X)+\frac{R^{B}}{\pi^{B}_{k}(X)}\{Y-m_{k}(X)\}\right].$
Also define $\bar{m}_{0}=\frac{1}{n}\sum_{i=1}^{n}m_{0}(X_{i})$.
Now, by a Taylor-series expansion,
$\sqrt{M}\;\{\hat{\theta}_{2}(\underline{\hat{\pi}}^{B},\underline{\hat{m}},\hat{\tau})-\bar{Y}\}$
$\hskip 14.22636pt=\sqrt{M}\;\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U^{\dagger}_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k},1)+\sqrt{M}\;\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\left.\frac{\partial U^{\dagger}_{i}}{\partial\tau}(\hat{\pi}^{B}_{k},\hat{m}_{k},\tau)\right|_{\tau=1}\;(\hat{\tau}-1)$
$\hskip 25.6073pt-\sqrt{M}\;\bar{Y}+o_{p}(1)$
$\hskip 14.22636pt=\sqrt{M}\;\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U^{\dagger}_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k},1)-\sqrt{M}\;\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U^{\dagger}_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k},1)\times(\hat{\tau}-1)$
$\hskip 25.6073pt-\sqrt{M}\;\bar{Y}+o_{p}(1)$
$\hskip 14.22636pt=\sqrt{M}\;\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-\sqrt{M}\;\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})\times(\hat{\tau}-1)$
$\hskip 25.6073pt-\sqrt{M}\;\bar{Y}+o_{p}(1)$
$\hskip 14.22636pt=\sqrt{M}\;\left[\frac{1}{n}\sum_{i=1}^{n}U_{i}(\pi^{B}_{0},m_{0})+o_{p}(M^{-1/2})\right]$
$\hskip 25.6073pt-\sqrt{M}\;\left[\frac{1}{n}\sum_{i=1}^{n}U_{i}(\pi^{B}_{0},m_{0})+o_{p}(M^{-1/2})\right]\times(\hat{\tau}-1)$
$\hskip 25.6073pt-\sqrt{M}\;\bar{Y}+o_{p}(1)$ (A41)
$\hskip 14.22636pt=\sqrt{M}\;\frac{1}{n}\sum_{i=1}^{n}U_{i}(\pi^{B}_{0},m_{0})-\sqrt{M}\;\frac{1}{n}\sum_{i=1}^{n}U_{i}(\pi^{B}_{0},m_{0})\times(\hat{\tau}-1)-\sqrt{M}\;\bar{Y}+o_{p}(1)$
$\hskip 14.22636pt=\sqrt{M}\;\frac{1}{n}\sum_{i=1}^{n}\frac{R^{A}_{i}}{\pi^{A}_{i}}m_{0}(X_{i})+\sqrt{M}\;\frac{1}{n}\sum_{i=1}^{n}\frac{R^{B}_{i}}{\pi^{B}_{0}(X_{i})}\{Y_{i}-m_{0}(X_{i})\}$
$\hskip 25.6073pt-\sqrt{M}\;\{\bar{m}_{0}+o_{p}(1)\}\times(\hat{\tau}-1)-\sqrt{M}\;\bar{Y}+o_{p}(1)$
$\hskip 14.22636pt=\sqrt{M}\;\frac{1}{n}\sum_{i=1}^{n}\frac{R^{A}_{i}}{\pi^{A}_{i}}m_{0}(X_{i})+\sqrt{M}\;\frac{1}{n}\sum_{i=1}^{n}\frac{R^{B}_{i}}{\pi^{B}_{0}(X_{i})}\{Y_{i}-m_{0}(X_{i})\}$
$\hskip 25.6073pt-\{\bar{m}_{0}+o_{p}(1)\}\;\sqrt{M}\;\frac{1}{n}\sum_{i=1}^{n}\left(\frac{R^{A}_{i}}{\pi^{A}_{i}}-1\right)-\sqrt{M}\;\bar{Y}+o_{p}(1)$
$\hskip 14.22636pt=\sqrt{M}\;\frac{1}{n}\sum_{i=1}^{n}\frac{R^{A}_{i}}{\pi^{A}_{i}}\{m_{0}(X_{i})-\bar{m}_{0}\}+\sqrt{M}\;\frac{1}{n}\sum_{i=1}^{n}\frac{R^{B}_{i}}{\pi^{B}_{0}(X_{i})}\{Y_{i}-m_{0}(X_{i})\}$
$\hskip 25.6073pt+\sqrt{M}(\bar{m}_{0}-\bar{Y})+o_{p}(1)$ (A42)
This is the same as equation (7). Note that line (A41) uses the result that $\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})=\hat{\theta}_{1}(\pi^{B}_{0},m_{0})+o_{p}(M^{-1/2})$.
Seaman et al. (2025)[25] show that $\sqrt{M}\{\hat{\theta}_{{\rm CLW}2}-\bar{Y}\}$ also equals expression (A42). Hence, $\hat{\theta}_{2}=\hat{\theta}_{{\rm CLW}2}+o_{p}(M^{-1/2})$.
A8 Proof that $\hat{\theta}_{\rm TMLE1}=\hat{\theta}_{1}+o_{p}(M^{1/2})$
Using a Taylor series expansion, we obtain
$\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}U_{i}(\hat{\pi}^{B}_{k},\hat{m}^{*}_{k})$
$=\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})+\left.\frac{d}{d\epsilon_{k}}\left[\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}U_{i}\{\hat{\pi}^{B}_{k},\hat{m}_{k}(\epsilon_{k})\}\right]\right|_{\epsilon_{k}=0}\hat{\epsilon}_{k}+o_{p}(\hat{\epsilon}_{k}^{2}).$
Now, if $\hat{m}_{k}(x;\epsilon_{k})$ is defined by equation (8), then
$\left.\frac{d}{d\epsilon_{k}}\left[\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}U_{i}\{\hat{\pi}^{B}_{k},\hat{m}_{k}(\epsilon_{k})\}\right]\right|_{\epsilon_{k}=0}$ $=$ $\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{R^{A}_{i}}{\pi^{A}_{i}}-\frac{R^{B}_{i}}{\hat{\pi}^{B}(X_{i})}\right\}\frac{1}{\hat{\pi}^{B}(X_{i})}$ (A43)
$=$ $\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}\left\{1-\frac{R^{B}_{i}}{\hat{\pi}^{B}(X_{i})}\right\}\frac{1}{\hat{\pi}^{B}(X_{i})}$
$+\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}\left(\frac{R^{A}_{i}}{\pi^{A}_{i}}-1\right)\frac{1}{\hat{\pi}^{B}(X_{i})}$
Now, for any fixed $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$,
$\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}\left\{1-\frac{R^{B}_{i}}{\hat{\pi}^{B}(X_{i})}\right\}\frac{1}{\hat{\pi}^{B}(X_{i})}$
$\hskip 8.5359pt=\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}\left\{1-\frac{\pi^{B}(X_{i})}{\hat{\pi}^{B}(X_{i})}\right\}\frac{1}{\hat{\pi}^{B}(X_{i})}$
$\hskip 22.76228pt-\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}\{R^{B}_{i}-\pi^{B}(X_{i})\}\;\frac{1}{\hat{\pi}^{B}(X_{i})^{2}}$
$\hskip 8.5359pt=E_{X}\left[\left\{1-\frac{\pi^{B}(X)}{\hat{\pi}^{B}(X)}\right\}\frac{1}{\hat{\pi}^{B}(X)}\right]+O_{p}(M^{-1/2})$
$\hskip 22.76228pt-E_{X,R^{B}}\left[\frac{1}{n}\sum_{i\in\mathcal{S}}\{R^{B}-\pi^{B}(X)\}\;\frac{1}{\hat{\pi}^{B}(X_{i})}\right]+O_{p}(M^{-1/2})$
$\hskip 8.5359pt=E_{X}\left[\left\{1-\frac{\pi^{B}(X)}{\hat{\pi}^{B}(X)}\right\}\frac{1}{\hat{\pi}^{B}(X)}\right]+O_{p}(M^{-1/2})$
$\hskip 8.5359pt\leq\sqrt{E_{X}\left[\left\{1-\frac{\pi^{B}(X)}{\hat{\pi}^{B}(X)}\right\}^{2}\right]\;E_{X}\left[\frac{1}{\hat{\pi}^{B}(X)^{2}}\right]}+O_{p}(M^{-1/2})$ (A44)
$\hskip 8.5359pt=\sqrt{o_{p}(M^{-c_{\pi}})}+O_{p}(M^{-1/2})$
$\hskip 8.5359pt=o_{p}(M^{-c_{\pi}/2}).$ (A45)
Note that line (A44) uses the Cauchy-Schwartz inequality.
If $\hat{m}(x;\epsilon_{k})$ is instead defined by equation (9), then
$\left.\frac{d}{d\epsilon_{k}}\left[\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}U_{i}\{\hat{\pi}^{B}_{k},\hat{m}_{k}(\epsilon_{k})\}\right]\right|_{\epsilon_{k}=0}$ $=$ $\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}\left\{\frac{R^{A}_{i}}{\pi^{A}_{i}}-\frac{R^{B}_{i}}{\hat{\pi}^{B}(X_{i})}\right\}\frac{1}{\hat{\pi}^{B}(X_{i})}\frac{\hat{m}(X_{i})}{\{1+\hat{m}(X_{i})\}^{2}},$
and the same argument shows that this is $o_{p}(M^{-c_{\pi}/2})$.
By using a similar argument to that used in Appendices A1 and A2, the term
$\frac{1}{n}\sum_{i\in\mathcal{S}}\left(\frac{R^{A}_{i}}{\pi^{A}_{i}}-1\right)\frac{1}{\hat{\pi}^{B}(X_{i})}$
in expression (A43) can be shown to be $O_{p}(M^{-1/2})$ given any fixed $\hat{\pi}^{B}_{k}$.
We shall now show that $\hat{\epsilon}_{k}$ is $o_{p}(M^{-c_{m}/2})$. First, consider the case where $\hat{m}_{k}(x;\epsilon_{k})$ is defined by equation (8). For any fixed $\hat{\pi}^{B}$ and $\hat{m}$, the first iteration of a Newton-Raphson algorithm to find the maximum likelihood estimate of $\epsilon_{k}$ starting from an initial value of $\hat{\epsilon}_{k}=0$ would set $\hat{\epsilon}_{k}$ equal to
$\hat{\epsilon}_{k}$ $=$ $\frac{1}{n_{k}}\left.\sum_{i\in\mathcal{S}_{k}}\frac{R^{B}_{i}}{\hat{\pi}^{B}_{k}(X_{i})}\{Y_{i}-\hat{m}_{k}(X_{i})\}\right/\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{R^{B}_{i}}{\hat{\pi}^{B}_{k}(X_{i})^{2}}$
$=$ $\left.E_{X}\left[\frac{\pi^{B}(X)}{\hat{\pi}^{B}_{k}(X)}\{m(X)-\hat{m}_{k}(X)\}\right]\right/E_{X}\left\{\frac{\pi^{B}(X)}{\hat{\pi}^{B}_{k}(X)^{2}}\right\}\;+O_{p}(M^{-1/2}).$
Now, by the Cauchy-Schwartz inequality,
$E_{X}\left[\frac{\pi^{B}(X)}{\hat{\pi}^{B}_{k}(X)}\{m(X)-\hat{m}_{k}(X)\}\right]$ $\leq$ $\sqrt{E_{X}\left\{\frac{\pi^{B}(X)^{2}}{\hat{\pi}^{B}_{k}(X)^{2}}\right\}E_{X}[\{m(X)-\hat{m}_{k}(X)\}^{2}]}$
$=$ $\sqrt{O_{p}(1)\times o_{p}(N^{-c_{m}})}$
$=$ $o_{p}(N^{-c_{m}/2}).$
Also,
$E_{X}\left\{\frac{\pi^{B}(X)}{\hat{\pi}^{B}_{k}(X)^{2}}\right\}=E_{X}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X)}\right\}+o_{p}(1).$
Hence,
$\hat{\epsilon}_{k}$ $=$ $\frac{o_{p}(N^{-c_{m}/2})}{E_{X}\left\{\frac{1}{\hat{\pi}^{B}_{k}(X)}\right\}+o_{p}(1)}+O_{p}(M^{-1/2})$
$=$ $o_{p}(N^{-c_{m}/2}).$
Since the first iteration of the Newton-Raphson algorithm leads to $\hat{\epsilon}_{k}=o_{p}(N^{-c_{m}/2})$, the value of $\hat{\epsilon}_{k}$ at convergence of the algorthm will also be $o_{p}(N^{-c_{m}/2})$.
Second, consider the case where $\hat{m}_{k}(x;\epsilon_{k})$ is defined by equation (9). Now, for any fixed $\hat{\pi}^{B}$ and $\hat{m}$, the first iteration of a Newton-Raphson algorithm to find the maximum likelihood estimate of $\epsilon_{k}$ starting from an initial value of $\hat{\epsilon}_{k}=0$ would set $\hat{\epsilon}_{k}$ equal to
$\hat{\epsilon}_{k}$ $=$ $\frac{1}{n_{k}}\left.\sum_{i\in\mathcal{S}_{k}}\frac{R^{B}_{i}}{\hat{\pi}^{B}_{k}(X_{i})}\{Y_{i}-\hat{m}_{k}(X_{i})\}\right/\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{R^{B}_{i}}{\hat{\pi}^{B}_{k}(X_{i})^{2}}\frac{\hat{m}_{k}(X_{i})}{\{1+\hat{m}_{k}(X_{i})\}^{2}}.$
The same logic that was used in the case of where $\hat{m}_{k}(x;\epsilon_{k})$ is defined by equation (8) again applies, and hence $\hat{\epsilon}_{k}=o_{p}(N^{-c_{m}/2})$.
Putting this together, we have
$\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})+\left.\frac{\partial}{\epsilon_{k}}\left[\frac{1}{n}\sum_{i\in\mathcal{S}_{k}}U_{i}\{\hat{\pi}^{B}_{k},\hat{m}_{k}(\epsilon_{k})\}\right]\right|_{\epsilon_{k}=0}\hat{\epsilon}_{k}$ $=$ $o_{p}(M^{-c_{\pi}/2})\times o_{p}(N^{-c_{m}/2})$
$=$ $o_{p}(M^{-1/2}).$
Hence, $\hat{\theta}_{\rm TMLE1}=\hat{\theta}_{1}+o_{p}(M^{1/2})$.

### v2

@@
 The third class, known as double robust (DR) methods, combine IPW and mass imputation.
-IPW methods require a consistent estimator of the selection probability function for Sample B, i.e.
-the conditional probability that an individual belongs to Sample B given $X$.
+IPW methods require a consistent estimator of the conditional probability that an individual belongs to Sample B given $X$.
 Mass imputation methods require a consistent estimator of the conditional expectation of $Y$ given $X$.
@@
 A parametric model could be used for each of the two nuisance functions, but the resulting estimator would typically be inconsistent if that parametric model — or, in the case of DR estimators, both parametric models — were misspecified.
-There is therefore considerable interest in using more flexible ‘data-adaptive’ or ‘machine-learning’ estimators of the nuisance functions, in order to minimise this risk of model misspecification.
+There is therefore considerable interest in using more flexible ‘data-adaptive’ or ‘machine-learning’ estimators of the nuisance functions in order to minimise this risk of model misspecification.
 A number of researchers have done this.
@@
 (2023)[14] and Rueda et al.
-(2024)[12] use IPW and estimate the conditional probability nuisance function with k-nearest neighbours, random forest or XGboost.
+(2024)[13] use IPW and estimate the conditional probability nuisance function with k-nearest neighbours, random forest or generalised boosting models (specifically, XGBoost).
 Castro-Martin et al.
@@
 They note, however, that this method is subject to the curse of dimensionality.
-A closely related problem to that of estimating the mean of $Y$ in a finite population by combining data from probability and nonprobability samples is that of estimating $E(Y)$, the expectation of $Y$ in an infinite population, when data $(X_{i},Y_{i})$ $(i=1,\ldots,n)$ are independent and identically distributed (iid), $X$ is fully observed and $Y$ is missing at random given $X$.
-IPW, imputation and DR methods exist for this problem.
+Chlebicki et al.
+(2024)[10] use mass imputation with predictive mean matching.
+This matching provides some limited robustness to misspecification of a parametric model for the nuisance conditional expectation function, because it requires only that the true conditional expectation be a monotone function of the limiting value of the estimated conditional expectation.
+Chlebicki et al.
+suggest that a data-adaptive method could be used to estimate this nuisance function, but do not prove that the resulting estimator of $\bar{Y}$ would be $n^{1/2}$-consistent.
+A closely related problem to that of estimating $\bar{Y}$ for a finite population by combining data from probability and nonprobability samples is that of estimating $E(Y)$, the expectation of $Y$ in an infinite population, when data $(X_{i},Y_{i})$ $(i=1,\ldots,n)$ are independent and identically distributed (iid), $X$ is fully observed and $Y$ is missing at random given $X$.
+IPW, regression imputation and DR methods exist for this problem.
 The first two require consistent estimators of, respectively, the conditional probability that $Y$ is observed given $X$ and the conditional expectation of $Y$ given $X$.
-DR methods involve estimating both these nuisance functions, but require only that (at least) one of them is estimated consistently.
+DR methods involve estimating both these nuisance functions, but require only that one of them is estimated consistently.
 For IPW and imputation, it is known that using data-adaptive methods to estimate the nuisance function can cause problems.
@@
 The DR approach (also known as ‘debiased machine learning’) can overcome this issue, because DR estimators do not inherit the slow convergence of the nuisance function estimators[27, 15, 24, 21, 8, 7, 9].
-Given the similarity of the problem of estimating $E(Y)$ for an infinite population with iid data to the problem of estimating the mean of $Y$ in a finite population using a nonprobability sample, there is reason to be concerned that the same issue may well also apply to the latter problem.
-In this article, we study the use of a DR estimator of the mean of $Y$ in a finite population when combining data from a probability sample and a nonprobability sample and using data-adaptive estimators of the nuisance functions.
+Given the similarity of the problem of estimating $E(Y)$ for an infinite population with iid data to the problem of estimating $\bar{Y}$ for a finite population using a nonprobability sample, there is reason to be concerned that the same issue may well also apply to the latter problem.
+In this article, we study the use of a DR estimator of $\bar{Y}$ for a finite population when combining data from a probability sample and a nonprobability sample and using data-adaptive estimators of the nuisance functions.
 Our goal is to provide a theoretical justification for the validity of this approach.
-An important difference between the problem of estimating $E(Y)$ for an infinite population with iid data and the problem we study here is that when, as is common, the probability sample is obtained using sampling without replacement, Sample A data are not iid.
+An important difference between the problem of estimating $E(Y)$ for an infinite population with iid data and the problem we study here is that when, as is common, the probability sample is obtained using sampling without replacement (SWOR), Sample A data are not independent.
 We build on the work of Chen, Li and Wu (2020)[6] (henceforth, CLW).
@@
 Yang et al.
-(2020)[29] extended this method to allow for variable selection by smoothly clipped absolute deviation (SCAD).
+(2020)[30] extended this method to allow for variable selection by smoothly clipped absolute deviation.
 We generalise the work of CLW (and Yang et al.) to allow for general data-adaptive estimation of the nuisance functions.
@@
 The structure of the article is as follows.
-In Section 2, we describe our design-model-based inference framework.
-In Section 3, we describe a von Mises expansion and cross-fitting, both fundamental to our approach.
-In Section 4, we consider the scenario where clusters (or ‘primary sampling units’) are sampled using simple random sampling (i.e.
-with equal probabilities) without replacement and then individuals within clusters are sampled using an arbitrary (possibly multi-stage) sampling design that obeys the independence and invariance conditions[24].
+In Section 2, we describe our assumed data-generating mechanism and design-model-based inference framework.
+In Section 3, we describe a DR estimator for $\bar{Y}$, its von Mises expansion and cross-fitting, which is fundamental to our approach.
+In Section 4, we consider the scenario where clusters (or ‘primary sampling units’) are sampled using simple random sampling without replacement (SRSWOR) and then individuals within clusters are sampled using an arbitrary (possibly multi-stage) sampling design that obeys the independence and invariance conditions[25].
 Section 5 considers the more general scenario where the cluster sampling probabilities vary.
-In Section 6, we describe a ratio-style DR estimator of $\bar{Y}$ that is typically more efficient than the DR estimator described in earlier sections.
-Section 7 covers targeted maximum likelihood estimators (TMLEs), which are an alternative to the estimating equations-style DR estimators described in earlier sections.
+In Section 6, we describe a ratio-type DR estimator of $\bar{Y}$ that is typically more efficient than the DR estimator described in earlier sections.
+Section 7 covers targeted maximum likelihood estimators (TMLEs), which are an alternative to the estimating equations-type DR estimators described in earlier sections.
 Possible data-adaptive estimators of the nuisance functions are discussed in Section 8.
@@
 This, along with other issues, is discussed in Section 10.
-2 Data-generating mechanism and estimator
+2 Assumed data-generating mechanism
 Our assumed data-generating mechanism consists of a superpopulation model for generating the finite population, a model for drawing Sample A from the finite population, and a model for drawing Sample B from the finite population.
@@
 The variable $R^{A}_{ji}$ will be a binary variable indicating whether individual $i$ in cluster $j$ ($i=1,\ldots,N_{j};\;j=1,\ldots,J$) is included in Sample A.
-First, a sample of size $M$ ($M<J$) clusters is drawn from the $J$ clusters in the population, with the probability that cluster $j$ is included in the sample being proportional to $h(N_{j},X_{j1},\ldots,X_{jN_{j}})$ for some function $h$ of $N_{j}$ and $X_{j1},\ldots,X_{jN_{i}}$.
-A possible example would be $h(N_{j},X_{j1},\ldots,X_{jN_{j}})=N_{j}$.
-Let $R^{C}_{j}=1$ if cluster $j$ is drawn, and $R^{C}_{j}=0$ otherwise.
+First, a sample of size $M$ ($M<J$) clusters is drawn without replacement from the $J$ clusters in the population, with the probability that cluster $j$ is included in the sample being proportional to $h(N_{j},X_{j1},\ldots,X_{jN_{j}})$ for some function $h$ of $N_{j}$ and $X_{j1},\ldots,X_{jN_{i}}$.
+An example would be $h(N_{j},X_{j1},\ldots,X_{jN_{j}})=N_{j}$.
+Let $R^{C}_{j}=1$ if cluster $j$ is drawn; $R^{C}_{j}=0$ otherwise.
 If $R^{C}_{j}=0$, then $R^{A}_{ji}=0$ ($i=1,\ldots,N_{j}$).
-If $R^{C}_{j}=1$, then $(R^{A}_{j1},\ldots,R^{A}_{jN_{j}})$ is drawn according to some (possibly multistage) sampling design that can depend on $N_{j}$ and $(X_{j1},\ldots,X_{jN_{j}})$ but which, given $N_{j}$, $(X_{j1},\ldots,X_{jN_{j}})$ and $R^{C}_{j}=1$, does not depend on $(Y_{j1},\ldots,Y_{jN_{j}})$, $\{(N_{k},X_{k1},\ldots,X_{kN_{k}},Y_{k1},\ldots,Y_{kN_{k}}):\;k=1,\ldots,j-1,j+1,\ldots J\}$ or $(R^{C}_{1},\ldots,R^{C}_{j-1},R^{C}_{j+1},\ldots,R^{C}_{J}$).
+If $R^{C}_{j}=1$, then $(R^{A}_{j1},\ldots,R^{A}_{jN_{j}})$ is drawn according to some (possibly multistage) SWOR design that can depend on $N_{j}$ and $(X_{j1},\ldots,X_{jN_{j}})$ but which, given $N_{j}$, $(X_{j1},\ldots,X_{jN_{j}})$ and $R^{C}_{j}=1$, does not depend on $(Y_{j1},\ldots,Y_{jN_{j}})$, $\{(N_{k},X_{k1},\ldots,X_{kN_{k}},Y_{k1},\ldots,Y_{kN_{k}}):\;k=1,\ldots,j-1,j+1,\ldots J\}$ or $(R^{C}_{1},\ldots,R^{C}_{j-1},R^{C}_{j+1},\ldots,R^{C}_{J}$).
 This sampling of $(R^{A}_{j1},\ldots,R^{A}_{jN_{j}})$ is done independently of the sampling of $(R^{A}_{k1},\ldots,R^{A}_{kN_{k}})$ within any other cluster $k$ with $R^{C}_{k}=1$.
@@
 Given $\mathcal{F}_{J}$, each $R^{B}_{ji}$ ($i=1,\ldots,N_{j};\;j=1,\ldots,J$) is independently sampled from a Bernoulli distribution with probability parameter $\pi^{B}_{0}(X_{ji})$ independently of $Y_{11},\ldots,Y_{1N_{1}},\ldots Y_{J1},\ldots,Y_{JN_{J}}$, $R^{C}_{1},\ldots,R^{C}_{J}$ and $R^{A}_{11},\ldots,R^{A}_{1N_{1}},\ldots R^{A}_{J1},\ldots,R^{A}_{JN_{J}}$.
-Now relabel the $\sum_{j=1}^{J}N_{j}$ individuals so that, for each $j=1,\ldots,J$, the $N_{j}$ individuals in cluster $j$ are labelled as individuals $\sum_{k=1}^{j-1}N_{j}+1,\sum_{k=1}^{j-1}N_{j}+2,\ldots,\sum_{k=1}^{j}N_{j}$.
-Let $D_{i}$ ($1\leq D_{i}\leq J$) denote the number of the original cluster to which individual $i$ belongs.
+Now relabel the $\sum_{j=1}^{J}N_{j}$ individuals so that, for each $j=1,\ldots,J$, the $N_{j}$ individuals in cluster $j$ are labelled as individuals $\sum_{k=1}^{j-1}N_{k}+1,\sum_{k=1}^{j-1}N_{k}+2,\ldots,\sum_{k=1}^{j}N_{k}$.
+Let $D_{i}$ ($1\leq D_{i}\leq J$) denote the number of the cluster to which individual $i$ belongs.
 Finally, let $n=n(\mathcal{F}_{J})=\sum_{j=1}^{J}N_{j}$ denote the total number of individuals in the population.
-We can now write the population mean as $\bar{Y}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$.
-The observed data are
-$\{(R^{A}_{i},R^{A}_{i}X_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}D_{i},R^{B}_{i},R^{B}_{i}X_{i},R^{B}Y_{i},R^{B}_{i}D_{i});\;i=1,\ldots,n\},$
-i.e.
-the $X$ values and sampling probabilities of individuals in Sample A, the $X$ and $Y$ values of individuals in Sample B, and the index of the cluster to which each of these individuals belongs.
+We can now write the population mean as $\bar{Y}=n^{-1}\sum_{i=1}^{n}Y_{i}$.
+The observed data are the $X$ values and sampling probabilities $\pi^{A}$ of individuals in Sample A, the $X$ and $Y$ values of individuals in Sample B, and the index of the cluster to which each of these individuals belongs, i.e.
+$\{(R^{A}_{i},R^{A}_{i}X_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}D_{i},R^{B}_{i},R^{B}_{i}X_{i},R^{B}Y_{i},R^{B}_{i}D_{i}):\;i=1,\ldots,n\}.$
 Note that our notation suggests that we know whether an individual in Sample A also appears in Sample B, and vice versa.
 In fact, this is unnecessary; this notation is used only for convenience.
-We shall, however, need to know, for each individual in Sample A and each individual in Sample B, whether they belong to the same cluster (this is used in the cross-fitting procedure — see Section 3).
-[…]
-$\hat{\theta}_{1}(\pi^{B},m)=\frac{1}{n}\sum_{i=1}^{n}U_{i}(\pi^{B},m),$
-$U(\pi^{B},m)=U(\pi^{B},m;X,Y,R^{A},R^{B},\pi^{A})=\frac{R^{B}}{\pi^{B}(X)}Y+\left\{\frac{R^{A}}{\pi^{A}}-\frac{R^{B}}{\pi^{B}(X)}\right\}m(X),$
-[…]
-3 von Mises expansion
-3.1 Independent, identically distributed data and infinite population
-[…]
-$\hat{\theta}_{\rm iid}(\pi_{\rm iid},m_{\rm iid})=\frac{1}{n}\sum_{i=1}^{n}\frac{R_{i}}{\pi_{\rm iid}(X_{i})}Y_{i}+\left\{1-\frac{R_{i}}{\pi_{\rm iid}(X_{i})}\right\}m_{\rm iid}(X_{i}).$
-Here, $R$ is an indicator that $Y$ is observed, $\hat{\pi}_{\rm iid}(X)$ is an estimator of $\pi_{\rm iid0}(X)=P(R=1\mid X)$ and $\hat{m}_{\rm iid}(X)$ is an estimator of $m_{\rm iid0}(X)=E(Y\mid X)$.
+We shall, however, need to know, for each individual in Sample A and each individual in Sample B, whether they belong to the same cluster (this information is used in the cross-fitting procedure — see Section 3).
+We shall consider the asymptotic repeated-sampling distributions of several DR estimators of $\bar{Y}$ conditional on $\mathcal{F}_{J}$ in an asymptotic framework in which $J\rightarrow\infty$ and $M\rightarrow\infty$ with $M/J$ converging to a constant.
+This means we do not need to distinguish between $M\rightarrow\infty$ and $J\rightarrow\infty$.
+The data-generating models for drawing Samples A and B from the finite population do not change as $M\rightarrow\infty$.
+3 DR estimator and its von Mises expansion
+3.1 iid data and infinite population
+It is instructive to consider the simpler context of estimating $E(Y)$ for an infinite population using a sample of $n$ iid individuals with $X$ observed for all $n$ individuals, $Y$ observed for only some of them, and these data being missing at random.
+In this context, a DR estimator for $E(Y)$ is $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$, where
+$\hat{\theta}_{\rm iid}(\pi_{\rm iid},m_{\rm iid})=\frac{1}{n}\sum_{i=1}^{n}m_{\rm iid}(X_{i})+\frac{R_{i}}{\pi_{\rm iid}(X_{i})}\{Y_{i}-m_{\rm iid}(X_{i})\},$
+$R$ is an indicator that $Y$ is observed, $\hat{\pi}_{\rm iid}(X)$ is an estimator of $\pi_{\rm iid0}(X)=P(R=1\mid X)$ and $\hat{m}_{\rm iid}(X)$ is an estimator of $m_{\rm iid0}(X)=E(Y\mid X)$.
 Hines et al.
-(2022)[20] give an account of the properties of $\hat{\theta}_{\rm iid}$ and of the von Mises expansion that can be used to derive these properties.
-They describe how, if $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ converge to $\pi_{\rm iid0}$ and $m_{\rm iid0}$ sufficiently rapidly as $n\rightarrow\infty$, then $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})=\hat{\theta}_{\rm iid}(\pi_{\rm iid0},m_{\rm iid0})+o_{p}(n^{-1/2})$, and so $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$ has the same asymptotic normal distribution as $\hat{\theta}_{\rm iid}(\pi_{\rm iid0},m_{\rm iid0})$.
-The required convergence rates of $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ are slower than the parametric $n^{1/2}$ rate, which enables the use of data-adaptive estimators.
-Often, $\hat{\theta}_{\rm iid}$ is used with cross-fitting, a technique similar to cross-validation.
-This involves splitting the sample of $n$ individuals into subsamples, called folds, calculating estimates $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ using the data on all but fold $k$, and then calculating $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$ using these estimates and the data on fold $k$.
-This procedure is repeated for $k=1,\ldots,K$, and the resulting $K$ values of $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$ are averaged.
-This ensures that $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ are independent of the data in fold $k$, which avoids the need for $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ to satisfy the Donsker condition[20].
+(2022)[21] describe how a von Mises expansion can be used to show that if $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ converge to $\pi_{\rm iid0}$ and $m_{\rm iid0}$ sufficiently fast as $n\rightarrow\infty$, then $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})=\hat{\theta}_{\rm iid}(\pi_{\rm iid0},m_{\rm iid0})+o_{p}(n^{-1/2})$, and so $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$ has the same asymptotic normal distribution as $\hat{\theta}_{\rm iid}(\pi_{\rm iid0},m_{\rm iid0})$.
+The required convergence rates of $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ are slower than the parametric $n^{1/2}$ rate, which enables the use of data-adaptive estimators, provided that either these estimators satisfy the Donsker condition or cross-fitting is used.
 3.2 Sampling without replacement from a finite population
-In this article, we obtain an analogous asymptotic equivalence result for our sample survey setting.
+In this article, we obtain an analogous result for our sample survey setting.
 We make use of cross-fitting.
@@
 $K=5$) subsets, called folds.
-The way in which is done will be described later.
-Let $\mathcal{S}_{k}\subset\{1,\ldots,n\}$ denotes the set of indices of the individuals belonging to fold $k$ ($k=1,\ldots,K$).
-Let $n_{k}=|\mathcal{S}_{k}|$ denote the number of individuals in $\mathcal{S}_{k}$.
-Let $\mathcal{S}_{-k}$ denote the indices of all individuals except those in fold $k$ (so, $\mathcal{S}_{k}\cup\mathcal{S}_{-k}=\{1,\ldots,n\}$).
+The way this is done will be described in Sections 4 and 5.
+Let $\mathcal{S}_{k}\subset\{1,\ldots,n\}$ denote the set of indices of the individuals belonging to fold $k$ ($k=1,\ldots,K$), and let $n_{k}=|\mathcal{S}_{k}|$ denote the number of these individuals.
+Let $\mathcal{S}_{-k}$ denote the indices of the individuals not in fold $k$ (so, $\mathcal{S}_{k}\cup\mathcal{S}_{-k}=\{1,\ldots,n\}$).
 Let $\hat{m}_{k}$ be an estimator of $m_{0}$ calculated using only the data $\{(X_{i},Y_{i}):R^{B}_{i}=1\mbox{ and }i\in\mathcal{S}_{-k}\}$.
-Let $\hat{\pi}^{B}_{k}$ be an estimator of $\pi^{B}_{0}$ obtained using only the data $\{(R^{A}_{i},R^{B}_{i}):i\in\mathcal{S}_{-k}\}$ and $\{X_{i}:R^{A}_{i}+R^{B}_{i}\geq 1\mbox{ and }i\in\mathcal{S}_{-k}\}$.
+Let $\hat{\pi}^{B}_{k}$ be an estimator of $\pi^{B}_{0}$ obtained using only the data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i}):i\in\mathcal{S}_{-k}\}$.
+We discuss possible estimators in Section 8.
 Write $\underline{\pi}^{B}=(\pi^{B}_{1},\ldots,\pi^{B}_{K})$, $\underline{m}=(m_{1},\ldots,m_{K})$, $\underline{\hat{\pi}}^{B}=(\hat{\pi}^{B}_{1},\ldots,\hat{\pi}^{B}_{K})$ and $\underline{\hat{m}}=(\hat{m}_{1},\ldots,\hat{m}_{K})$.
-In a slight abuse of notation, we define
-$\hat{\theta}_{1}(\underline{\pi}^{B},\underline{m})=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U_{i}(\pi^{B}_{k},m_{k}).$
 […]
+$\hat{\theta}_{1}(\underline{\pi}^{B},\underline{m})=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U_{i}(\pi^{B}_{k},m_{k})$
+$U_{i}(\pi^{B},m)=\frac{R^{A}_{i}}{\pi^{A}_{i}}m(X_{i})+\frac{R^{B}_{i}}{\pi^{B}(X_{i})}\{Y_{i}-m(X_{i})\}.$
+We shall use $\hat{\theta}_{1}(\pi^{B},m)$ as shorthand for $\hat{\theta}_{1}\big((\pi^{B},\ldots,\pi^{B}),(m,\ldots,m)\big)$.
+Note that $\hat{\theta}_{1}$ is analogous to $\hat{\theta}_{\rm iid}$, but with the weighted mean of $m(X)$ in Sample A substituting for the unknown mean of $m(X)$.
+We shall assume that $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ satisfy the following condition.
 Condition C1: There exist $c_{\pi}>0$ and $c_{m}>0$ such that $c_{\pi}+c_{m}=1$,
@@
 $E_{X}\left[\{\hat{m}_{k}(X)-m_{0}(X)\}^{2}\right]=o_{p}(M^{-c_{m}}).$ (2)
-Condition C1 allows convergence rates that are considerably slower than those of parametric estimators [20].
-For example, Condition C1 is satisfied when both $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ converge faster than the $M^{1/4}$ rate.
-Our proof uses the following von Mises expansion of $\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}$.
+Condition C1 allows convergence rates for $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ considerably slower than those of parametric estimators[21].
+For example, it is satisfied when both $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ converge faster than the $M^{1/4}$ rate.
+We shall exploit the following von Mises expansion of $\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}$:
 $\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}$
@@
 $\hskip 113.81102pt-\left.E\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}\right]$
-$\hskip 28.45274pt+\sum_{k=1}^{K}\frac{n_{k}}{n}\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}E\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}.$
+$\hskip 28.45274pt+\sum_{k=1}^{K}\frac{n_{k}}{n}\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}E\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}.$ (3)
 The first term on the right-hand side of equation (3) is just $\hat{\theta}_{1}(\pi^{B}_{0},m_{0})-\bar{Y}$.
@@
 We shall prove that these two terms are $o_{p}(M^{-1/2})$ provided that Condition C1 holds.
-It follows from this that
+It then follows that $\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}$ has the same asymptotic distribution as $\hat{\theta}_{1}(\pi^{B}_{0},m_{0})-\bar{Y}$, i.e.
 $\sqrt{M}\{\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}\}$
 $\hskip 28.45274pt=\sqrt{M}\{\hat{\theta}_{1}(\pi^{B}_{0},m_{0})-\bar{Y}\}+o_{p}(1)$
-$\hskip 28.45274pt=\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}m_{0}(X_{i})+\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{B}_{i}}{\pi^{B}_{0}(X_{i})}\{Y_{i}-m_{0}(X_{i})\}+o_{p}(1).$
+$\hskip 28.45274pt=\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}m_{0}(X_{i})+\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{B}_{i}}{\pi^{B}_{0}(X_{i})}\{Y_{i}-m_{0}(X_{i})\}$
+$\hskip 42.67912pt-\sqrt{M}\bar{Y}+o_{p}(1).$ (4)
 The asymptotic variance of $\hat{\theta}_{1}(\pi^{B}_{0},m_{0})-\bar{Y}$ is given by CLW, along with an estimator of this variance.
@@
 Moreover, subject to regularity conditions, $\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}$ is asymptotically normally distributed[4].
-An important complication when proving that the empirical process and remainder term are $o_{p}(M^{-1/2})$ is that, in general, $\hat{\pi}^{B}_{k}(X)$ is not independent of the data in fold $k$.
-This is because $\hat{\pi}^{B}_{k}(X)$ is calculated using the $R^{A}_{i}$ values of individuals in $\mathcal{S}_{-k}$, $\sum_{i\in\mathcal{S}_{k}}U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})$ is calculated using the $R^{A}_{i}$ values of individuals in $\mathcal{S}_{k}$, and $\{R^{A}_{i}:\;i\in\mathcal{S}_{-k}\}$ and $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ are, in general, not independent when clusters are sampled without replacement.
+An important complication when proving that the empirical process and remainder term are $o_{p}(M^{-1/2})$ is that, in general, $\hat{\pi}^{B}_{k}$ is not independent of the data in fold $k$.
+This is because $\hat{\pi}^{B}_{k}(X)$ is calculated using the $R^{A}_{i}$ values of individuals in $\mathcal{S}_{-k}$, and $\sum_{i\in\mathcal{S}_{k}}U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})$ is calculated using the $R^{A}_{i}$ values of individuals in $\mathcal{S}_{k}$, and $\{R^{A}_{i}:\;i\in\mathcal{S}_{-k}\}$ and $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ are not independent when clusters are selected using SWOR.
 For example, the number of clusters with $R^{C}_{j}=1$ in fold $k$ must equal $M$ minus the number of clusters with $R^{C}_{j}=1$ in the remaining folds.
-4 Simple random sampling of clusters without replacement
-Consider the case of simple random sampling of clusters without replacement (so, $\pi^{C}_{1}=\ldots=\pi^{C}_{J}$).
+We address this complication in Sections 4 and 5.
+4 Simple random sampling without replacement of clusters
+First, consider the case of SRSWOR of clusters (so, $\pi^{C}_{1}=\ldots=\pi^{C}_{J}$).
 As described in Section 2, sampling of individuals within sampled clusters is done using some (possibly multistage) sampling design.
-Suppose that $M$ and $J$ are both integer multiples of $K$ (the case where $M$ and/or $J$ are not multiples of $K$ is considered as a special case in Section 5).
+Suppose that $M$ and $J$ are both integer multiples of $K$; the case where $M$ and/or $J$ are not multiples of $K$ is considered as a special case in Section 5.
 Randomly partition the $M$ sampled clusters (i.e.
@@
 those with $R^{C}_{j}=0$) evenly into the $K$ folds, with all such partitions being equally probable.
-Each fold contains $M/K$ sampled clusters and (JM)/K(J-M)/K unsampled clusters.
-This ensures that the conditional distribution of $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J}$, $\mathcal{S}_{k}$ and $\{R^{A}_{i}:\;i\in\mathcal{S}_{-k}\}$ is the same as the distribution of $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$, and corresponds to simple random sampling of $M/K$ clusters without replacement from the $J/K$ clusters in fold $k$ followed by the original second-stage sampling mechanism for individuals within clusters.
+Thus, each fold contains $M/K$ sampled clusters and (JM)/K(J-M)/K unsampled clusters.
+This ensures that the conditional distribution of $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J}$, $\mathcal{S}_{k}$ and $\{R^{A}_{i}:\;i\in\mathcal{S}_{-k}\}$ is the same as the distribution of $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$, and corresponds to SRSWOR of $M/K$ clusters from the $J/K$ clusters in fold $k$ followed by the original second-stage sampling mechanism for individuals within clusters.
 Recall that $Y_{i}$ is assumed to be conditionally independent of $R^{A}_{i^{\prime}}$, $R^{C}_{j}$ and $Y_{i^{\prime\prime}}$ given $\mathcal{F}_{J}$ for all $i$, $i^{\prime}$ and $j$ and all $i^{\prime\prime}\neq i$.
@@
 Because the choice of folds does not depend on $Y$ or $R^{B}$ values, these conditional independences also hold conditional on $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$.
-All this implies that the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{k}\}$ on fold $k$ are conditionally independent of the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{-k}\}$ on the remaining folds given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$.
-Hence, since $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ are calculated using only the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{-k}\}$, the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{k}\}$ are conditionally independent of $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$.
-In Appendix A1 we show that this ensures that the empirical process and remainder terms are $o_{p}(M^{-1/2})$, as required, provided that Condition C1 holds.
-5 Sampling of clusters without replacement with varying cluster sampling probabilities
-Now consider the more general case where there are a small number, $L$, of distinct values of the cluster sampling probability $\pi^{C}_{j}$, which we denote as $\pi^{C(1)},\ldots,\pi^{C(L)}$.
-Simple random sampling of clusters without replacement is the special case of this with $L=1$.
+All this implies that the data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{k}\}$ on fold $k$ are conditionally independent of the data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{-k}\}$ on the remaining folds given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$.
+Hence, since $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ are calculated using only the data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{-k}\}$, the data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{k}\}$ are conditionally independent of $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$.
+In Appendix A1 we show that this ensures that the empirical process and remainder terms are $o_{p}(M^{-1/2})$, provided that Condition C1 holds.
+5 Varying cluster sampling probabilities
+Now consider the more general case where there are $L$ distinct values of the cluster sampling probability $\pi^{C}_{j}$, which we denote as $\pi^{C(1)},\ldots,\pi^{C(L)}$.
+Assume that $L$ is small relative to $M$.
+SRSWOR of clusters is the special case of this with $L=1$.
 Let $M^{(l)}=M^{(l)}(R^{C}_{1},\ldots,R^{C}_{J})=\sum_{j=1}^{J}R^{C}_{j}\;I(\pi^{C}_{j}=\pi^{C(l)})$ and $J^{(l)}=\sum_{j=1}^{J}I(\pi^{C}_{j}=\pi^{C(l)})$ denote, respectively, the number of sampled clusters and the total number of clusters with cluster sampling probability $\pi^{C(l)}$ ($l=1,\ldots,L$).
 Clearly, $\sum_{l=1}^{L}J^{(l)}=J$ and $\sum_{l=1}^{L}M^{(l)}=M$.
-Assume that the cluster sampling mechanism satisfies the following condition:
+Assume that the cluster sampling mechanism satisfies the following symmetry condition:
 $P(R^{C}_{1}=r_{1},\ldots,R^{C}_{1}=r_{J}\mid\mathcal{F}_{J})=P(R^{C}_{1}=r_{1}^{\prime},\ldots,R^{C}_{1}=r_{J}^{\prime}\mid\mathcal{F}_{J})$
@@
 This condition is satisfied by, for example, conditional Poisson sampling, Sampford sampling, Pareto sampling and randomised systematic sampling[1, 20].
-For each of $l=1,\ldots,L$, randomly partition the $M^{(l)}$ sampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ evenly into $K$ folds, and randomly partition the $J^{(l)}-M^{(l)}$ unsampled clusters evenly into the same $K$ folds.
-For each $l$, this partition begins by randomly choosing a set of $K\times\lfloor M^{(l)}/K\rfloor$ sampled clusters and a set of $K\times\lfloor(J^{(l)}-M^{(l)})/K\rfloor$ unsampled clusters ($\lfloor x\rfloor$ denotes the integer part of $x$) and evenly partitioning each of these two sets between the folds.
-This leaves fewer than $K$ sampled clusters and fewer than $K$ unsampled clusters so far unassigned to folds.
-Each of the unassigned sampled clusters is randomly assigned to the $K$ folds in such a way that no fold receives more than one of these clusters, and the same is done with the unassigned unsampled clusters.
+For each of $l=1,\ldots,L$, randomly partition the $M^{(l)}$ sampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ evenly into $K$ folds, and randomly partition the $J^{(l)}-M^{(l)}$ unsampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ evenly into the same $K$ folds.
+For each $l$, this partition begins by randomly choosing a set of $K\times\lfloor M^{(l)}/K\rfloor$ sampled clusters ($\lfloor x\rfloor$ denotes the integer part of $x$) and evenly partitioning these between the folds.
+This leaves fewer than $K$ sampled clusters so far unassigned to folds.
+Each of these is randomly assigned to the $K$ folds in such a way that no fold receives more than one of them.
+The same is done with the unassigned unsampled clusters.
 […]
-As in Section 4, $\hat{m}_{k}$ is calculated using the $Y$ values of all individuals with $R^{B}=1$ not in fold $k$, and $\hat{\pi}^{B}_{k}$ is calculated using $R^{B}$ values of all the individuals not in fold $k$.
-However, unlike in Section 4, this calculation of $\hat{\pi}^{B}_{k}$ uses the $R^{A}$ values of only individuals in a subset of the clusters not in fold $k$.
-We shall refer to this subset as the ‘active’ subset.
-The reason for using only this subset is to ensure that information about the $R^{A}$ values of individuals in fold $k$ provided by the value of $\hat{\pi}^{B}_{k}$ vanishes asymptotically.
-We now describe the procedure for choosing the active subset, first for the case where $L=1$ (simple random sampling without replacement) and then for the case where $L>1$.
-If $L=1$, then at most $\lceil M/K\rceil$ of the $M$ sampled clusters are in fold $k$ (here, $\lceil x\rceil$ denotes the smallest integer that is greater than or equal to $x$).
-Hence, at least $M-\lceil M/K\rceil$ sampled clusters must be in the remaining folds.
-The active subset is chosen by randomly subsampling $M-\lceil M/K\rceil$ of the $M-M_{k}$ sampled clusters not in fold $k$.
-To compensate for the subsampling, we multiply the $\pi^{A}$ values of individuals not in fold $k$ by $(M-\lceil M/K\rceil)/(M-M/K)$ when calculating $\hat{\pi}^{B}_{k}$.
-Now consider the case where $L>1$. Choose a number $0<\delta<1$ and define
+As in Section 4, $\hat{m}_{k}$ is calculated using the $Y$ and $X$ values of all individuals with $R^{B}=1$ not in fold $k$, and $\hat{\pi}^{B}_{k}$ uses the data $\{(R^{B}_{i},R^{B}_{i}X_{i}):i\notin\mathcal{S}_{k}\}$ on all individuals not in fold $k$.
+However, unlike in Section 4, where $\hat{\pi}^{B}_{k}$ used the data $\{(R^{A}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i}):i\notin\mathcal{S}_{k}\}$ on all individuals not in fold $k$, here we discard some of these data when calculating $\hat{\pi}^{B}_{k}$.
+Specifically, we only use data $(R^{A},R^{A}\pi^{A},R^{A}X)$ on those individuals in Sample A who belong to a subset of the clusters not in fold $k$.
+We refer to this subset of clusters as the ‘active’ subset.
+In effect, we temporarily set $R^{A}=R^{A}X=0$ for any individual in Sample A and not in fold $k$ who does not belong to an active cluster while calculating $\hat{\pi}^{B}_{k}$.
+As explained below, the reason for only using this subset is to ensure that information about the $R^{A}$ values of individuals in fold $k$ provided by the value of $\hat{\pi}^{B}_{k}$ vanishes asymptotically.
+To choose the active subset, first choose a number $0<\delta<1$ and define
 $C^{(l)}=\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor\wedge(M^{(l)}-M^{(l)}_{k})\hskip 28.45274pt(l=1,\ldots,L)$ (5)
-($a\wedge b$ denotes the minimum of $a$ and $b$).
-Choose the active subset by, for each $l=1,\ldots,L$, randomly subsampling $C^{(l)}$ of the $M^{(l)}-M^{(l)}_{k}$ sampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ not in fold $k$ To compensate for the subsampling, multiply the $\pi^{A}$ values of individuals with $\pi^{C}=\pi^{C(l)}$ not in fold $k$ by
+(here, $a\wedge b$ denotes the minimum of $a$ and $b$).
+For each $l=1,\ldots,L$, randomly subsample $C^{(l)}$ of the $M^{(l)}-M^{(l)}_{k}$ sampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ not in fold $k$.
+The set of $\sum_{l=1}^{L}M^{(l)}-M^{(l)}_{k}$ subsampled clusters is the active subset.
+To compensate for the subsampling, we temporarily multiply the $\pi^{A}$ values of individuals with $\pi^{C}=\pi^{C(l)}$ not in fold $k$ by
 $\frac{\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor}{\pi^{C(l)}(J^{(l)}-J^{(l)}_{k})}$ (6)
 when calculating $\hat{\pi}^{B}_{k}$.
-The motivation for calculating $\hat{\pi}^{B}_{k}$ using the $R^{A}$ values only of individuals in the active clusters is that when $M$ is large, $C^{(l)}$ is very likely to equal $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor$.
-If $C^{(l)}$ were guaranteed to equal $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor$, then knowing $\hat{\pi}^{B}_{k}$ would tell us nothing about the $(R^{A},R^{B},Y)$ values of individuals in fold $k$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$.
-Since, when $M$ is large, $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor\approx(M^{(l)}-M^{(l)}_{k})(1-\delta)$, the price we pay to achieve this noninformativeness of $\hat{\pi}^{B}_{k}$ when $M$ is large is the discarding of about $100\delta$% of the individuals with $R^{A}_{i}=1$ when calculating $\hat{\pi}^{B}_{k}$.
+The motivation for calculating $\hat{\pi}^{B}_{k}$ using the $(R^{A},R^{A}X)$ values only of individuals in the active clusters is that when $M$ is large, $C^{(l)}$ is very likely to equal $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor$.
+If $C^{(l)}$ were guaranteed to equal $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor$ for all $l$, then knowing $\hat{\pi}^{B}_{k}$ would tell us nothing about the $(R^{A},R^{B},Y)$ values of individuals in fold $k$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$.
+Hence, information about the $R^{A}$ values of individuals in fold $k$ provided by the value of $\hat{\pi}^{B}_{k}$ vanishes asymptotically.
+When $M$ is large, $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor\approx(M^{(l)}-M^{(l)}_{k})(1-\delta)$.
+So, the price we pay to achieve this noninformativeness of $\hat{\pi}^{B}_{k}$ when $M$ is large is the discarding of about $100\delta$% of the individuals with $R^{A}_{i}=1$ when calculating $\hat{\pi}^{B}_{k}$.
 This motivates us to choose $\delta$ close to zero, e.g.
 $\delta=0.01$.
-In Appendix A2 we prove that when the folds and active subset are chosen in the way described above, the empirical process and remainder terms in the von Mises expansion (equation (3)) are $o_{p}(M^{-1/2})$, as required, provided that Condition C1 is satisfied.
+In Appendix A2 we prove that when the folds and active subset are chosen in the way described above, the empirical process and remainder terms in the von Mises expansion (equation (3)) are $o_{p}(M^{-1/2})$, provided that Condition C1 is satisfied.
+In the special case of $L=1$ (SRSWOR), we can instead choose the active subset to be a random subset of $M-\lceil M/K\rceil$ of the $M-M_{k}$ sampled clusters not in fold $k$ ($\lceil x\rceil$ denotes the smallest integer greater than or equal to $x$).
+This ensures noninformativeness of $\hat{\pi}^{B}_{k}$, because there must be at least $M-\lceil M/K\rceil$ sampled clusters not in fold $k$, and ensures that at most K1K-1 sampled clusters are excluded from the active subset.
+To compensate for the subsampling, we temporarily multiply the $\pi^{A}$ values of individuals not in fold $k$ by $(M-\lceil M/K\rceil)/(M-M/K)$ when calculating $\hat{\pi}^{B}_{k}$.
+Clearly, if $M$ is an integer multiple of $K$, then all sampled clusters not in fold $k$ are in the active subset, i.e.
+there is no subsampling.
 In practice, the number of distinct cluster sampling probabilities will often not be small.
@@
 $\hat{n}^{B}$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\frac{R^{B}_{i}}{\hat{\pi}^{B}(X_{i})}.$
-This estimator $\hat{\theta}_{{\rm CLW}2}$ may be more efficient than $\hat{\theta}_{1}$, for the same reason that the Hajek estimator may be more efficient than the Horwitz-Thompson estimator.
+This may be more efficient than $\hat{\theta}_{1}$, for the same reason that the Hajek estimator may be more efficient than the Horwitz-Thompson estimator.
 In CLW’s simulation study with parametric nuisance models, $\hat{\theta}_{{\rm CLW}2}$ did indeed have smaller variance than $\hat{\theta}_{1}$.
@@
 and $\hat{\theta}_{2}=\hat{\theta}_{{\rm CLW}2}+o_{p}(M^{-1/2})$, i.e. $\hat{\theta}_{2}$ is asymptotically equivalent to $\hat{\theta}_{{\rm CLW}2}$.
-The mean-zero term $\sqrt{M}n^{-1}\sum_{i=1}^{n}\{m_{0}(X_{i})-Y_{i}\}$ in expression (7) is $O_{p}(1)$ and does not vanish asymptotically.
-However, when the population is large compared to Samples A and B, the variance of this term is small compared to the variance of the other terms in expression (7).
+The mean-zero term $\sqrt{M}n^{-1}\sum_{i=1}^{n}\{m_{0}(X_{i})-Y_{i}\}$ in equation (7) is $O_{p}(1)$ and does not vanish asymptotically.
+However, when the population is large compared to Samples A and B, its variance is small compared to that of the other terms in expression (7).
 We see that if this term is ignored, equation (7) is the same as equation (4) but with $m_{0}(X_{i})$ in the first term replaced by $m_{0}(X_{i})-n^{-1}\sum_{i^{\prime}=1}^{n}m_{0}(X_{i^{\prime}})$.
-Ignoring the term $\sqrt{M}n^{-1}\sum_{i=1}^{n}\{m_{0}(X_{i})-Y_{i}\}$ in equation (7), the asymptotic variance of $\hat{\theta}_{{\rm CLW}2}-\bar{Y}$ (and hence of $\hat{\theta}_{2}-\bar{Y}$) is given by CLW, along with an estimator of this variance.
+The asymptotic variance of $\hat{\theta}_{{\rm CLW}2}-\bar{Y}$ (and hence of $\hat{\theta}_{2}-\bar{Y}$) is given by CLW (who also ignored the $\sqrt{M}n^{-1}\sum_{i=1}^{n}\{m_{0}(X_{i})-Y_{i}\}$ term), along with an estimator of this variance.
 This is described in more detail by Seaman et al.
@@
 A popular alternative to estimator $\hat{\theta}_{\rm iid}$ for iid data is the targeted maximum likelihood estimator (TMLE).
-The TMLE method involves modifying initial estimate $\hat{m}_{\rm iid}$ of $E(Y\mid X)$ to $\hat{m}^{*}_{\rm iid}$ in such a way that the simple regression imputation estimator $n^{-1}\sum_{i=1}^{n}\hat{m}_{\rm iid}^{*}(X_{i})$ of $E(Y)$ can be used.
+The TMLE method involves modifying an initial estimate $\hat{m}_{\rm iid}$ of $E(Y\mid X)$ to $\hat{m}^{*}_{\rm iid}$ in such a way that the simple regression imputation estimator $n^{-1}\sum_{i=1}^{n}\hat{m}_{\rm iid}^{*}(X_{i})$ of $E(Y)$ can be validly used even if $\hat{m}_{\rm iid}$ was obtained using a data-adaptive method.
 We now describe an analogous TMLE method for our survey sample situation.
@@
 $\hat{m}_{k}(x;\epsilon_{k})=\hat{m}_{k}(x)+\epsilon_{k}\;\frac{1}{\hat{\pi}^{B}_{k}(X_{i})}$ (8)
-for some $\epsilon_{k}$. Alternatively, if $Y$ is bounded by zero and one, i.e. $P(0<Y<1)=1$, then $\hat{m}_{k}(x;\epsilon_{k})$ can be defined by equation (8) or equation (9):
+for some $\epsilon_{k}$. Alternatively, if $Y$ is bounded by zero and one, i.e. $P(0<Y<1)=1$, then $\hat{m}_{k}(x;\epsilon_{k})$ can be defined by either equation (8) or equation (9):
 $\hat{m}_{k}(x;\epsilon_{k})=\frac{\exp\{\mbox{logit }\hat{m}_{k}(x)+\epsilon_{k}/\hat{\pi}^{B}_{k}(X_{i})\}}{1+\exp\{\mbox{logit }\hat{m}_{k}(x)+\epsilon_{k}/\hat{\pi}^{B}_{k}(X_{i})\}}.$ (9)
@@
 8 Calculating $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$
-So far, we have been agnostic about how to calculate $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$, except to specify that they use, respectively, $(X,R^{A},R^{B})$ values and $(X,Y)$ values of individuals not in fold $k$.
+So far, we have been agnostic about how to calculate $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$, except to specify that they only use, respectively, data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i}):i\notin\mathcal{S}_{k}\}$ and data $\{(R^{B}_{i},R^{B}_{i}X_{i},R^{B}_{i}Y_{i}):i\notin\mathcal{S}_{k}\}$ not in fold $k$ (and, when calculating $\hat{\pi}^{B}$, setting $R^{A}=R^{A}X=0$ for individuals not in active clusters and compensating for this by scaling the $\pi^{A}_{i}$’s).
 Estimating $m_{0}(X)=E(Y\mid X)$ is relatively straightforward, because $Y_{1},\ldots,Y_{n}$ are assumed to be independent given $X_{1},\ldots,X_{n}$.
 This is a standard problem for which many data-adaptive methods could be used.
+In the simulation study of Section 9 we use both XGBoost and Highly Adaptive LASSO (HAL).
 Estimating $\pi^{B}_{0}(X)=P(R^{B}=1\mid X)$ is more complicated, because, although $R^{B}_{1},\ldots,R^{B}_{n}$ are assumed independent given $X_{1},\ldots,X_{n}$, we only observe $X$ for individuals with $R^{A}=1$ or $R^{B}=1$.
@@
 $\sum_{i=1}^{n}R^{B}_{i}\;\log\pi^{B}(X_{i})-\left(\frac{R^{A}_{i}}{\pi^{A}}-R^{B}_{i}\right)\log\{1-\pi^{B}(X_{i})\}.$ (11)
-This is the log likelihood for a binomial regression model of $R^{B}$ on $X$ in the population of $n$ individuals, using the individuals in Sample A weighted by their inverse sampling probabilities to represent the population, so that the ‘failures’ or ‘non-events’ are represented by the weighted Sample A minus the individuals in Sample B.
+This is the log likelihood for a binomial regression model of $R^{B}$ on $X$ in the population of $n$ individuals, using the individuals in Sample A weighted by their inverse sampling probabilities to represent the population.
 Ferri-Garcia et al.
 (2022)[16] (see also [11]) use the pseudo-likelihood together with a classification and regression tree to estimate $\pi^{B}_{0}(X)$.
-The approximate log pseudo-likelihood is the same as expression (11) but omitting the $R^{B}_{i}\{1-\pi^{B}(X_{i})\}$ term.
+In Section 9, we use expression (11) as an objective function (to be maximised) for XGBoost.
+The approximate log pseudo-likelihood is the same as expression (11) but omitting the $R^{B}_{i}\log\{1-\pi^{B}(X_{i})\}$ term.
 When Sample B is small compared to the population, this term will be negligible compared to the other terms in expression (11).
-The approximate pseudo-likelihood has the advantage that it can be used quite generally with data-adaptive methods that predict a binary outcome and that allow for weights.
+The approximate pseudo-likelihood has the advantage that it can be used quite generally with data-adaptive methods that predict a binary outcome and allow for weights.
+In Section 9, we use the approximate pseudo-likelihood with HAL.
 We acknowledge CLW’s criticism of the approximate pseudo-likelihood in situations where Sample B is not a small fraction of the population, but we focus in Section 9 on scenarios where Sample B is small compared to the population, a situation which would be quite common in practice.
@@
 Hence, the expected number of individuals in a cluster is 600, and the expected population size is 600,000.
-For each individual $i$ in cluster $j$, continuous variables $X_{1i}$ and $X_{2i}$ and binary variables $X_{3i}$ and $X_{4i}$ were generated independently, with expectations that depend on the total number of households in the cluster $j$, the size $q$ of the household to which individual $i$ belongs, and a cluster-level random effect.
+For each individual $i$ in cluster $j$, continuous variables $X_{1i}$ and $X_{2i}$ and binary variables $X_{3i}$ and $X_{4i}$ were generated independently, with expectations that depend on the total number of households in the cluster $j$, the size (1, 2 or 3) of the household to which individual $i$ belongs, and a cluster-level random effect.
+See Appendix A9 for details.
 This fixed population was used for all the simulations.
-For each simulated dataset, $Y_{1},\ldots,Y_{n}$ were generated independently from a normal distribution with mean that is a function of $X_{i}=(X_{1i},X_{2i},X_{3i},X_{4i})$ and with variance 1.
-Sample A was drawn by sampling $M$ clusters with replacement from the $J$ clusters using Sampford sampling, with the probability of sampling cluster $j$ being proportional to the total number of households in cluster $j$.
-From each of these $M$ sampled clusters, $n_{\rm house}$ households were sampled using simple random sampling without replacement.
+For each simulated dataset, $Y_{1},\ldots,Y_{n}$ were generated independently from a normal distribution with mean that depends on $X=(X_{1},X_{2},X_{3},X_{4})$ and variance 1.
+Sample A was drawn using a three-stage sampling design.
+First $M$ clusters were selected using Sampford sampling, with the probability of selecting cluster $j$ being proportional to the total number of households in cluster $j$.
+From each of these $M$ selected clusters, $n_{\rm house}$ households were selected using SRSWOR.
 Then one individual was sampled at random from each of these $n_{\rm house}$ sampled households.
 Thus, Sample A contained $M\times n_{\rm house}$ individuals.
-The probability of sampling individual $i$ for Sample B was a function of $X_{i}$.
-We considered six scenarios.
-In Scenarios 1 and 2, $m_{0}(X)$ is linear in $X$ and $\pi^{B}_{0}(X)$ follows a logistic regression model with only main effects for $X_{1}$, $X_{2}$, $X_{3}$ and $X_{4}$.
+We considered 24 scenarios.
+In Scenarios 1–12, $m_{0}(X)$ is linear in $X$ and $\pi^{B}_{0}(X)$ follows a logistic regression model with only main effects.
 Specifically, $m_{0}(X)=X_{1}+X_{2}+2X_{3}+X_{4}$ and $\mbox{logit }\pi^{B}_{0}(X)=\alpha_{\rm int}+0.5X_{1}+X_{2}+0.5X_{3}+X_{4}$.
-Both these scenarios use $M=150$ and $n_{\rm house}=20$.
-They differ in the value of $\alpha_{\rm int}$, and hence the expected size of Sample B.
-In Scenario 1, $\alpha_{\rm int}=-6.2$ and the expected size of Sample B is 7000; in Scenario 2, $\alpha_{\rm int}=-7.5$ and the expected size of Sample B is 2000.
-In Scenarios 3–6, $m_{0}(X)$ is a non-linear function of $X$ and the logistic regression model for $\pi^{B}_{0}(X)$ includes an interaction and a quadratic term.
-Specifically, $m_{0}(X)=0.5X_{1}+0.5X_{2}+2X_{3}+X_{4}+2X_{1}X_{3}+X_{2}^{2}$ and $\mbox{logit }\pi^{B}_{0}(X)=-6.4+0.25X_{1}+0.5X_{2}+0.5X_{3}+X_{4}+X_{1}X_{3}+0.5X_{2}^{2}$.
-The expected size of Sample B is 7500.
-These four scenarios differ in the values of $M$ and $n_{\rm house}$.
-Specifically, Scenarios 3–6 use, respectively, $(M,n_{\rm house})=(150,20)$, $(50,20)$, $(150,5)$ and $(50,5)$.
-The following estimators were applied to each of 100 simulated datasets for each of the six scenarios.
+In Scenarios 13–24, $m_{0}(X)$ and $\pi^{B}_{0}(X)$ each include an interaction and a quadratic term.
+Specifically, $m_{0}(X)=0.5X_{1}+0.5X_{2}+2X_{3}+X_{4}+2X_{1}X_{3}+X_{2}^{2}$ and $\mbox{logit }\pi^{B}_{0}(X)=\alpha_{\rm int}+0.25X_{1}+0.5X_{2}+0.5X_{3}+X_{4}+X_{1}X_{3}+0.5X_{2}^{2}$.
+The value of $\alpha_{\rm int}$ was chosen to achieve a desired expected size $E(n^{B})$ of Sample B.
+Each of these two sets of 12 scenarios corresponded to two choices of $M$ (150 and 50), two choices of $n_{\rm house}$ (20 and 5), and three choices of $E(n^{B})$ (5000, 1000 and 500).
+The following estimators were applied to each of 400 simulated datasets for each of the 24 scenarios.
 HT
@@
 : $\hat{\theta}_{{\rm CLW}2}$ using parametric nuisance models
-DR2
-: $\hat{\theta}_{2}$ using parametric nuisance models
 TMLE1
@@
 DR1.hal5
-: $\hat{\theta}_{1}$ using Highly Adaptive LASSO (HAL) to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
+: $\hat{\theta}_{1}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting
 DR2.hal5
-: $\hat{\theta}_{2}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
+: $\hat{\theta}_{2}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting
 TMLE1.hal5
-: $\hat{\theta}_{\rm TMLE1}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
+: $\hat{\theta}_{\rm TMLE1}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting
 TMLE2.hal5
-: $\hat{\theta}_{\rm TMLE2}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
+: $\hat{\theta}_{\rm TMLE2}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting
 Of these, DR1, DR2clw and KH were described by CLW.
-For the methods using HAL, we set $\delta=0.01$ and used $L=4$ sets of clusters based on quartiles of the distribution of $\pi^{C}_{j}$ (as described at the end of Section 5).
-In addition, we calculated the estimators using generalised boosting models (GBM), in place of HAL, to estimate the nuisance models.
-These will be denoted ‘DR1.gbm5’ and so on.
-We also applied the estimators that used HAL or GBM without cross-fitting.
-These will be denoted ‘DR1.hal1’, ‘DR1.gbm1’, and so on.
-Results are shown in Tables 1, 2, 3, 4, 5 and 6 for Scenarios 1–6, respectively.
-In Scenarios 1 and 2, the methods using parametric nuisance models, viz.
-DR1, DR2clw, DR2, TMLE1, TMLE2 and KH, yield approximately unbiased point estimates and standard error estimates, and the coverage of 95% confidence intervals is correct.
-The estimators DR2clw and TMLE2 are more efficient than the DR1 and TMLE1 estimators, and DR2 is, as expected, as efficient as DR2clw.
-Using HAL to estimate the nuisance functions provides no benefit in these scenarios, where the parametric nuisance models are correctly specified, but also no loss of efficiency or coverage.
-Using GBM leads to some bias and a small deterioriation in coverage, but bias is still small and coverage not far from its nominal level.
-In Scenarios 3–6, the methods using parametric nuisance models yield biased estimates and confidence intervals with poor coverage.
-When HAL is used, the bias is small and the coverage close to its nominal level.
-Results from GBM are similar, although the bias can be slightly greater than for HAL.
-Using cross-fitting with five folds appears to provide no performance improvement over not using cross-fitting in these scenarios but also no loss of efficiency.
-bias empSE SEhat cover
-HT 0.002 0.046 0.046 94
-Haj 0.000 0.037 0.037 95
-naive -0.288 0.035
-DR1 0.001 0.044 0.044 94
-DR2clw -0.000 0.035 0.035 95
-DR2 -0.000 0.035 0.035 95
-TMLE1 0.001 0.044 0.044 94
-TMLE2 -0.000 0.035 0.035 95
-KH 0.001 0.044 0.044 94
-DR1.hal5 0.002 0.044 0.044 93
-DR2.hal5 0.001 0.035 0.034 94
-TMLE1.hal5 0.002 0.044 0.044 93
-TMLE2.hal5 0.000 0.035 0.034 94
-DR1.hal1 0.002 0.044 0.044 93
-DR2.hal1 0.000 0.034 0.034 94
-TMLE1.hal1 0.002 0.044 0.044 93
-TMLE2.hal1 0.000 0.034 0.034 94
-DR1.gbm5 0.012 0.044 0.043 92
-DR2.gbm5 0.010 0.034 0.033 91
-TMLE1.gbm5 0.009 0.044 0.043 93
-TMLE2.gbm5 0.008 0.034 0.033 92
-DR1.gbm1 0.010 0.044 0.043 92
-DR2.gbm1 0.009 0.034 0.033 92
-TMLE1.gbm1 0.007 0.044 0.043 92
-TMLE2.gbm1 0.006 0.034 0.033 92
-Table 1: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 1.
-bias empSE SEhat cover
-HT 0.002 0.046 0.046 94
-Haj 0.000 0.037 0.037 95
-naive -0.288 0.035
-DR1 0.001 0.052 0.051 96
-DR2clw -0.001 0.044 0.043 95
-DR2 -0.001 0.044 0.043 95
-TMLE1 0.001 0.052 0.051 96
-TMLE2 -0.001 0.044 0.043 95
-KH 0.001 0.052 0.051 96
-DR1.hal5 0.002 0.052 0.049 95
-DR2.hal5 0.001 0.044 0.041 94
-TMLE1.hal5 0.002 0.052 0.050 95
-TMLE2.hal5 0.000 0.044 0.042 94
-DR1.hal1 0.002 0.052 0.050 95
-DR2.hal1 0.001 0.044 0.041 94
-TMLE1.hal1 0.001 0.052 0.050 95
-TMLE2.hal1 0.000 0.044 0.042 94
-DR1.gbm5 0.021 0.058 0.053 91
-DR2.gbm5 0.019 0.051 0.045 90
-TMLE1.gbm5 0.020 0.052 0.054 92
-TMLE2.gbm5 0.019 0.044 0.045 91
-DR1.gbm1 0.017 0.052 0.048 92
-DR2.gbm1 0.016 0.043 0.039 90
-TMLE1.gbm1 0.013 0.052 0.049 94
-TMLE2.gbm1 0.011 0.044 0.040 92
-Table 2: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 2.
-bias empSE SEhat cover
-HT 0.001 0.062 0.063 94
-Haj -0.001 0.053 0.054 95
-naive -0.399 0.047
-DR1 0.222 0.068 0.078 16
-DR2clw 0.184 0.053 0.069 18
-DR2 0.220 0.058 0.069 8
-TMLE1 0.125 0.064 0.070 58
-TMLE2 0.124 0.053 0.060 46
-KH 0.089 0.062 0.063 72
-DR1.hal5 0.002 0.063 0.062 94
-DR2.hal5 0.001 0.053 0.053 94
-TMLE1.hal5 0.002 0.063 0.066 96
-TMLE2.hal5 0.001 0.053 0.053 94
-DR1.hal1 0.002 0.062 0.062 94
-DR2.hal1 0.001 0.053 0.053 94
-TMLE1.hal1 0.002 0.062 0.067 96
-TMLE2.hal1 0.001 0.053 0.053 94
-DR1.gbm5 0.020 0.063 0.062 94
-DR2.gbm5 0.018 0.053 0.052 92
-TMLE1.gbm5 0.016 0.063 0.066 96
-TMLE2.gbm5 0.015 0.053 0.052 92
-DR1.gbm1 0.019 0.063 0.062 94
-DR2.gbm1 0.018 0.053 0.052 92
-TMLE1.gbm1 0.016 0.063 0.066 95
-TMLE2.gbm1 0.014 0.053 0.052 92
-Table 3: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 3.
-bias empSE SEhat cover
-HT -0.002 0.109 0.110 94
-Haj -0.001 0.091 0.094 95
-naive -0.399 0.081
-DR1 0.238 0.119 0.126 56
-DR2clw 0.190 0.084 0.110 60
-DR2 0.239 0.099 0.110 42
-TMLE1 0.118 0.110 0.116 86
-TMLE2 0.119 0.090 0.098 82
-KH 0.085 0.108 0.106 87
-DR1.hal5 0.000 0.109 0.106 93
-DR2.hal5 0.001 0.090 0.089 95
-TMLE1.hal5 -0.000 0.108 0.113 94
-TMLE2.hal5 0.001 0.090 0.089 95
-DR1.hal1 0.000 0.108 0.106 93
-DR2.hal1 0.001 0.090 0.089 95
-TMLE1.hal1 -0.000 0.108 0.113 94
-TMLE2.hal1 0.001 0.090 0.089 95
-DR1.gbm5 0.024 0.109 0.105 93
-DR2.gbm5 0.025 0.090 0.088 93
-TMLE1.gbm5 0.019 0.109 0.113 95
-TMLE2.gbm5 0.020 0.090 0.088 94
-DR1.gbm1 0.020 0.109 0.105 94
-DR2.gbm1 0.021 0.090 0.088 94
-TMLE1.gbm1 0.016 0.109 0.113 96
-TMLE2.gbm1 0.016 0.090 0.088 94
-Table 4: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 4.
-bias empSE SEhat cover
-HT -0.011 0.111 0.112 94
-Haj -0.013 0.092 0.093 94
-naive -0.406 0.081
-DR1 0.239 0.129 0.123 52
-DR2clw 0.186 0.090 0.104 57
-DR2 0.237 0.111 0.104 38
-TMLE1 0.117 0.108 0.115 85
-TMLE2 0.115 0.087 0.094 79
-KH 0.082 0.105 0.107 90
-DR1.hal5 -0.003 0.105 0.106 95
-DR2.hal5 -0.005 0.086 0.086 94
-TMLE1.hal5 -0.003 0.105 0.113 96
-TMLE2.hal5 -0.005 0.086 0.086 94
-DR1.hal1 -0.003 0.105 0.106 95
-DR2.hal1 -0.005 0.086 0.086 94
-TMLE1.hal1 -0.003 0.105 0.113 96
-TMLE2.hal1 -0.005 0.086 0.086 94
-DR1.gbm5 0.022 0.106 0.106 96
-DR2.gbm5 0.021 0.087 0.085 94
-TMLE1.gbm5 0.018 0.106 0.112 97
-TMLE2.gbm5 0.016 0.087 0.086 94
-DR1.gbm1 0.019 0.105 0.106 96
-DR2.gbm1 0.017 0.086 0.085 94
-TMLE1.gbm1 0.015 0.106 0.112 97
-TMLE2.gbm1 0.013 0.087 0.086 94
-Table 5: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 5.
-bias empSE SEhat cover
-HT 0.008 0.195 0.196 95
-Haj 0.004 0.164 0.164 95
-naive -0.394 0.142
-DR1 0.329 0.262 0.212 68
-DR2clw 0.224 0.147 0.178 80
-DR2 0.325 0.246 0.178 58
-TMLE1 0.123 0.186 0.202 94
-TMLE2 0.119 0.149 0.165 91
-KH 0.095 0.178 0.184 94
-DR1.hal5 0.010 0.180 0.184 95
-DR2.hal5 0.007 0.146 0.149 96
-TMLE1.hal5 0.011 0.180 0.196 97
-TMLE2.hal5 0.007 0.145 0.149 96
-DR1.hal1 0.011 0.180 0.184 95
-DR2.hal1 0.007 0.145 0.149 96
-TMLE1.hal1 0.010 0.180 0.196 97
-TMLE2.hal1 0.006 0.146 0.149 96
-DR1.gbm5 0.035 0.180 0.184 94
-DR2.gbm5 0.031 0.146 0.149 95
-TMLE1.gbm5 0.036 0.181 0.196 96
-TMLE2.gbm5 0.032 0.146 0.149 94
-DR1.gbm1 0.032 0.181 0.184 95
-DR2.gbm1 0.029 0.146 0.149 95
-TMLE1.gbm1 0.032 0.181 0.196 96
-TMLE2.gbm1 0.028 0.146 0.149 95
-Table 6: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 6.
+The parametric nuisance models we used were linear regression (for $m_{0}$) and logistic regression (for $\pi^{B}_{0}$) models with only main effects for $X_{1}$, $X_{2}$, $X_{3}$ and $X_{4}$.
+Hence, these models are correctly specified in Scenarios 1–12 and misspecified in Scenarios 13–24.
+For DR1, DR2clw, TMLE1 and TMLE2, we estimated the standard error (SE) and calculated 95% confidence intervals (CIs) using variance estimators that are valid when the model for $\pi^{B}_{0}$ (but not necessarily $m_{0}$) is correctly specified[26].
+In addition, we calculated $\hat{\theta}_{1}$, $\hat{\theta}_{2}$, $\hat{\theta}_{\rm TMLE1}$ and $\hat{\theta}_{\rm TMLE2}$ using XGBoost, instead of HAL, to estimate $\hat{\pi}^{B}$ and $\hat{m}$.
+These will be denoted ‘DR1.xgb5’ and so on.
+For methods using cross-fitting, we used $K=5$ folds, set $\delta=0.01$ and used $L=4$ sets of clusters based on quartiles of the distribution of $\pi^{C}_{j}$ (as described at the end of Section 5).
+We also applied the estimators that used HAL or XGBoost without cross-fitting (denoted ‘DR1.hal1’, ‘DR1.xgb1’, and so on).
+Full results are given in Appendix A10.
+Here, we highlight the main findings.
+In Scenarios 1–4, where $E(n^{B})=5000$, all estimators except Naive were approximately unbiased with CIs that had approximately correct coverage.
+The estimators that used HAL or XGBoost with cross-fitting to estimate the nuisance functions had very similar SE to the corresponding estimators that used parametric nuisance models (see Figure 1).
+Figure 1: Empirical standard error of $\hat{\theta}_{1}$ when $\pi^{B}_{0}$ and $m_{0}$ are estimated using parametric models (DR1), HAL (DR1.hal5) or XGBoost (DR1.xgb5), for Scenarios 1–12.
+The Hajek, DR2 and TMLE2 estimators were more efficient than the corresponding HT, DR1 and TMLE1 estimators; the SEs of the former were approximately half that of the former.
+Findings for Scenarios 5–8, where $E(n^{B})=1000$, were similar, although CI coverage of DR2.xgb5 and TMLE2.xgb5 dropped to 89% in Scenario 6, where $M=150$ and $n_{\rm house}=5$.
+Coverage for DR2.hal5, TMLE2.hal5, DR2.xgb5 and TMLE2.xgb5 dropped further in Scenarios 9–12, where $E(n^{B})=500$, but usually remained above 90% and was never less than 87%.
+In Scenarios 13–24, the estimators using parametric nuisance models were biased and had CIs with undercoverage (see Figure 2).
+Figure 2: Left: Coverage of 95% confidence interval for $\bar{Y}$ when $\hat{\theta}_{1}$ used with parametric nuisance models (DR1), HAL (DR1.hal5), XGBoost (DR1.xgb5) or Kim and Haziza’s method (KH), for Scenarios 13–24.
+Right: Coverage when $\hat{\theta}_{{\rm CLW}2}$ used with parametric nuisance models (DR2clw) or $\hat{\theta}_{2}$ used with HAL (DR2.hal5) or XGBoost (DR2.xgb5).
+Coverage was particularly bad (in one case, as low as 11%) for the DR2clw and TMLE2 estimators.
+Estimators using HAL or XGBoost had much smaller bias, especially in Scenarios 13–16, where $E(n^{B})=5000$; here bias was almost zero.
+CI coverage for these estimators was approximately 95% and never dropped below 90%.
+Using HAL or XGBoost without cross-fitting appeared to have little effect on bias and virtually no effect on the empirical SE.
+For XGBoost (but less obviously for HAL), omitting the cross-fitting step caused CI coverage to deteriorate in many of the scenarios, especially for DR2 and TMLE2 (see Figure 3).
+This lack of obvious effect on bias and this deterioration of coverage is in line with the findings of Ellul et al.
+(2025)[15].
+Figure 3: Coverage of 95% confidence interval for $\bar{Y}$ when $\hat{\theta}_{1}$ (top) or $\hat{\theta}_{2}$ (bottom) is used with HAL (left) or XGBoost (right).
+Each plot compares coverage when 5-fold cross-fitting is used versus when no cross-fitting is used.
+Each dot corresponds to one of the 24 scenarios.
+Broken lines indicate perfect 95% coverage.
+Finally, we also calculated $\hat{\theta}_{2}$ using parametric nuisance models.
+As expected, the biases and SEs of this estimator were very similar to those of DR2clw.
+Over the 24 scenarios, the mean (maximum) absolute difference in their biases was 0.0011 (0.0036), and the ratio of their SEs varied from 0.99 to 1.01.
 10 Discussion
@@
 This method yields an asymptotically normally distributed estimator with a standard error that can estimated as easily as can the standard error of a Horwitz-Thompson estimator.
-To prove this, we have relied on cross-fitting and have (unless cluster sampling probabilities are all equal) discarded some information when estimating the nuisance function $\hat{\pi}^{B}$.
+To prove this, we have relied on cross-fitting.
 For the closely related problem of estimating $E(Y)$ for an infinite population using iid data, an alternative to cross-fitting is to rely on the Donsker condition[21].
@@
 Nevertheless, it would be interesting to establish whether the Donsker condition would suffice in the setting we study in this article.
-Some researchers have previously used DR estimators with data-adaptive estimation of nuisance functions, e.g.
+Some researchers have previously used DR estimators with data-adaptive estimation of nuisance functions for combining probability and non-probability survey data, e.g.
 [3] and [12].
@@
 Indeed [12] comment (page 276) that “there are hardly any results on the theoretical properties of variance estimators when more complex machine learning models and techniques are used.” Moreover, these researchers used data-adaptive estimators that do not necessarily satisfy the Donsker condition and without using cross-fitting, which is potentially problematic even in the simpler situation of estimating $E(Y)$ for an infinite population using iid data.
-To establish asymptotic properties of our estimators, we have assumed that the finite population is generated from a particular superpopulation model, so that it grows in a particular way.
-It would be interesting to investigate whether asymptotic properties can be established when the finite population is allowed to grow in another way.
-We also used subsampling of an ‘active subset’ of clusters when calculating $\hat{\pi}^{B}$.
+In general, we have used subsampling of an ‘active subset’ of clusters when calculating $\hat{\pi}^{B}$.
 This should have no effect on asymptotic efficiency, and in our simulation study we observed no loss of efficiency from doing it.
@@
 As with other debiased machine learning estimators using cross-fitting, repeated cross-fitting would also have the advantage of reducing the finite-sample Monte Carlo variation induced by the randomness of the process of partitioning units (in our case, clusters) among folds, although at the cost of increasing computation time.
+As discussed in Section 4, if SRSWOR is used to select the $M$ clusters for Sample A and $M$ is an integer multiple of the number of folds $K$, there is no need to subsample an active subset.
+Appendix A11 shows this is true also when the clusters are sampled using conditional Poisson sampling.
 The approach we have presented is for probability samples that do not use stratified sampling.
 It is straightforward in principle to generalise it to deal with stratified sampling.
-Now the assignment of clusters to the $K$ folds should be done within each stratum.
-This should be fairly unproblematic when the probability sample uses simple random sampling of clusters without replacement within each stratum or when the number of clusters sampled from each stratum is large.
+Now the assignment of clusters to the $K$ folds would be done within each stratum.
+This should be fairly unproblematic when the probability sample uses SRSWOR of clusters within each stratum or when the number of clusters sampled from each stratum is large.
 When, however, the cluster sampling probabilities within strata vary and the number of sampled clusters within strata is small, there may be two problems.
@@
 The scenario of sampling few clusters per stratum is a problem worthy of further research.
-Finally, several researchers have considered the problem of efficiently combining an IPW or DR estimator of the mean of $Y$ with a Horwitz-Thompson or Hajek estimator of this same mean based on Sample A alone.
-Gao and Yang [18] consider the DR estimator $\hat{\theta}_{1}$ and use parametric nuisance models.
+To establish asymptotic properties of our estimators, we have assumed that the finite population is generated from a particular superpopulation model, so that it grows in a particular way.
+It would be interesting to investigate whether asymptotic properties can be established when the finite population is allowed to grow in another way.
+Finally, several researchers have considered the problem of efficiently combining an IPW or DR estimator of the mean of $Y$ with a Horwitz-Thompson or Hajek estimator of this same mean based on Sample A alone when $Y$ is observed in both samples.
+Gao and Yang (2023)[19] consider the DR estimator $\hat{\theta}_{1}$ and use parametric nuisance models.
 Seaman et al.
-(2025) discuss this approach and extend it to $\hat{\theta}_{{\rm CLW}2}$ and IPW estimators, again using parametric nuisance models.
+(2025)[26] discuss this approach and extend it to $\hat{\theta}_{{\rm CLW}2}$ and IPW estimators, again using parametric nuisance models.
 Rueda et al.
@@
 Since, when Condition C1 is satisfied, the debiased machine learning estimators of $\bar{Y}$ that we have presented here have the same asymptotic distribution as the corresponding DR estimators of $\bar{Y}$ that use two correctly specified parametric nuisance models, the combining methods discussed by Seaman et al.
-(2025) are equally appropriate.
+(2025) should apply.
 Acknowledgements
@@
 Shaun R. Seaman
-A1 Proof that empirical process and remainder terms are $o_{p}(M^{-1/2})$ for simple random sampling of clusters without replacement
+A1 Proof that empirical process and remainder terms are $o_{p}(M^{-1/2})$ for SRSWOR of clusters
 A1.1 Empirical process term
@@
 Now consider the second variance in expression (A4).
-As we argued in Section 4, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ corresponds to simple random sampling of $M/K$ clusters without replacement from the $J/K$ clusters in $\mathcal{S}_{k}$ followed by the original second-stage sampling of individuals within clusters.
+As we argued in Section 4, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ corresponds to SRSWOR of $M/K$ clusters from the $J/K$ clusters in $\mathcal{S}_{k}$ followed by the original second-stage sampling of individuals within clusters.
 Hence, for any fixed function $\hat{m}_{k}$,
 ${\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$ (A6)
-is the variance of the Horwitz-Thompson estimator of the population mean of $\hat{m}_{k}(X)-m_{0}(X)$ in individuals in $\mathcal{S}_{k}$ given $\mathcal{F}_{J}$ when the clusters in $\mathcal{S}_{k}$ are selected by simple random sampling without replacement and individuals within clusters are sampled according to the original second-stage sampling procedure.
+is the variance of the Horwitz-Thompson estimator of the population mean of $\hat{m}_{k}(X)-m_{0}(X)$ in individuals in $\mathcal{S}_{k}$ given $\mathcal{F}_{J}$ when the clusters in $\mathcal{S}_{k}$ are selected by SRSWOR and individuals within clusters are sampled according to the original second-stage sampling procedure.
 Subject to some regularity conditions, this variance is $O_{p}(M^{-1})$ (Proposition 2 of Chauvet and Vallee, 2020[4]).
@@
 $\hskip 28.45274pt-\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\frac{E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})}{\pi^{A}_{i}}\;m_{0}(X_{i}).$ (A7)
-As argued in Section 4, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ corresponds to simple random sampling of $M/K$ clusters without replacement from the $J/K$ clusters in $\mathcal{S}_{k}$ followed by the original second-stage sampling of individuals within clusters.
+As argued in Section 4, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}$ corresponds to SRSWOR of $M/K$ clusters from the $J/K$ clusters in $\mathcal{S}_{k}$ followed by the original second-stage sampling of individuals within clusters.
 Hence, $E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})=\pi^{A}_{i}$.
@@
 So, if Condition C1 holds, then expression (A8) is $o_{p}(M^{-1/2})$, as required.
-A2 Proof that empirical process and remainder terms are $o_{p}(M^{-1/2})$ for random sampling of clusters without replacement with unequal probabilities
+A2 Proof that empirical process and remainder terms are $o_{p}(M^{-1/2})$ for SWOR of clusters with unequal probabilities
 A2.1 Empirical process term
-Consider the empirical process term for fold $k$.
-Equations (A2) and (A3) still apply when clusters are sampled without replacement with unequal probabilities.
-Consider the variance in expression (A3).
+Consider the empirical process term for fold $k$. Equations (A2) and (A3) still apply when clusters are selected by SWOR with unequal probabilities. Consider the variance in expression (A3).
 ${\rm Var}_{\mathcal{G}_{J}}\left[\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right]$
@@
 $\hskip 56.9055pt\times\left.\{\hat{m}_{k}(X_{i})-m_{0}(X_{i})\}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}\right$]. (A12)
-By a very similar argument to that used in Section 4, we see that, because of the way that fold $S_{k}$ has been chosen, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}$ corresponds to stratified simple random sampling of $M^{(l)}_{k}$ clusters without replacement from the stratum of $J^{(l)}_{k}$ clusters with the same value of $\pi^{C(l)}$ followed by the original second-stage sampling of individuals within clusters.
+By a very similar argument to that used in Section 4, we see that, because of the way that fold $S_{k}$ has been chosen, the distribution of $\{R_{i}^{A}:i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k},M^{(.)}_{k}$ corresponds to stratified SRSWOR of $M^{(l)}_{k}$ clusters from the stratum of $J^{(l)}_{k}$ clusters with the same value of $\pi^{C(l)}$ followed by the original second-stage sampling of individuals within clusters.
 The probability of sampling individual $i$ is the original probability $\pi^{A}_{i}$ multiplied by $M^{(l)}_{k}/(J^{(l)}_{k}\pi^{C(l)})$, because the fraction of clusters sampled from this stratum is $M^{(l)}_{k}/J^{(l)}_{k}$, rather than $\pi^{C(l)}$.
@@
 $\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}E_{\mathcal{G}_{J}}\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}.$
-Equation (A7) still applies when clusters are sampled without replacement with unequal probabilities. As before, because of the way that the folds have been chosen, and using equation (A14), we have
+Equation (A7) still applies when clusters are selected using SWOR with unequal probabilities. As before, because of the way that the folds have been chosen, and using equation (A14), we have
 $\frac{E(R^{A}_{i}\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k})}{\pi^{A}_{i}}$
@@
 The rest of expression (A21) is the same as expression (A8).
-Hence, we are back to the situation we have where sampling of clusters is by simple random sampling without replacement.
+Hence, we are back to the situation we have where clusters are selected by SRSWOR.
 Thus, under the conditions given there, the remainder term is $o_{p}(M^{-1/2})$, as required.
@@
 Hence, $\hat{\theta}_{\rm TMLE1}=\hat{\theta}_{1}+o_{p}(M^{1/2})$.
+A9 Data-generating mechanism for simulation study
+Let $H_{jq}\sim\mbox{NegativeBinomial}(\mbox{mean}=100,\mbox{variance}=400)$ be the number of households of $q$ individuals in cluster $j$ ($q=1,2,3$ and $j=1,\ldots,J$).
+Thus, the number of individuals in cluster $j$ is $N_{j}=\sum_{q=1}^{3}qH_{j}$, and the population size is $n=\sum_{j=1}^{J}N_{j}$, which has expectation $600J$ and standard deviation $\sqrt{2400J}$.
+We use $J=1000$, and so this expectation and standard deviation are 600000 and 1549, respectively.
+Let $H_{j}=\sum_{q=1}^{3}H_{jq}$ be the total number of households in cluster $j$.
+Let $H_{j}^{*}=H_{j}\times J\left/\sum_{j=1}^{J}H_{j}\right.+\mbox{Normal}(0,0.1^{2})$ be the ratio of the number of households in cluster $j$ to the mean number of households per cluster plus a cluster-specific random effect.
+Recall that $D_{i}$ denotes the index of the cluster to which individual $i$ belongs, and let $B_{i}$ denote the number of individuals in the household to which individual $i$ belongs divided by the maximum household size.
+So, $B=1/3$, 2/3 or 1.
+Let $\mu_{i}=-1.3+0.5H_{D_{i}}^{*}+0.5B_{i}$.
+Given $\mu_{1},\ldots,\mu_{n}$, the $4n$ random variables $X_{11},X_{21},X_{31},X_{41},\ldots,X_{1n},X_{2n},X_{3n},X_{4n}$ are independently distributed with
+$X_{1i}\mid\mu_{i}$ $\sim$ $\mbox{Normal}(\mu_{i},0.4^{2})$
+$X_{2i}\mid\mu_{i}$ $\sim$ $\mbox{Normal}(\mu_{i},0.2^{2})$
+$X_{3i}\mid\mu_{i}$ $\sim$ $\mbox{Bernoulli}\big(\mbox{expit}(-1.35+\mu_{i})\big)$
+$X_{4i}\mid\mu_{i}$ $\sim$ $\mbox{Bernoulli}\big(\mbox{expit}(2\mu_{i})\big)$
+Note that marginally (over $\mu_{i}$) $P(X_{3i}=1)=0.2$ and $P(X_{3i}=1)=0.5$.
+A10 Results for simulation study
+Table A1 shows the values of $E(n^{B})$, $M$ and $n_{\rm house}$ in the 24 scenarios.
+Scenarios $E(n^{B})$ $M$ $n_{\rm house}$
+1 13 5000 150 20
+2 14 5000 150 5
+3 15 5000 50 20
+4 16 5000 50 5
+5 17 1000 150 20
+6 18 1000 150 5
+7 19 1000 50 20
+8 20 1000 50 5
+9 21 500 150 20
+10 22 500 150 5
+11 23 500 50 20
+12 24 500 50 5
+Table A1: For each of the 24 scenarios in the simulation study, the expected number of individuals in Sample B ($E(n^{B})$), and the number of sampled clusters ($M$) and of sampled houses per sampled cluster ($n_{\rm house}$) in Sample A.
+Tables A2A25 shows the bias, empirical SE, mean estimated SE and coverage of 95% for each estimator in each of Scenarios 1–24.
+bias empSE SEhat cover
+HT 0.005 0.069 0.068 95
+Haj 0.001 0.035 0.035 94
+naive -0.306 0.034
+DR1 0.004 0.068 0.068 94
+DR2clw 0.001 0.035 0.035 95
+TMLE1 0.004 0.068 0.068 94
+TMLE2 0.001 0.035 0.035 95
+KH 0.004 0.068 0.068 94
+DR1.hal5 0.008 0.068 0.067 94
+DR2.hal5 0.004 0.035 0.034 95
+TMLE1.hal5 0.007 0.068 0.067 94
+TMLE2.hal5 0.004 0.035 0.034 95
+DR1.hal1 0.007 0.068 0.067 94
+DR2.hal1 0.004 0.035 0.034 94
+TMLE1.hal1 0.007 0.068 0.067 94
+TMLE2.hal1 0.004 0.035 0.034 94
+DR1.xgb5 0.011 0.069 0.067 95
+DR2.xgb5 0.007 0.035 0.033 92
+TMLE1.xgb5 0.009 0.069 0.067 94
+TMLE2.xgb5 0.006 0.035 0.034 93
+DR1.xgb1 0.011 0.068 0.067 95
+DR2.xgb1 0.008 0.035 0.033 92
+TMLE1.xgb1 0.010 0.068 0.067 95
+TMLE2.xgb1 0.006 0.035 0.033 94
+Table A2: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 1, where $E(n^{B})=5000$, $M=150$, $n_{\rm house}=20$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT -0.006 0.127 0.127 95
+Haj -0.009 0.064 0.064 96
+naive -0.313 0.061
+DR1 -0.001 0.120 0.122 94
+DR2clw -0.005 0.053 0.055 94
+TMLE1 -0.001 0.120 0.122 94
+TMLE2 -0.005 0.053 0.055 94
+KH -0.001 0.120 0.122 94
+DR1.hal5 0.005 0.120 0.121 94
+DR2.hal5 0.001 0.053 0.053 94
+TMLE1.hal5 0.004 0.120 0.121 94
+TMLE2.hal5 0.001 0.053 0.053 94
+DR1.hal1 0.004 0.120 0.121 94
+DR2.hal1 0.000 0.053 0.053 94
+TMLE1.hal1 0.003 0.120 0.121 94
+TMLE2.hal1 -0.000 0.053 0.053 94
+DR1.xgb5 0.008 0.120 0.122 95
+DR2.xgb5 0.004 0.053 0.053 93
+TMLE1.xgb5 0.007 0.120 0.122 95
+TMLE2.xgb5 0.003 0.053 0.053 94
+DR1.xgb1 0.008 0.120 0.122 94
+DR2.xgb1 0.004 0.053 0.053 93
+TMLE1.xgb1 0.006 0.120 0.122 94
+TMLE2.xgb1 0.003 0.053 0.053 93
+Table A3: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 2, where $E(n^{B})=5000$, $M=150$, $n_{\rm house}=5$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT -0.003 0.120 0.119 93
+Haj -0.000 0.060 0.062 94
+naive -0.308 0.058
+DR1 -0.004 0.118 0.115 93
+DR2clw -0.001 0.055 0.055 94
+TMLE1 -0.004 0.118 0.115 93
+TMLE2 -0.001 0.055 0.055 94
+KH -0.004 0.118 0.115 93
+DR1.hal5 0.001 0.118 0.114 93
+DR2.hal5 0.004 0.054 0.053 94
+TMLE1.hal5 0.001 0.117 0.114 93
+TMLE2.hal5 0.004 0.054 0.054 94
+DR1.hal1 0.001 0.118 0.114 93
+DR2.hal1 0.004 0.055 0.054 94
+TMLE1.hal1 0.000 0.118 0.114 93
+TMLE2.hal1 0.003 0.055 0.054 94
+DR1.xgb5 0.005 0.118 0.114 92
+DR2.xgb5 0.008 0.055 0.053 93
+TMLE1.xgb5 0.004 0.118 0.114 93
+TMLE2.xgb5 0.007 0.055 0.054 93
+DR1.xgb1 0.004 0.118 0.114 93
+DR2.xgb1 0.007 0.055 0.053 93
+TMLE1.xgb1 0.003 0.118 0.114 93
+TMLE2.xgb1 0.005 0.055 0.053 93
+Table A4: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 3, where $E(n^{B})=5000$, $M=50$, $n_{\rm house}=20$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT 0.010 0.208 0.220 97
+Haj 0.004 0.112 0.112 94
+naive -0.303 0.102
+DR1 0.010 0.198 0.210 96
+DR2clw 0.004 0.089 0.092 94
+TMLE1 0.009 0.198 0.210 96
+TMLE2 0.003 0.089 0.091 94
+KH 0.009 0.198 0.210 96
+DR1.hal5 0.017 0.196 0.209 96
+DR2.hal5 0.011 0.089 0.089 94
+TMLE1.hal5 0.016 0.196 0.208 96
+TMLE2.hal5 0.010 0.089 0.090 94
+DR1.hal1 0.015 0.197 0.209 96
+DR2.hal1 0.009 0.089 0.090 94
+TMLE1.hal1 0.015 0.197 0.209 96
+TMLE2.hal1 0.009 0.089 0.090 94
+DR1.xgb5 0.020 0.197 0.209 96
+DR2.xgb5 0.014 0.089 0.089 94
+TMLE1.xgb5 0.019 0.197 0.209 96
+TMLE2.xgb5 0.013 0.088 0.089 94
+DR1.xgb1 0.019 0.198 0.209 96
+DR2.xgb1 0.013 0.089 0.089 94
+TMLE1.xgb1 0.018 0.198 0.209 96
+TMLE2.xgb1 0.012 0.089 0.089 94
+Table A5: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 4, where $E(n^{B})=5000$, $M=50$, $n_{\rm house}=5$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT 0.005 0.069 0.068 95
+Haj 0.001 0.035 0.035 94
+naive -0.306 0.034
+DR1 0.003 0.075 0.078 96
+DR2clw 0.000 0.051 0.051 94
+TMLE1 0.003 0.075 0.078 96
+TMLE2 0.000 0.051 0.051 94
+KH 0.003 0.075 0.077 96
+DR1.hal5 0.013 0.075 0.075 96
+DR2.hal5 0.010 0.051 0.048 94
+TMLE1.hal5 0.013 0.076 0.075 95
+TMLE2.hal5 0.010 0.051 0.048 93
+DR1.hal1 0.013 0.075 0.075 96
+DR2.hal1 0.009 0.051 0.048 94
+TMLE1.hal1 0.011 0.075 0.075 96
+TMLE2.hal1 0.008 0.051 0.048 94
+DR1.xgb5 0.017 0.078 0.076 95
+DR2.xgb5 0.014 0.053 0.048 92
+TMLE1.xgb5 0.016 0.077 0.076 94
+TMLE2.xgb5 0.013 0.053 0.048 91
+DR1.xgb1 0.021 0.077 0.074 94
+DR2.xgb1 0.018 0.052 0.045 89
+TMLE1.xgb1 0.018 0.077 0.074 95
+TMLE2.xgb1 0.014 0.053 0.045 89
+Table A6: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 5, where $E(n^{B})=1000$, $M=150$, $n_{\rm house}=20$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT -0.006 0.127 0.127 95
+Haj -0.009 0.064 0.064 96
+naive -0.313 0.061
+DR1 0.002 0.130 0.128 94
+DR2clw -0.002 0.069 0.067 93
+TMLE1 0.002 0.129 0.128 94
+TMLE2 -0.002 0.069 0.067 93
+KH 0.002 0.129 0.128 94
+DR1.hal5 0.016 0.127 0.125 94
+DR2.hal5 0.013 0.068 0.062 92
+TMLE1.hal5 0.016 0.128 0.125 94
+TMLE2.hal5 0.012 0.069 0.062 91
+DR1.hal1 0.014 0.127 0.125 94
+DR2.hal1 0.011 0.068 0.062 91
+TMLE1.hal1 0.013 0.128 0.125 94
+TMLE2.hal1 0.009 0.068 0.062 90
+DR1.xgb5 0.024 0.129 0.125 93
+DR2.xgb5 0.021 0.069 0.062 88
+TMLE1.xgb5 0.023 0.129 0.125 94
+TMLE2.xgb5 0.020 0.069 0.062 90
+DR1.xgb1 0.025 0.128 0.125 93
+DR2.xgb1 0.021 0.068 0.060 89
+TMLE1.xgb1 0.022 0.128 0.124 93
+TMLE2.xgb1 0.018 0.068 0.060 90
+Table A7: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 6, where $E(n^{B})=1000$, $M=150$, $n_{\rm house}=5$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT -0.003 0.120 0.119 93
+Haj -0.000 0.060 0.062 94
+naive -0.308 0.058
+DR1 -0.003 0.123 0.121 93
+DR2clw -0.000 0.067 0.067 96
+TMLE1 -0.003 0.123 0.121 93
+TMLE2 -0.000 0.067 0.067 96
+KH -0.003 0.123 0.121 93
+DR1.hal5 0.010 0.122 0.118 93
+DR2.hal5 0.013 0.067 0.063 91
+TMLE1.hal5 0.010 0.122 0.119 93
+TMLE2.hal5 0.013 0.067 0.064 93
+DR1.hal1 0.009 0.123 0.118 93
+DR2.hal1 0.012 0.067 0.062 92
+TMLE1.hal1 0.008 0.123 0.118 92
+TMLE2.hal1 0.011 0.067 0.063 92
+DR1.xgb5 0.018 0.121 0.118 94
+DR2.xgb5 0.021 0.065 0.062 93
+TMLE1.xgb5 0.017 0.121 0.119 94
+TMLE2.xgb5 0.020 0.065 0.063 94
+DR1.xgb1 0.019 0.122 0.117 92
+DR2.xgb1 0.022 0.065 0.060 90
+TMLE1.xgb1 0.017 0.122 0.118 92
+TMLE2.xgb1 0.019 0.065 0.061 91
+Table A8: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 7, where $E(n^{B})=1000$, $M=50$, $n_{\rm house}=20$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT 0.010 0.208 0.220 97
+Haj 0.004 0.112 0.112 94
+naive -0.303 0.102
+DR1 0.009 0.203 0.214 96
+DR2clw 0.003 0.099 0.101 94
+TMLE1 0.009 0.203 0.214 96
+TMLE2 0.003 0.099 0.100 94
+KH 0.009 0.202 0.213 95
+DR1.hal5 0.028 0.200 0.209 95
+DR2.hal5 0.022 0.098 0.093 92
+TMLE1.hal5 0.027 0.201 0.209 95
+TMLE2.hal5 0.021 0.099 0.094 92
+DR1.hal1 0.025 0.201 0.209 95
+DR2.hal1 0.019 0.098 0.093 92
+TMLE1.hal1 0.024 0.201 0.209 95
+TMLE2.hal1 0.018 0.099 0.094 92
+DR1.xgb5 0.037 0.203 0.210 95
+DR2.xgb5 0.031 0.099 0.093 92
+TMLE1.xgb5 0.036 0.203 0.210 95
+TMLE2.xgb5 0.030 0.099 0.094 92
+DR1.xgb1 0.036 0.203 0.210 95
+DR2.xgb1 0.030 0.098 0.092 91
+TMLE1.xgb1 0.033 0.203 0.210 95
+TMLE2.xgb1 0.027 0.098 0.092 91
+Table A9: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 8, where $E(n^{B})=1000$, $M=50$, $n_{\rm house}=5$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT 0.005 0.069 0.068 95
+Haj 0.001 0.035 0.035 94
+naive -0.306 0.034
+DR1 0.007 0.088 0.088 96
+DR2clw 0.003 0.066 0.066 94
+TMLE1 0.007 0.088 0.088 96
+TMLE2 0.003 0.066 0.066 94
+KH 0.007 0.088 0.087 95
+DR1.hal5 0.020 0.094 0.087 94
+DR2.hal5 0.017 0.075 0.065 92
+TMLE1.hal5 0.023 0.090 0.085 92
+TMLE2.hal5 0.020 0.069 0.062 91
+DR1.hal1 0.022 0.088 0.083 94
+DR2.hal1 0.019 0.067 0.060 91
+TMLE1.hal1 0.020 0.088 0.083 94
+TMLE2.hal1 0.016 0.067 0.060 91
+DR1.xgb5 0.026 0.091 0.086 92
+DR2.xgb5 0.023 0.070 0.064 90
+TMLE1.xgb5 0.027 0.092 0.086 91
+TMLE2.xgb5 0.023 0.071 0.063 90
+DR1.xgb1 0.034 0.089 0.080 88
+DR2.xgb1 0.031 0.068 0.056 86
+TMLE1.xgb1 0.030 0.090 0.080 90
+TMLE2.xgb1 0.027 0.069 0.056 88
+Table A10: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 9, where $E(n^{B})=500$, $M=150$, $n_{\rm house}=20$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT -0.006 0.127 0.127 95
+Haj -0.009 0.064 0.064 96
+naive -0.313 0.061
+DR1 0.001 0.137 0.136 94
+DR2clw -0.002 0.082 0.080 93
+TMLE1 0.002 0.137 0.136 94
+TMLE2 -0.002 0.082 0.080 93
+KH 0.002 0.137 0.135 94
+DR1.hal5 0.023 0.137 0.131 92
+DR2.hal5 0.019 0.083 0.073 90
+TMLE1.hal5 0.023 0.138 0.130 92
+TMLE2.hal5 0.019 0.085 0.073 87
+DR1.hal1 0.022 0.136 0.130 92
+DR2.hal1 0.019 0.083 0.071 89
+TMLE1.hal1 0.019 0.137 0.130 92
+TMLE2.hal1 0.015 0.083 0.072 89
+DR1.xgb5 0.031 0.137 0.131 92
+DR2.xgb5 0.027 0.084 0.073 88
+TMLE1.xgb5 0.031 0.137 0.131 91
+TMLE2.xgb5 0.027 0.084 0.073 88
+DR1.xgb1 0.036 0.135 0.128 91
+DR2.xgb1 0.033 0.081 0.067 84
+TMLE1.xgb1 0.033 0.136 0.128 92
+TMLE2.xgb1 0.030 0.082 0.068 85
+Table A11: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 10, where $E(n^{B})=500$, $M=150$, $n_{\rm house}=5$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT -0.003 0.120 0.119 93
+Haj -0.000 0.060 0.062 94
+naive -0.308 0.058
+DR1 -0.002 0.131 0.129 93
+DR2clw 0.000 0.080 0.080 96
+TMLE1 -0.002 0.130 0.129 93
+TMLE2 0.000 0.080 0.080 96
+KH -0.002 0.131 0.128 94
+DR1.hal5 0.013 0.133 0.125 93
+DR2.hal5 0.016 0.083 0.075 92
+TMLE1.hal5 0.016 0.131 4.014 92
+TMLE2.hal5 0.019 0.081 3.964 92
+DR1.hal1 0.016 0.131 0.123 92
+DR2.hal1 0.018 0.081 0.072 90
+TMLE1.hal1 0.013 0.131 0.123 92
+TMLE2.hal1 0.016 0.081 0.072 90
+DR1.xgb5 0.024 0.132 0.125 92
+DR2.xgb5 0.027 0.081 0.074 91
+TMLE1.xgb5 0.025 0.132 0.125 92
+TMLE2.xgb5 0.028 0.081 0.075 91
+DR1.xgb1 0.030 0.130 0.121 92
+DR2.xgb1 0.033 0.079 0.068 87
+TMLE1.xgb1 0.027 0.130 0.121 92
+TMLE2.xgb1 0.030 0.080 0.068 88
+Table A12: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 11, where $E(n^{B})=500$, $M=50$, $n_{\rm house}=20$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT 0.010 0.208 0.220 97
+Haj 0.004 0.112 0.112 94
+naive -0.303 0.102
+DR1 0.010 0.206 0.220 95
+DR2clw 0.004 0.106 0.111 95
+TMLE1 0.010 0.205 0.219 96
+TMLE2 0.004 0.106 0.110 95
+KH 0.010 0.205 0.217 95
+DR1.hal5 0.040 0.213 0.214 96
+DR2.hal5 0.030 0.106 0.099 91
+TMLE1.hal5 0.034 0.203 0.211 95
+TMLE2.hal5 0.028 0.106 0.100 91
+DR1.hal1 0.033 0.203 0.211 96
+DR2.hal1 0.027 0.105 0.099 93
+TMLE1.hal1 0.030 0.204 0.211 96
+TMLE2.hal1 0.024 0.105 0.099 93
+DR1.xgb5 0.047 0.205 0.213 96
+DR2.xgb5 0.041 0.107 0.100 92
+TMLE1.xgb5 0.047 0.206 0.213 96
+TMLE2.xgb5 0.041 0.106 0.101 92
+DR1.xgb1 0.049 0.206 0.211 96
+DR2.xgb1 0.043 0.105 0.096 91
+TMLE1.xgb1 0.047 0.206 0.211 96
+TMLE2.xgb1 0.041 0.106 0.097 91
+Table A13: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 12, where $E(n^{B})=500$, $M=50$, $n_{\rm house}=5$ and parametric nuisance models are correctly specified.
+bias empSE SEhat cover
+HT 0.003 0.045 0.044 95
+Haj 0.001 0.033 0.032 95
+naive -0.151 0.029
+DR1 0.102 0.043 0.050 45
+DR2clw 0.099 0.030 0.031 11
+TMLE1 0.098 0.043 0.049 48
+TMLE2 0.096 0.030 0.031 12
+KH 0.085 0.043 0.044 51
+DR1.hal5 0.006 0.043 0.043 94
+DR2.hal5 0.004 0.031 0.031 96
+TMLE1.hal5 0.006 0.043 0.043 94
+TMLE2.hal5 0.005 0.031 0.031 96
+DR1.hal1 0.006 0.044 0.043 95
+DR2.hal1 0.004 0.031 0.031 95
+TMLE1.hal1 0.006 0.044 0.043 95
+TMLE2.hal1 0.004 0.031 0.031 95
+DR1.xgb5 0.011 0.044 0.043 95
+DR2.xgb5 0.009 0.031 0.031 92
+TMLE1.xgb5 0.010 0.044 0.043 94
+TMLE2.xgb5 0.009 0.031 0.031 94
+DR1.xgb1 0.011 0.044 0.043 94
+DR2.xgb1 0.009 0.031 0.030 93
+TMLE1.xgb1 0.011 0.044 0.043 94
+TMLE2.xgb1 0.009 0.031 0.030 93
+Table A14: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 13, where $E(n^{B})=5000$, $M=150$, $n_{\rm house}=20$ and parametric nuisance models are misspecified.
+bias empSE SEhat cover
+HT -0.007 0.084 0.084 96
+Haj -0.008 0.063 0.062 94
+naive -0.158 0.056
+DR1 0.099 0.074 0.087 84
+DR2clw 0.095 0.048 0.051 52
+TMLE1 0.094 0.074 0.087 86
+TMLE2 0.092 0.048 0.050 54
+KH 0.080 0.075 0.077 85
+DR1.hal5 0.003 0.075 0.076 96
+DR2.hal5 0.001 0.051 0.051 96
+TMLE1.hal5 0.002 0.075 0.076 95
+TMLE2.hal5 0.001 0.051 0.051 96
+DR1.hal1 0.002 0.074 0.076 96
+DR2.hal1 0.000 0.051 0.051 96
+TMLE1.hal1 0.002 0.074 0.076 96
+TMLE2.hal1 0.000 0.051 0.051 96
+DR1.xgb5 0.007 0.074 0.076 95
+DR2.xgb5 0.005 0.050 0.051 96
+TMLE1.xgb5 0.006 0.074 0.076 96
+TMLE2.xgb5 0.005 0.050 0.051 96
+DR1.xgb1 0.006 0.074 0.076 96
+DR2.xgb1 0.005 0.050 0.051 96
+TMLE1.xgb1 0.006 0.074 0.076 96
+TMLE2.xgb1 0.004 0.050 0.051 96
+Table A15: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 14, where $E(n^{B})=5000$, $M=150$, $n_{\rm house}=5$ and parametric nuisance models are misspecified.
+bias empSE SEhat cover
+HT 0.000 0.076 0.077 95
+Haj 0.001 0.054 0.057 96
+naive -0.151 0.047
+DR1 0.099 0.070 0.082 82
+DR2clw 0.099 0.043 0.047 43
+TMLE1 0.094 0.071 0.081 83
+TMLE2 0.096 0.043 0.047 45
+KH 0.081 0.072 0.071 82
+DR1.hal5 0.004 0.072 0.071 94
+DR2.hal5 0.005 0.048 0.048 95
+TMLE1.hal5 0.004 0.072 0.071 93
+TMLE2.hal5 0.006 0.047 0.048 95
+DR1.hal1 0.003 0.072 0.071 94
+DR2.hal1 0.005 0.048 0.048 95
+TMLE1.hal1 0.003 0.072 0.071 94
+TMLE2.hal1 0.005 0.048 0.048 95
+DR1.xgb5 0.008 0.073 0.071 93
+DR2.xgb5 0.010 0.048 0.048 95
+TMLE1.xgb5 0.008 0.072 0.071 93
+TMLE2.xgb5 0.010 0.048 0.048 95
+DR1.xgb1 0.008 0.072 0.071 93
+DR2.xgb1 0.010 0.048 0.047 95
+TMLE1.xgb1 0.008 0.072 0.071 93
+TMLE2.xgb1 0.009 0.048 0.048 95
+Table A16: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 15, where $E(n^{B})=5000$, $M=50$, $n_{\rm house}=20$ and parametric nuisance models are misspecified.
+bias empSE SEhat cover
+HT 0.009 0.142 0.146 96
+Haj 0.005 0.109 0.109 94
+naive -0.147 0.092
+DR1 0.112 0.120 0.149 94
+DR2clw 0.106 0.075 0.084 79
+TMLE1 0.104 0.122 0.148 94
+TMLE2 0.101 0.077 0.083 81
+KH 0.089 0.124 0.131 92
+DR1.hal5 0.013 0.124 0.131 96
+DR2.hal5 0.010 0.084 0.087 94
+TMLE1.hal5 0.012 0.123 0.130 96
+TMLE2.hal5 0.009 0.083 0.086 95
+DR1.hal1 0.012 0.124 0.130 96
+DR2.hal1 0.009 0.083 0.086 94
+TMLE1.hal1 0.012 0.124 0.130 96
+TMLE2.hal1 0.009 0.083 0.086 94
+DR1.xgb5 0.017 0.125 0.131 95
+DR2.xgb5 0.014 0.084 0.087 95
+TMLE1.xgb5 0.017 0.125 0.131 95
+TMLE2.xgb5 0.014 0.084 0.087 95
+DR1.xgb1 0.017 0.125 0.131 96
+DR2.xgb1 0.014 0.084 0.086 94
+TMLE1.xgb1 0.016 0.125 0.131 96
+TMLE2.xgb1 0.013 0.084 0.087 95
+Table A17: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 16, where $E(n^{B})=5000$, $M=50$, $n_{\rm house}=5$ and parametric nuisance models are misspecified.
+bias empSE SEhat cover
+HT 0.003 0.045 0.044 95
+Haj 0.001 0.033 0.032 95
+naive -0.151 0.029
+DR1 0.103 0.057 0.064 64
+DR2clw 0.100 0.048 0.050 48
+TMLE1 0.099 0.056 0.063 66
+TMLE2 0.097 0.048 0.050 50
+KH 0.084 0.056 0.056 67
+DR1.hal5 0.009 0.055 0.055 96
+DR2.hal5 0.007 0.047 0.046 95
+TMLE1.hal5 0.010 0.055 0.055 95
+TMLE2.hal5 0.008 0.047 0.046 94
+DR1.hal1 0.009 0.055 0.054 96
+DR2.hal1 0.007 0.046 0.045 94
+TMLE1.hal1 0.008 0.055 0.054 96
+TMLE2.hal1 0.007 0.047 0.045 94
+DR1.xgb5 0.016 0.055 0.054 95
+DR2.xgb5 0.015 0.047 0.045 94
+TMLE1.xgb5 0.016 0.056 0.055 95
+TMLE2.xgb5 0.015 0.047 0.045 95
+DR1.xgb1 0.019 0.055 0.053 94
+DR2.xgb1 0.017 0.047 0.043 93
+TMLE1.xgb1 0.017 0.055 0.053 94
+TMLE2.xgb1 0.016 0.047 0.043 93
+Table A18: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 17, where $E(n^{B})=1000$, $M=150$, $n_{\rm house}=20$ and parametric nuisance models are misspecified.
+bias empSE SEhat cover
+HT -0.007 0.084 0.084 96
+Haj -0.008 0.063 0.062 94
+naive -0.158 0.056
+DR1 0.105 0.083 0.096 87
+DR2clw 0.101 0.061 0.065 67
+TMLE1 0.099 0.084 0.095 88
+TMLE2 0.097 0.061 0.064 70
+KH 0.083 0.084 0.085 87
+DR1.hal5 0.011 0.083 0.083 94
+DR2.hal5 0.009 0.063 0.061 93
+TMLE1.hal5 0.011 0.083 0.083 94
+TMLE2.hal5 0.009 0.062 0.061 93
+DR1.hal1 0.009 0.083 0.083 94
+DR2.hal1 0.007 0.062 0.060 93
+TMLE1.hal1 0.009 0.083 0.083 94
+TMLE2.hal1 0.007 0.062 0.060 93
+DR1.xgb5 0.018 0.083 0.083 94
+DR2.xgb5 0.016 0.062 0.061 94
+TMLE1.xgb5 0.017 0.083 0.083 94
+TMLE2.xgb5 0.015 0.062 0.061 94
+DR1.xgb1 0.018 0.083 0.082 94
+DR2.xgb1 0.016 0.063 0.059 92
+TMLE1.xgb1 0.016 0.083 0.082 94
+TMLE2.xgb1 0.015 0.063 0.059 93
+Table A19: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 18, where $E(n^{B})=1000$, $M=150$, $n_{\rm house}=5$ and parametric nuisance models are misspecified.
+bias empSE SEhat cover
+HT 0.000 0.076 0.077 95
+Haj 0.001 0.054 0.057 96
+naive -0.151 0.047
+DR1 0.103 0.079 0.092 85
+DR2clw 0.103 0.056 0.062 64
+TMLE1 0.097 0.079 0.090 85
+TMLE2 0.099 0.056 0.061 66
+KH 0.082 0.079 0.080 83
+DR1.hal5 0.009 0.078 0.079 94
+DR2.hal5 0.011 0.058 0.059 95
+TMLE1.hal5 0.011 0.077 0.079 94
+TMLE2.hal5 0.012 0.056 0.059 96
+DR1.hal1 0.009 0.077 0.078 94
+DR2.hal1 0.010 0.056 0.057 95
+TMLE1.hal1 0.009 0.077 0.078 94
+TMLE2.hal1 0.010 0.056 0.058 95
+DR1.xgb5 0.017 0.079 0.078 94
+DR2.xgb5 0.019 0.058 0.058 93
+TMLE1.xgb5 0.017 0.079 0.079 94
+TMLE2.xgb5 0.018 0.058 0.059 94
+DR1.xgb1 0.019 0.079 0.077 92
+DR2.xgb1 0.020 0.057 0.056 92
+TMLE1.xgb1 0.017 0.079 0.077 94
+TMLE2.xgb1 0.018 0.057 0.056 92
+Table A20: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 19, where $E(n^{B})=1000$, $M=50$, $n_{\rm house}=20$ and parametric nuisance models are misspecified.
+bias empSE SEhat cover
+HT 0.009 0.142 0.146 96
+Haj 0.005 0.109 0.109 94
+naive -0.147 0.092
+DR1 0.118 0.129 0.155 93
+DR2clw 0.112 0.087 0.094 81
+TMLE1 0.109 0.131 0.154 93
+TMLE2 0.106 0.088 0.093 82
+KH 0.092 0.133 0.136 92
+DR1.hal5 0.022 0.135 0.136 95
+DR2.hal5 0.018 0.098 0.094 93
+TMLE1.hal5 0.022 0.131 0.135 96
+TMLE2.hal5 0.019 0.093 0.093 94
+DR1.hal1 0.020 0.130 0.134 95
+DR2.hal1 0.017 0.092 0.091 94
+TMLE1.hal1 0.020 0.130 0.134 95
+TMLE2.hal1 0.017 0.092 0.091 94
+DR1.xgb5 0.031 0.132 0.135 95
+DR2.xgb5 0.027 0.094 0.092 92
+TMLE1.xgb5 0.030 0.132 0.135 95
+TMLE2.xgb5 0.027 0.094 0.092 94
+DR1.xgb1 0.029 0.132 0.134 95
+DR2.xgb1 0.026 0.093 0.091 93
+TMLE1.xgb1 0.028 0.132 0.134 95
+TMLE2.xgb1 0.024 0.093 0.091 93
+Table A21: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 20, where $E(n^{B})=1000$, $M=50$, $n_{\rm house}=5$ and parametric nuisance models are misspecified.
+bias empSE SEhat cover
+HT 0.003 0.045 0.044 95
+Haj 0.001 0.033 0.032 95
+naive -0.151 0.029
+DR1 0.105 0.070 0.078 75
+DR2clw 0.102 0.063 0.067 69
+TMLE1 0.100 0.069 0.077 77
+TMLE2 0.098 0.063 0.066 70
+KH 0.084 0.068 0.069 79
+DR1.hal5 0.016 0.068 0.068 95
+DR2.hal5 0.014 0.061 0.060 94
+TMLE1.hal5 0.015 0.067 0.067 94
+TMLE2.hal5 0.014 0.061 0.060 94
+DR1.hal1 0.012 0.065 0.065 95
+DR2.hal1 0.010 0.059 0.057 94
+TMLE1.hal1 0.011 0.067 0.065 95
+TMLE2.hal1 0.010 0.061 0.058 94
+DR1.xgb5 0.020 0.068 0.067 94
+DR2.xgb5 0.019 0.062 0.059 93
+TMLE1.xgb5 0.020 0.069 0.067 93
+TMLE2.xgb5 0.019 0.063 0.059 93
+DR1.xgb1 0.024 0.066 0.062 92
+DR2.xgb1 0.022 0.060 0.055 91
+TMLE1.xgb1 0.021 0.066 0.062 92
+TMLE2.xgb1 0.020 0.060 0.055 92
+Table A22: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 21, where $E(n^{B})=500$, $M=150$, $n_{\rm house}=20$ and parametric nuisance models are misspecified.
+bias empSE SEhat cover
+HT -0.007 0.084 0.084 96
+Haj -0.008 0.063 0.062 94
+naive -0.158 0.056
+DR1 0.109 0.095 0.107 88
+DR2clw 0.105 0.076 0.079 75
+TMLE1 0.103 0.095 0.105 88
+TMLE2 0.101 0.075 0.078 77
+KH 0.086 0.095 0.094 88
+DR1.hal5 0.019 0.095 0.091 94
+DR2.hal5 0.018 0.076 0.072 92
+TMLE1.hal5 0.019 0.095 0.091 93
+TMLE2.hal5 0.017 0.076 0.071 92
+DR1.hal1 0.016 0.094 0.090 94
+DR2.hal1 0.014 0.075 0.070 92
+TMLE1.hal1 0.015 0.095 0.091 94
+TMLE2.hal1 0.013 0.078 0.071 92
+DR1.xgb5 0.026 0.096 0.091 94
+DR2.xgb5 0.024 0.077 0.071 91
+TMLE1.xgb5 0.026 0.097 0.091 94
+TMLE2.xgb5 0.025 0.078 0.071 91
+DR1.xgb1 0.029 0.094 0.088 93
+DR2.xgb1 0.027 0.076 0.067 90
+TMLE1.xgb1 0.027 0.094 0.088 93
+TMLE2.xgb1 0.025 0.076 0.067 90
+Table A23: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 22, where $E(n^{B})=500$, $M=150$, $n_{\rm house}=5$ and parametric nuisance models are misspecified.
+bias empSE SEhat cover
+HT 0.000 0.076 0.077 95
+Haj 0.001 0.054 0.057 96
+naive -0.151 0.047
+DR1 0.102 0.091 0.102 86
+DR2clw 0.101 0.073 0.076 74
+TMLE1 0.096 0.090 0.101 87
+TMLE2 0.097 0.072 0.075 76
+KH 0.080 0.089 0.090 86
+DR1.hal5 0.011 0.087 0.087 94
+DR2.hal5 0.013 0.070 0.069 94
+TMLE1.hal5 0.014 0.087 0.088 94
+TMLE2.hal5 0.016 0.070 0.070 95
+DR1.hal1 0.010 0.087 0.086 95
+DR2.hal1 0.012 0.069 0.068 94
+TMLE1.hal1 0.010 0.087 0.086 95
+TMLE2.hal1 0.011 0.069 0.068 94
+DR1.xgb5 0.019 0.088 0.087 94
+DR2.xgb5 0.021 0.070 0.069 94
+TMLE1.xgb5 0.020 0.087 0.088 94
+TMLE2.xgb5 0.022 0.070 0.070 94
+DR1.xgb1 0.024 0.087 0.084 94
+DR2.xgb1 0.025 0.069 0.065 92
+TMLE1.xgb1 0.022 0.087 0.084 94
+TMLE2.xgb1 0.023 0.069 0.065 92
+Table A24: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 23, where $E(n^{B})=500$, $M=50$, $n_{\rm house}=20$ and parametric nuisance models are misspecified.
+bias empSE SEhat cover
+HT 0.009 0.142 0.146 96
+Haj 0.005 0.109 0.109 94
+naive -0.147 0.092
+DR1 0.118 0.134 0.163 93
+DR2clw 0.111 0.095 0.106 83
+TMLE1 0.108 0.135 0.161 94
+TMLE2 0.105 0.095 0.104 84
+KH 0.091 0.137 0.142 92
+DR1.hal5 0.027 0.138 0.139 94
+DR2.hal5 0.024 0.103 0.098 93
+TMLE1.hal5 0.026 0.138 0.139 94
+TMLE2.hal5 0.023 0.102 0.099 92
+DR1.hal1 0.024 0.137 0.138 93
+DR2.hal1 0.021 0.101 0.097 93
+TMLE1.hal1 0.023 0.137 0.138 93
+TMLE2.hal1 0.020 0.101 0.098 93
+DR1.xgb5 0.039 0.140 0.139 94
+DR2.xgb5 0.036 0.103 0.099 92
+TMLE1.xgb5 0.038 0.139 0.140 94
+TMLE2.xgb5 0.035 0.103 0.100 92
+DR1.xgb1 0.038 0.139 0.137 92
+DR2.xgb1 0.034 0.102 0.096 91
+TMLE1.xgb1 0.036 0.139 0.137 93
+TMLE2.xgb1 0.033 0.102 0.096 91
+Table A25: Bias, empirical SE (‘empSE’), mean estimated SE (‘SEhat’), and CI coverage (‘cover’) for Scenario 24, where $E(n^{B})=500$, $M=50$, $n_{\rm house}=5$ and parametric nuisance models are misspecified.
+A11 Conditional Poisson sampling of clusters
+For simplicity of notation in this section, we shall use $R_{j}$ to denote $R^{C}_{j}$, i.e. the binary indicator that cluster $j$ is included in Sample A.
+If $M$ clusters are sampled from $J$ clusters using conditional Poisson sampling with binomial sampling probabilities $\pi_{1},\ldots,\pi_{J}$, then
+$P(R_{1}=r_{1},\ldots,R_{J}=r_{J}\mid\mathcal{F}_{J})$ $=$ $\frac{\prod_{j=1}^{J}\pi_{j}^{r_{j}}(1-\pi_{j})^{1-r_{j}}}{\sum_{(r_{1}^{\prime},\ldots,r_{J}^{\prime})}I\left(\sum_{j=1}^{J}r_{j}^{\prime}=M\right)\prod_{j=1}^{J}\pi_{j}^{r_{j}^{\prime}}(1-\pi^{C}_{j})^{1-r_{j}^{\prime}}}$
+if $\sum_{j=1}^{J}r_{j}=M$, and $P(R_{1}=r_{1},\ldots,R_{J}=r_{J}\mid\mathcal{F}_{J})=0$ otherwise. We shall denote this distribution as
+$(R_{1},\ldots,R_{J})\mid\mathcal{F}_{J}\sim\mbox{condPois}\big((\pi_{1},\ldots,\pi_{J}),\;M\big).$
+Henceforth, for convenience, we shall omit explicit conditioning on $\mathcal{F}_{J}$ in expressions for probabilities, but it should be understood that this conditioning is implicit.
+It follows from equation (LABEL:eq:cond.Pois) that, for any $(r_{1},\ldots,r_{J})$ such that $\sum_{j=1}^{J}r_{j}=M$,
+$P(R_{1}=r_{1},\ldots,R_{k}=r_{k}\mid R_{k+1}=r_{k+1},\ldots,R_{J}=r_{J})$
+$\hskip 14.22636pt=\frac{P(R_{1}=r_{1},\ldots,R_{J}=r_{J})}{P(R_{k+1}=r_{k+1},\ldots,R_{J}=r_{J})}$
+$\hskip 14.22636pt=\frac{\prod_{j=1}^{J}\pi_{j}^{r_{j}}(1-\pi_{j})^{1-r_{j}}}{\sum_{(r_{1}^{\prime},\ldots,r_{k}^{\prime})}I\left(\sum_{j=1}^{k}r_{j}^{\prime}=M-\sum_{j=k+1}^{J}r_{j}\right)\prod_{j=1}^{J}\pi_{j}^{r_{j}^{\prime}}(1-\pi_{j})^{1-r_{j}^{\prime}}}$
+$\hskip 14.22636pt=\frac{\prod_{j=1}^{J}\pi_{j}^{r_{j}}(1-\pi_{j})^{1-r_{j}}}{\prod_{j=k+1}^{J}\pi_{j}^{r_{j}}(1-\pi_{j})^{1-r_{j}}\sum_{(r_{1}^{\prime},\ldots,r_{k}^{\prime})}I\left(\sum_{j=1}^{k}r_{j}^{\prime}=M-\sum_{j=1}^{k}r_{j}\right)\prod_{j=1}^{k}\pi_{j}^{r_{j}^{\prime}}(1-\pi_{j})^{1-r_{j}^{\prime}}}$
+$\hskip 14.22636pt=\frac{\prod_{j=1}^{k}\pi_{j}^{r_{j}}(1-\pi_{j})^{1-r_{j}}}{\sum_{(r_{1}^{\prime},\ldots,r_{k}^{\prime})}I\left(\sum_{j=1}^{k}r_{j}^{\prime}=M-\sum_{j=1}^{k}r_{j}\right)\prod_{j=1}^{k}\pi_{j}^{r_{j}^{\prime}}(1-\pi_{j})^{1-r_{j}^{\prime}}}.$
+We recognise this as
+$R_{1},\ldots,R_{k}\mid R_{k+1},\ldots,R_{J},\mathcal{F}_{J}\sim\mbox{condPois}\left((\pi_{1},\ldots,\pi_{k}),\;M-\sum_{j=k+1}^{J}R_{j}\right).$
+More generally, for any subset $\mathcal{S}$ of $\{1,\ldots,J\}$, we have
+$\{R_{j}:j\in\mathcal{S}\}\mid\{R_{j}:j\notin\mathcal{S}\},\mathcal{F}_{J}\sim\mbox{condPois}\left(\{\pi_{j}:j\in\mathcal{S}\},\;M-\sum_{j\notin\mathcal{S}}R_{j}\right).$
+Now assume that $M$ is an integer multiple of $K$.
+Let $\mathcal{S}^{C}_{k}$ denote the set of clusters in fold $k$.
+Because of the way the folds are chosen, we know that fold $k$ will contain $M/K$ sampled clusters.
+So, for any $\{r_{1},\ldots,r_{J}\}$ such that $\sum_{j=1}^{J}r_{j}=M$ and for any set $\mathcal{S}$ of $\lfloor J/K\rfloor$ or $\lceil J/K\rceil$ clusters such that $\sum_{j\in\mathcal{S}}r_{j}=M/K$ (and thus $\sum_{j\notin\mathcal{S}}r_{j}=M(K-1)/K$), we can use Bayes’ Rule to show
+$P(\{R_{j}=r_{j}:j\in\mathcal{S}\}\mid\{R_{j}=r_{j}:j\notin\mathcal{S}\},\mathcal{S}^{C}_{k}=\mathcal{S})$
+$\hskip 5.69046pt=\frac{P(\mathcal{S}_{k}^{C}=\mathcal{S}\mid R_{1}=r_{1},\ldots,R_{J}=r_{j})\times P(\{R_{j}=r_{j}:j\in\mathcal{S}\}\mid\{R_{j}=r_{j}:j\notin\mathcal{S}\})}{P(\mathcal{S}^{C}_{k}=\mathcal{S}\mid\{R_{j}=r_{j}:j\notin\mathcal{S}\})}$
+$\hskip 5.69046pt=\frac{P(\mathcal{S}^{C}_{k}=\mathcal{S}\mid\sum_{j\in\mathcal{S}}R_{j}=M/K)\times P(\{R_{j}=r_{j}:j\in\mathcal{S}\}\mid\{R_{j}=r_{j}:j\notin\mathcal{S}\})}{P(\mathcal{S}^{C}_{k}=\mathcal{S}\mid\sum_{j\notin\mathcal{S}}R_{j}=M(K-1)/K)}$
+$\hskip 5.69046pt=\frac{P(\mathcal{S}^{C}_{k}=\mathcal{S}\mid\sum_{j\in\mathcal{S}}R_{j}=M/K)\times P(\{R_{j}=r_{j}:j\in\mathcal{S}\}\mid\{R_{j}=r_{j}:j\notin\mathcal{S}\})}{P(\mathcal{S}^{C}_{k}=\mathcal{S}\mid\sum_{j\in\mathcal{S}}R_{j}=M/K)}$
+$\hskip 5.69046pt=P(\{R_{j}=r_{j}:j\in\mathcal{S}\}\mid\{R_{j}=r_{j}:j\notin\mathcal{S}\}).$
+Therefore
+$\{R_{j}:j\in\mathcal{S}^{C}_{k}\}\mid\{R_{j}:j\notin\mathcal{S}^{C}_{k}\},\mathcal{F}_{J},\mathcal{S}^{C}_{k}\sim\mbox{condPois}\left(\{\pi_{j}:j\in\mathcal{S}^{C}_{k}\},\;M-\sum_{j\notin\mathcal{S}^{C}_{k}}R_{j}\right).$
+By the way the folds have been chosen, $\sum_{j\notin\mathcal{S}^{C}_{k}}R_{j}$ must equal M(K1)/KM(K-1)/K, and so we have
+$\{R_{j}:j\in\mathcal{S}^{C}_{k}\}\mid\{R_{j}:j\notin\mathcal{S}^{C}_{k}\},\mathcal{F}_{J},\mathcal{S}^{C}_{k}\sim\mbox{condPois}\left(\{\pi_{j}:j\in\mathcal{S}^{C}_{k}\},\;M/K\right).$
+We see therefore that $\{R_{j}:j\in\mathcal{S}^{C}_{k}\}\perp\!\!\!\perp\{R_{j}:j\notin\mathcal{S}^{C}_{k}\}\mid\mathcal{F}_{J},\mathcal{S}^{C}_{k}$.
+This implies that $\{R^{A}_{i}:i\in\mathcal{S}_{k}\}\perp\!\!\!\perp\{R^{A}_{i}:i\notin\mathcal{S}_{k}\}\mid\mathcal{F}_{J},\mathcal{S}_{k}$.
+Hence, $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ (which depend on $\{R^{A}_{i}:i\notin\mathcal{S}_{k}\}$) tell us nothing about $\{R^{A}_{i}:i\in\mathcal{S}_{k}\}$.